A Complete Bibliography of Publications in Ecological Modelling (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.04

Title word cross-reference

$125.00/Dfl [Mid84a]. $139.50 [Str84a]. $16.95 [LU82a]. $21.50 [Ush81b]. $25.00 [Str82a]. $26.00 [Ush84c]. $29.50 [Jør82a]. 3/2 [Bar89]. $31.00 [Jør80c]. $31.65 [Gro83a]. $32.00 [Jør80a]. $34.95 [Gra88a]. $35.00 [Gat83]. $35.50 [Col83]. $43.00 [Jør81d]. $43.80 [Ush81a]. $46.10 [Nih81]. $47.50 [Koh89]. $48.00 [Gro80]. $49.00 [Jør87d]. $51.50 [Ush84a]. $58.50 [Gro83b]. $59.50 [Jør87a]. $61.00 [Gro81]. $75.00 [Saw83]. $75.50/Dfl [Ano82h]. $78.00 [Sko86b]. $93.00/Dfl [Nih84a]. _{2} [Cou81, Dah85, KA84]. _{3} [RBM89, Sch85]. _{c} [GÅ89]. K [Bar89]. × [KHF88]. x [Fit84, Ush84a].

/0 [Jør82c]. /0-521-29525-4 [Jør82c].

0 [Ano82h, Ano87i, Baz86, BK86, Col83, ES85, Gat83, Gat87, Gat88, Gra88a, Gra89, Gro81, Gro83b, Gro84b, Gro84c, Gro85, Gue83, Hal86a, Hof84, Jac87, Jam87, Jør81d, Jør82a, Jør83, Jør86c, Jør86d, Jør87a, Jør87b, Jør87c, Jør87d, Jør87e, Jør87f, Koh89, Leg84b, Lho87, Log87, Lon86, Lon87, Mau87, Mau88, Mej83, Mid84a, Mid84b, Mit87, Nih82, Nih83b, Nih84d, Nih84a, Nih84b, Pac88, Saw83, SB85, Sko86a, Sko86c, Sko88, Ste88, Str82a, Str84a, Svi88a, Svi88c, SV89b, Svi89, Tre89, Ush84a, Ush84b, Ush84c, Wel88, Zuc88, Zuc89b, Zuc89a, LU82a, Ush84d]. 0-02-948040-X [Gra88a]. 0-12-013913-8 [Ush84b]. 0-12-040906-2 [Ush84a]. 0-12-040909-7 [Jør87a]. 0-12-053040-6 [Jør87c]. 0-12-071350-0 [Lon87]. 0-12-318750-8 [Jør81d]. 0-12-348780-3 [Gat83]. 0-12-385280-3 [Ush84c]. 0-12-543615-7 [Ste88]. 0-12-554520-7 [Gat87]. 0-12-583980-4 [Jør87e]. 0-12-583980-4/0-12-583982-0 [Jør87e]. 0-12-583982-0 [Jør87e]. 0-13-709197-4 [Jør87f]. 0-306-40494-X [Gro84c]. 0-306-40589-X [LU82a]. 0-306-40820-1 [Saw83]. 0-306-41460-0 [Gro85]. 0-306-41877-0 [Jør87b]. 0-309-03484-1 [Hal86a]. 0-309-03675-5 [Sko88]. 0-333-30721-6 [Mej83]. 0-387-12677-5 [Ush85]. 0-387-13631-2 [Svi88b]. 0-387-16088-4 [Wel88]. 0-408-10637-9 [ES85]. 0-408-10868-1 [Sko86c]. 0-412-24360-1 [Nih83b]. 0-412-26750-0 [Pac88]. 0-44-42314-1 [Jør86c]. 0-444-41793-1 [Ano82h]. 0-444-41827-X [Gro81]. 0-444-41969-1 [Gro83b]. 0-444-42155-6 [Mid84a]. 0-444-42156-4 [Lho87]. 0-444-42179-3 [Baz86]. 0-444-42247-1 [Nih84d]. 0-444-42794-5 [Mau88]. 0-444-42936-0 [Svi89]. 0-444-86406-7 [Nih84a]. 0-444-99505-6 [Zuc89b]. 0-444-99567-6 [Mit87]. 0-444-99721-0 [Nih82]. 0-444-99731-8 [Str84a]. 0-471-04831-3 [Gro83a]. 0-471-10051-X [Gue83]. 0-471-10274-1 [Svi88a]. 0-471-10521-X [Jam87]. 0-471-10522-8 [Leg84b]. 0-471-27956-0 [Col83]. 0-471-27957-9 [Col83]. 0-471-65277-6 [Nih84b]. 0-471-83019-4 [Koh89]. 0-471-87394-2 [Gat88]. 0-471-87888-X [Sko86a]. 0-471-89236-X [Svi88c]. 0-471-90104-0 [Jør87d]. 0-471-90204-7 [SB85]. 0-471-90262-4 [Jør86d]. 0-471-90349-3 [Lon86]. 0-471-90558-5 [Ano87i]. 0-471-90625-5 [Jør87d]. 0-471-90625-5/Cloth [Jør87d]. 0-471-91462-2 [Zuc89a]. 0-521-22495-0 [Jør82c, Jør83]. 0-521-29525-4 [Jør82c]. 0-521-33390-3 [Gra89]. 0-8133-7145-7 [SV89b]. 0-8493-5952-X [Ush84d]. 0-85312-426-4 [Gro84b]. 0-85312-905-3 [Log87]. 0-87371-052-5 [Jac87]. 0-87933-347-2 [Jør82a]. 0-916150-44-5 [Mid84b]. 0739-5671 [Sko86b].

1 [Ano82h, Bla83, Gra88b, Gro83b, Hal86a, Jør86c, Nih83b, Nih84d, Saw83, Sko86c, Svi88a]. 1/ISBN [Sko86b]. 12.95/US$22.40. [Gro84b]. 120.- [Gro83b]. 12th [Gro83b]. 14.95/Cloth [Jør87d]. 15th [Jør86c]. 17.20/US [Ush84c]. 175.00/US$67.50 [Nih84d]. 18 [Ano83h]. 18th [Mau88]. 1940 [Ush81b]. 1978 [Gro80, Jør81a]. 1979 [Ano87h]. 1980 [Str82b]. 1982 [Str84b]. 1984 [Str87b]. 1UD [Hal86a].

2 [Gat88, Jør82a, Jør86b, Leg89a, MTC85, Nih85, Som84, Svi88b, Ush84a, Zuc89a]. 25.00/US$44.50 [Gra89]. 25/US [Ush84a]. 26 [Ano83h, Vid89]. 27.00/ [Ush84b]. 27.50/US$29.95 [Lon87]. 295 [Str82a]. 295-95719-0 [Str82a].

3 [Baz86, BK86, Gat83, Gra88b, Gra89, Gro80, Gro84a, Hof84, Jør86e, Leg89a, Log87, Lon86, Nih81, Nih85, Pal89, Sko85, Sko86b, Sko86d, Som84, Str87a, Svi88b, Ush81b, Ush81a, Ush84c, Ush85, Wul86, Zuc88, Ano87h]. 3$ [CMW84a]. 3-540 [Svi88b]. 3-540-08769-9 [Ush81b]. 3-540-09620-5 [Nih81]. 3-540-10566-2 [Som84]. 3-540-12013-0 [Hof84]. 3-540-12158-7 [Sko85]. 3-540-12454-3 [Jør86e]. 3-540-12677-5 [Ush85]. 3-540-12677-5/0-387-12677-5 [Ush85]. 3-540-12681-3 [Wul86]. 3-540-12804-2 [Nih85]. 3-540-12903-0 [BK86]. 3-540-13179-5 [Sko86d]. 3-540-15942-8 [Str87a]. 3-540-16089-2 [Leg89a]. 3-540-18243-8 [Pal89]. 3-540-90379-8 [Ush81a]. 3-540-96289-1 [Gra88b]. 3-540-96446-0 [Zuc88]. 3-540/0-387-13631-2 [Svi88b]. 3-7045-0061-5 [Gro84a]. 3-7045-0075-5 [Ano87h]. 3-7186-0187-7 [Sko86b]. 300.- [Str84a]. 34.00 [Jør87d]. 34.00/US [Jør87d]. 3rd [Gro84b].

4 [God86a, Gro84b, Jør86d, Jør87f, Koh89, Lho87, Wel88]. 4/0 [Jør87e]. 44 [Jør86e]. 49.00/approximately [Nih85]. 4A [Mid84a]. 4B [Nih84d].

5 [Ano87i, Fit84, Gro84a, Jac87, Mau88, Mid84b, Nih81, Sko86d, Sko88, Ush85, Ano87h]. 5/0 [Ush85]. 5/Cloth [Jør87d]. 53.00/US$68.00 [Lon87]. 540/0 [Svi88b].

6 [Gro83c, Jør87c, Mej83, Mid84a, Mit87, Nih84b, Zuc89b]. 65.00/US$32.50 [Tre89].

7 [Gat87, Jør87a, Nih84a, SB85, Sko85, Sko86b, Ste88, SV89b]. 78.00/U.S [Nih81].

8 [God86b, Jør81d, Leg84b, Leg84a, Pal89, Str84a, Str87a, Ush81a, Ush84b]. 87 [Gro80]. 87-87257-17-3 [Gro80].

9 [Col83, ES85, Ush81b]. 90 [Fit84, God86a, God86b, Jør86b, Tre89]. 90-220-0842-8 [God86b]. 90-220-0844-4 [God86a]. 90-220-0913-0 [Tre89]. 90-277-1753-2 [Jør86b]. 90-6193-636-5 [Fit84]. 91 [Mau87]. 91-86344-25-0 [Mau87]. 92 [Gro83c, Leg84a]. 92-3-101906-6 [Gro83c]. 92-3-101922-8 [Leg84a].

AAAS [SV89b]. ability [Bar82]. above [BHD89a, BHD89b]. above- [BHD89a, BHD89b]. aboveground [LHSS86]. abundance [Cer89]. abundant [Ush84c]. acacia [BS89]. Academic [Gat83, Gat87, Jør81d, Jør87a, Jør87c, Jør87e, Lon87, Sko86b, Ste88, Ush84a, Ush84b, Ush84c]. Academy [Hal86a, Sko88]. Acanthaster [Sey89]. acarid [WH87b]. access [KT89]. accessibility [MRBU89]. accompanied [Jør88]. accordance [Mau82b]. account [Smi87]. accumulating [Van87]. accumulation [Meh84, PG89, Urc84]. accumulations [AM87, AMB87]. accuracy [CS85, JCN85]. achievement [DLMP87c]. acid [Arp83, AR83, Fon84b, Fon84c, KEK85, Mau87]. acidic [GB88, KKP^{+}86]. Acidification [KKP^{+}86, GÅ89, LSSF89]. acidity [Reu80]. activated [ML84, VV83]. activity [Böl82, IUME85]. actual [NHTO86]. adaptation [Fle84]. Adapting [PS84]. adaptive [Rac87]. added [BvdD80]. Adenostoma [SR86]. adequacy [GGN83]. Advances [Fit84, Jør87a, Jør87b, Ush84a, Ush84b]. advection [DLMP87b]. advection-diffusion [DLMP87b]. adverse [MK86]. advisory [RWJ89]. aeration [TCL82]. aerobic [Vav82]. affected [SSJ89, Shi80]. afforested [Dye81]. African [CG83, Cou84, CMW84b, Gro86]. after [MTC85, OL88a]. Age [Wah81, FGS89, Ush81b, Vla88b]. age-class [Wah81]. age-structured [FGS89, Vla88b]. aged [LK87]. agent [KH87a]. Aggregated [ML84]. Aggregation [BCOG88, CO88, Gar88, IAL87, LSV83, Man82, AK88, LS86a, Luc84]. Agricultural [Jør81c, Mid84b, KMRV89, Maa88, MP83, Sko86a, Zuc89b]. Agriculture [Nih83a, Sko88, Sko86b, Sko86e, Sko86c]. agro [CWTH86]. agro-ecosystem [CWTH86]. agroecological [Rac87]. agroecosystems [KW84, VAH^{+}84]. AI [FPG89]. aid [Hus82, Rec85]. aided [Gal87]. air [DDM84, Saw83]. airborne [MRE87]. Akira [Nih81]. al [Jør81c, Nih85, BDVC88, GÅ89]. Alan [LU82a]. Alaskan [RJBL88]. Alberta [Ush85]. albipunctella [BMJ89]. Albrecht [Jol84, Mit87]. Alexander [Jør86b, SV89b]. algae [Sei80, Vyh87, WZE86]. algal [De 83, MK86, VH82]. algorithm [HHM89, NHTO86]. algorithmic [Log81]. Alison [Zuc88]. alkali [TN83]. all-wave [Var87a]. Allkin [Jør87c]. allocation [FGL82, FSTF86, GLLN82, MH86, MIK81, SHOP88]. alloregulated [Lom84]. allowing [BDVC88]. alluvial [PC86]. along [FGL82]. also [Lon87]. Alta [Ush85]. Alternative [Lom88, Bos83, Gla88, KS85, KBPR87, Mal85, Rag84]. alternatives [IHH81, Hal86a]. alterniflora [MHB84]. Alves [Nih83a]. America [Gra88b]. American [Mid84b]. americanum [BMJ89]. ammonia [CSG85b]. among [BBDK81, CSG85b, EEKM82, Gla88, Kar85, WSWP85]. Amsterdam [Ano82h, Baz86, Gro81, Gro83b, Jør81a, Jør86c, Lho87, Mau88, Mid84a, Mit87, Nih82, Nih84d, Nih84a, Str84a, Svi89, Zuc89b]. Amsterdam/Oxford/New [Jør86c, Lho87, Mid84a, Mit87, Nih84d]. Amur [EF82]. Analyses [HCR81, AM81, DJGR88, DK83b, VG82]. Analysis [Ano87h, Cas89, Gro84a, Har84, KA84, Lon86, OMA^{+}86, RS89, Ras87, Rec85, Rob89, SMR86, Ano82h, Atk87, Bai86, BGO84, Bax85, BM86, Bos81a, Bos81b, Bos86, BR82b, Bos89, Bra85, BRC84, CG83, Chr87, Cou84, DTL89, FGS89, FDC89, Fon84a, GOMC81, GRS88, GM86, Hal84, HIB85, Hig86a, HP86, Hip83, Hir85, HU86, Hus82, Joy85, Kaw86a, Kaw86b, KSW82, LS85, LPV87, LA84a, Log87, LMR84, LS86b, MSH85, MK86, Mah88, MRM81, MHB84, MRE87, Nih84a, OGC82, PM82, Rec84, RRW84, RMS89, SF83, SM82, SM81, Svi88c, Vid89, Wel88, WSWP85, Jør80a, Baz86]. analytical [Col83, Ons88, PG89, SBL88, Nih83a]. analyzing [KKP^{+}86]. Anchoa [PTS88]. anchovy [PTS88]. Andersson [Mau87]. angles [Cou84]. Animal [EP89, Wul80, FPG89, SSF^{+}88, Uch82, Wor87]. animal/habitat [SSF^{+}88]. animals [DH87b]. animations [HdJ89]. Announcement [Ano80a, Ano80b, Ano81a, Ano81b, Ano82a, Ano82b, Ano82c, Ano84a, Ano84b, Ano84c, Ano86a, Ano87a, Ano88a, Ano89a]. Announcements [Ano81c, Ano85a, Ano85b, Ano87b, Ano88b, Ano89b]. annual [GIG81, OA82, TD82, WC81]. annual-crop [GIG81]. anthrax [HF83]. anthropogenic [Koh89]. any [EM84]. any-deer [EM84]. Aphanizomenon [Var88b]. Apis [RM87]. Appalachian [RSS83]. apple [WH89]. Application [Bax85, BJ89, DABG84, Fon84a, KS89b, LPV87, MBM^{+}89, RU89, Smi80b, WMP^{+}87, AMB87, BGO84, BR82a, BS89, COS83, Fle84, FV84, HS84b, Jør81b, JG86, KYTS86, KKP^{+}86, KvdK88, NHTO86, PM82, Saw83, SSJ89, SF82, TH80, Var88a, Ver87, WSML89, CHK84, Svi88a, Zuc89b, Mid84a, Nih84d]. applications [BV86, CS85, Gro85, PH84, RI85, RC88, Sve85, Uch82, Log87]. Applied [Ano87h, Gro84a, HH86, Lon86, Ush84a, Cer89, God86a, Jør87a, Mah88, Meh84, RMS89]. Applying [Loe87a, DeA88]. approach [ÅA80, Aiz89, Arp83, Atk87, Bai86, BJ89, Bos83, CWTH86, DH87a, Fed84, Fon84b, Fon84c, GRS88, Gyd84, HCL83, LSSF89, LHSS86, LS86a, Lon88a, MIK81, OSW85, Pat86, RBK81, SSF^{+}88, SWPW85, SMR86, TI85, Tri87, Uri80, VAH^{+}84, Vid89, WRSZ87, WMP^{+}87, Str82a]. approaches [HPI85, MS87a, Som82]. approx [Som84]. approximately [Nih85, Ush85]. approximation [Luc88]. April [Ano81o, Ano82p, Ano83m, Ano85u, Ano88q, Ano89-29, Str84a]. aquaculture [CSG85a, CSG85b, CSG85c]. aquae [Var88b]. aquatic [AKL84, BS82, BS87, CH84, CHK84, DeA88, Fon84b, Fon84c, KS89a, KD82, KS89b, Rec85, SR84, Vol82, WO81, Gro83a, Str82a]. aquatica [RSG87]. Aquatische [Jol84]. arbitrariness [FVB81]. archipelagos [Aiz89]. Arctic [NNP84]. area [AW85]. arid [PC86, SWPW85]. ARMA [Gui86]. Aromatics [BGO84]. art [Ano83i, Baz86, Gro80]. arthropods [YW87]. Articulation [CS85]. Artificial [Ryk89, BH84, DNHP89, Loe87a, SSF^{+}88, TSOS86, VVK80]. artificially [Phi85]. artificially-drained [Phi85]. ascendency [Her89]. ASI [Leg89a, Wel88]. asio [Nor85, NBF88]. aspects [DTL89, DV86, HP86, LP83, OL88b]. aspen [RBK81]. Assessing [Sko88, Rag84]. Assessment [Gro86, BRC84, Hal88a, HM88, PU88, Phi85, YL87, Jør86b, Jør87f]. assimilation [Jør81c, URC88]. assist [Ter86b]. assisted [KBPR87]. Association [Jør87c, OMA^{+}86]. Aswan [Lho87]. Asymptotic [MRE87, HIB85]. Atlantic [SL84]. atmosphere [KE84]. Atmospheric [Gro84b]. attached [KGHM86]. attack [Bar87, SR89]. attempt [LS86a]. attraction [Mik86]. attributes [NO84]. August [Ano80j, Ano81n, Ano82r, Ano84r, Ano85v, Ano88z, Ano89w, Gro80]. Austin [Fit84]. Australia [HON86, PC86]. Austria [Ano87h, Gro84a]. Author [Ano80c, Ano81d, Ano81e, Ano81f, Ano82d, Ano82e, Ano82f, Ano82g, Ano83a, Ano83b, Ano83c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano85c, Ano85d, Ano85e, Ano85f, Ano86e, Ano86b, Ano86c, Ano86d, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88c, Ano88d, Ano88e, Ano88f, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g]. available [Gro84b, Lon87]. Averaging [LS86a]. aztecus [CG87, GG83].

B [God86b, Gro84b, Jør82c, Jør83, Leg84b, Mau87, Mej83, Pac88, Sko86a, Sko86d, Nih84d]. background [Bla83]. Bacon [Lon87]. bacterial [IUME85]. Baffins [Hal86a]. balance [Aok87, Cal88, Gro83c, RC88, Var88a]. balances [Mau82b]. Balaton [Ano87h]. Baltic [Vid89]. bands [PE81]. Bank [Bax85]. banksiana [MW80]. bark [DS85]. Barnes [Gro85]. Baruch [Jør80b]. Bary [Mic85]. based [AM87, AMB87, Bax85, Fet81, GOMC81, HP83, JB87, JKNC^{+}86, LSSF89, Lom88, MH86, MM87, PE81, QF85, QF88, SWPW85, SFT89, VAH^{+}84]. Basic [Jør81b]. Basin [Jen82, MC86, IA86, SMR86]. Basin-wide [MC86]. basis [Smi80a]. bass [DABG84]. Bay [GG83, PTS88]. Bean [Ano82h]. Bean-San [Ano82h]. Bechmark [Jør82a]. BEEPOP [DHRLE89]. bees [WHB83]. beetle [DS85, FCC81, FWS^{+}81, KWSM89, SH85, SH86, WGSC84]. before [OL88a]. behavior [AV81, BBC88, HIB85, KH88, KDC84, KS89a, NUK86, Ree83, Smi80b]. behaviour [Hal84, Woo88]. Belle [Jør80b]. belowground [BHD89a, BHD89b, MHB84]. Bennett [Jam87]. bent [HS81]. bent-over [HS81]. benthic [Sei80]. Bering [Pol85]. Berlin [BK86, Gra88b, Hof84, Jør80a, Jør86e, Leg89a, Nih85, Pal89, Sko85, Sko86d, Som84, Str87a, Svi88b, Ush81b, Ush81a, Ush85, Wel88, Wul86, Zuc88]. Berlin/Heidelberg [Sko85, Sko86d]. Berlin/Heidelberg/New [Hof84, Nih85, Str87a, Ush85, Wul86]. Berlin/New [Svi88b, Wel88]. Berryman [LU82a]. Bert [Gue83]. best [Hal83a, Hal83b, Hal85]. between [AR83, Ber85, BBD85, Day88, Dye81, EP88, Hir85, KP89, Lon88a, Mal85, OBPK89, Sko86d, SO88, Vyh87, WH86]. between-tree [KP89]. Beyer [Str82a]. BGS [HHS^{+}86]. BGS-II [HHS^{+}86]. BGS-III [HHS^{+}86]. Bias [SBL88]. biennials [KD83]. bioaccumulation [Ber85, BMOW82, RMS89]. biocenoses [SS82]. biochemical [Leg84b, VVK80]. biocontrol [KH87a]. bioeconomic [Cha86, GIG81, GGBH84, Tre89, Gat88]. Bioenergetic [DTL89]. bioenergetics [CSG85a, CSG85b, CSG85c]. biofilm [KGHM86]. biological [Ano82h, Gat83, Gra88b, HBB^{+}84, Leg84a, Log87, Nor87, SLG84, SR84, Vav82, WO81]. biology [Log87, Str82b, Str84b, Str87b, Ush85, Ush84a, ES85, Jør87a, Jør87e]. biomass [AM87, AMB87, Bax85, CE87, LMR84, Mah88, PS84, RW87]. biomass-based [Bax85]. Biomathematics [Som84, Svi88b, Ush85, Ush81b, BK86]. biomonitoring [RMS89]. Biophysical [Cle87, Kau87, WGSC84]. bird [OL88a, Smi83a, Smi83b]. Bisby [Jør87c]. bisinuatus [EGJ^{+}86]. Bisset [Jør86b]. Bistability [KB88].

Bivalvia [EGJ^{+}86]. Biwa [Aok87]. Black [Pod82]. Bleloch [Gra88a]. block [CL89]. bloom [KI86]. blue [BHD89a, BHD89b]. Board [Hal86a, Sko88, Ano80d, Ano80e, Ano81h, Ano81i, Ano81j, Ano82i, Ano82j, Ano82k, Ano83d, Ano83e, Ano83f, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano85h, Ano85i, Ano85j, Ano85k, Ano86f, Ano86g, Ano86h, Ano86i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q]. Boca [Ush84d]. BOD [CP88]. bodies [KSBM84, MK86]. body [Vol82]. bog [LA84a, LA84b]. Bolin [Gue83, Leg84b]. Book [Gro84a]. bootstrap [Hal89b]. boreal [Bon89]. borer [MSR87]. Bose [YW87]. Bosserman [Str84a]. bottomland [PMK85]. Boulder [SV89b]. bound [KB84]. boundary [DLMP87d, Sul88]. bowhead [RJBL88]. box [MF86]. Brachidontes [EGJ^{+}86]. brackish [Böl82]. Brakensiek [Mid84b]. breeding [Smi83b]. Brian [Jør86b, Pal89]. Brisbane [Jør80c]. Britain [Ano87i, BK86, Gat88, Gra89, Gro84b, Hal86a, Jac87, Jør86d, Jør87d, Leg84b, Log87, Lon86, Mid84b, Nih84b, SB85, Sko86a, Sko88, Svi88a, Svi88c, Zuc89a, RU89]. British [Smi83b]. broad [DJGR88]. broad-scale [DJGR88]. bromate [KEK85]. brown [CG87, GG83]. Bruk [Lij87]. Bruno [Mau88]. Brunswick [MW80]. budget [KvdK88, Var89b, Wul80, YL87]. budgets [TSOS86]. budworm [Cas82]. build [Vol82]. building [Vol85, Gra88a]. Bulletin [Nih83a]. Bulletins [Mau87]. burning [DK83a]. bush [FP86]. Butler [SB85]. butterfly [Loe89]. Butterworth [Sko86c]. Butterworths [ES85].

C [Gro81, Gro83b, Jør81c, Jør87b, Mau88, Mej83, Mid84b, Nih85, Saw83, SB85, Str87a, Ush85]. C.J [Jør86c]. Ca [BDVC88, GÅ89]. calcite [KBPR87]. CalComp [HHM89]. Calculating [Bro81, Bos83]. calculation [Leo83]. Calculations [SL84, Wul80]. calibration [RSKT88]. California [BH84, Kam81]. Call [Ano85g]. Cambridge [Gra89, Jør82c]. Canada [Nih84d, BGTH82, Joy85, MW80, Ush85]. Canadian [BJ89, EBSW84]. canopies [OBPK89]. canopy [Jia88, KP89, RC88]. canvasback [AT83]. capacity [HH86]. capensis [FVV86]. capture [MB88]. capture-recapture [MB88]. carbon [BHD89a, BHD89b, BS89, CE87, GE84, Har84, HP83, JBL^{+}89, Jør86d, KvdK88, LA84a, MH86, MRE87, RA85, SKMD80, SM81, Sum85, YL87, Bos81a, Gue83]. Carey [Wel88]. Carlo [Gro84a, OGC82, RMS89]. Carolina [Jør80b, BWKR87, PMK85]. carrying [HH86]. cartographical [SMR86]. case [HS84a, Hip83, HH86, HL85, IY80, LR82, Lho87, Wul86]. Catastrophe [Loe89, Loe85, OL88b]. Catastrophes [Cas82]. catch [SRM83]. catch-effort [SRM83]. catchment [HL85, Var88a, Mau87]. catchments [Dye81]. caterpillar [RSG87]. Cation [GB88]. cattle [Sen89]. causal [MFG82, Umb89]. cause [KSB^{+}84]. cause-effect [KSB^{+}84]. caused [Kin87]. CE [WC85]. CE-QUAL-R1 [WC85]. cell [UO84]. Celtic [Par86]. censuses [SRM83]. central [PC86]. Centre [Jør81c]. Cereal [SH86, SH85]. certain [AK88]. certainty [IDS81]. Cervus [SFS84]. chain [GMM88]. chains [Par82]. Challenges [Saw83]. changes [AW85, Ess89, GÅ89, Jør88, RM87, Shi80, SD83, WH87a, YL87]. Changing [KST^{+}88, Ano85n]. chaparral [SR86]. Chapman [Nih83b, Pac88]. Characteristic [Tay88]. characteristics [HCR81, Jør82c, Jør83, Kam81, LF85]. Characterizing [Cal88]. charge [Smi80b]. charge-discharge [Smi80b]. Chelsea [Jac87]. chemical [Bos83, Gro83a, KB81, LAF81, Tho84]. chemicals [BRC84, Gyd84, MBM^{+}89, MC86, PM89]. chemistry [Gro83a]. chemostat [BV81, KB88]. Chichester [Ano87i, Gat88, Gro84b, Hal86a, Jac87, Jam87, Jør80c, Jør86d, Jør87d, Leg84b, Log87, SB85, Sko88, Svi88a, Svi88c, Zuc89a]. chimney [HS81]. chlorinated [FD86]. chlorophyll [OFMS86, Par86, Smi80a]. Chthamalus [EGJ^{+}86]. circulation [BCJ82, KT89, Lei86, WH86]. Cirripedia [EGJ^{+}86]. Clark [Gat88, Jør86b]. class [LU82b, Wah81]. Classical [Chr87]. classification [Loe87b]. Cliffs [Jør87f]. climate [Gui86, SF82, Zuc89a, Koh89]. climatic [DW84, JBL^{+}89]. clonal [KB84]. Closed [Ker83, LS85]. Closed-form [Ker83]. Cloth [Hof84, Jør86e, Jør87d, Nih85, Nih81]. cluster [BM86]. clustering [FDC89]. CO [SV89b, Nih84a, Sko86c, Dah85]. Coaker [Jør87a, Ush84a]. Coastal [Jac87, Wul86, Zuc88, IUME85, Jør80a, KI86, KMRV89, Leg89a, LJZ82, SCD85, Jac87]. codling [GKS87, SG85]. coefficient [AEO83, Hus84]. coefficients [Bac85, Lom86]. coexistence [Lom85]. cognitive [OMA^{+}86]. cold [Fet81]. Coleoptera [DS85, WGSC84, Nut89]. Colin [Gat88]. collecting [Vol82]. Colloq [Mau88]. Colloquium [Gro83b, Jac87, Jør86c]. colonial [Kar85]. colonization [Sei80, Ter86a]. colony [MR86, RM87]. column [CM87]. Combinatorial [LS87]. combined [Cha88, Bos81b]. comment [Bar89, Wal85]. commercial [GMM88]. Commission [Hal86a]. Committee [Sko88]. communities [Dah85, Day89, GRS88, HSP88, Mik86, PMK85]. community [Day88, Fle87, Lai87, MM87, Ras87, Sei80, Wul80]. Company [Gro81]. comparative [AK88]. compared [VKDS86]. Comparing [Lom84, KS85, WG89, Hal83b]. Comparison [FNK81, HK84, JKNM82, AR83, DDS85, GOMC81, HPI85, Har84, Kar85, Lon88a, Som82, UO84, Jør81a]. compartmental [Hig86b, Hip83, Nem87]. Competing [SRM83, HK84, Lom85]. Competition [KB84, Abd80, BBDK81, CO88, CSG85c, DS85, Jen87, LW88, Lom86, Mal85, Sei80, WC81]. Competitive [Kar85, AEO83, Bar82, MDS84]. complete [LU82b]. complex [AEO83, BC82, Bel84, GB81, JKNM82, KR82, KSBM84, NHTO86, WH87b]. complexity [Lim85, SD83, YL87]. Components [Jør86e, EP89, Jen89, LS86b]. composition [Sen89]. comprehensive [GM86, Rec84]. computation [Gro83c]. Computer [Ano87i, DAA82, Ano89r, BMJ89, Bon89, Gal87, GRS88, GE84, HdJ89, Hal89a, Hal89b, KB81, KH87a, Nor85, NBF88, Ter86b, WZE86]. computer-aided [Gal87]. computing [Lij87, Var87b]. concentration [KEK85, SD83]. concentrations [SA86, TSJ87]. Concept [KMW84, AEO83, HH86]. concepts [DNHP89, SI84, Sko86a]. Conceptual [PU88, ZMRB84]. Conceptually [Las83, Ano83h]. concerns [Hir85]. concurrent [Bar80]. conditions [AR89, CTJ81, DLMP87d, DLMP87e, IA86, KHMBD86, Loe87c, Sul88, VVK80]. Conference [Gro80, Nih84a, Ano83i, Ano89r, Ush85, Vid89]. Confidence [PE81, SO88]. Confirmation [RC83]. congruence [Log81]. congruous [Mik86]. coniferous [KA84]. conjunctive [Tya88]. Connectance [Mar87]. connective [Nih83a]. connective-dispersive [Nih83a]. consequences [Rag84]. conservation [Gra88a, Pac88, Sko88, Gra89]. considerations [Gro88, Zuc83]. considering [PLRV88]. consistency [CO88, KvdK88]. constancy [Bar89]. constant [Hig86b, Par82]. constitutes [Hal88b]. constraints [Lai87, OGC82]. constructing [SO88]. construction [Loe87b]. consumer [DH87a]. consumers [Whi84a, Whi84b, Whi84c, Whi84c]. consumption [Bla83, VV83]. containing [LS86b]. Contaminant [Jen86]. contamination [HGG^{+}82]. Contents [Ano88g, Ano88h, Ano88i, Ano88j, Ano89h, Ano89i, Ano89j, Ano89l, Ano89k]. context [Vol86]. continental [HPKA80, Gro84c]. Continuous [UO84, OSW85, SWPW85, Vyh87]. continuous-time [OSW85, SWPW85]. continuum [Ree83]. Contribution [Fle84]. contributors [Ano86j]. Control [Smi87, TCL82, BS87, Cas82, Fle87, Gro86, IDS81, JUGH83, KOSS82, LP83, Lon86, OF81, Rec85, SSS81, Str83, TI85, Wal85, Wel88, Wil82, Jør80c]. controlled [JBL^{+}89]. conversion [SAT86]. Conway [Lon86]. Cook [Leg84b]. Cooke [Gat83]. cooling [Jen82]. Copenhagen [Gro80, Jør81a, Zuc89b]. Copenhagen/Elsevier [Zuc89b]. coregonid [DTL89]. corn [WS80]. corngro [KSW82]. cornstalk [MSR87]. Correlation [RPM86, BM86, Gar84]. Corrigendum [Ano81g]. Cost [MK86]. Cotton [GPD^{+}84]. Council [Hal86a, Sko88]. Councils [Mau87]. countries [Nih85]. country [BT82]. coupled [CS89]. coupling [SKMD80]. course [DKR84]. covariation [Dye81]. cover [AW85, HK84, Wul86]. coypu [RU89]. Crank [Mej83]. CRC [Ush84d]. credibility [Pal89]. Creek [HON86, Var88a]. cristata [FV84]. criteria [Hal88b, KST^{+}88]. criterion [WG89]. critical [EP89, Loe89, TH88, TH89]. critique [RA85]. crop [AEO83, GIG81, HM88, Kin87, Maa88, RS89, SH85]. crops [Jør81c]. crown [KP89]. Cryptolestes [KWSM89]. Crystophora [FV84]. Ctenopharyngodon [EF82]. CTM [OSW85]. cubs [GHV86]. Cuff [Wul86]. Cullen [Log87]. cultivation [Vyh87]. current [Cle87]. cusp [OL88b]. cutting [KYTS86]. Cybernetic [Str83, MS87a]. cycle [Cou81, DK83a, Gue83, Jør86d, LA84a, LA84b, MW80, RSS83]. cycles [GE84, Leg84b]. cyclicity [Lom88]. cycling [Bos89, CAJ^{+}87, CR80, HF80, Har84, HCR81, PH84, Pat85, Ush84d]. Cydia [GKS87, SG85]. Czechoslovakia [Str82b, Str84b, Str87b].

D [Ano87i, Gat83, God86b, Jør86b, Mej83, Mid84b, SB85, Sko85, Str87a, Tre89, Ush84b]. Dacus [CF88]. Dagan [Sko86d]. daily [FRCC89, OA82, Var88a]. Dale [Lho87]. dam [Lho87]. Dame [Jør80b]. Daphnia [Wul80]. data [BM86, CS89, DW84, FM88, Hal88a, HP83, Lom86, Lom88, LS86b, MB88, OA82, RSKT88, RMS89, Smi80a, SFO^{+}86, TH80, WC88, WHB83, WC85, YL87]. Databases [Jør87c]. David [Zuc88]. DC [Hal86a, Sko88]. DeAngelis [Str87a]. December [Ano80k, Ano82o, Ano83p, Ano84u, Ano85s, Ano86m, Ano87v, Ano88r, Ano89t]. deciduous [JBL^{+}89]. deciduous-type [JBL^{+}89]. decision [Ver87]. decisionmaking [Zuc88]. decisions [JUGH83, Ano87i]. decisive [Gro88]. decline [Gro88]. decomposition [FFMB82, GB88, PE81, Ush84d, Cas89]. decreases [Lom88]. dedicated [Hal89a]. deeply [CS89]. deer [EM84, SFS84]. Definition [Hus84, KvdK88]. defoliation [CMW84b]. Degradation [SFP^{+}89, DSSMZG88]. degradative [DSSM87]. Delaware [AM81]. delay [Agn82, GPD^{+}84]. delays [Pat86]. Delhi [Bis88]. demand [RRW84]. demise [Nor87]. Democratic [Jol84]. demographic [CS89]. Demography [Car82]. demonstration [BDVC88]. Dendroctonus [FCC81, FWS^{+}81, WGSC84]. denitrification [MB84]. Denmark [Gro80]. Density [Cer89, AK88, Bou85, Bur86, CP88]. density-dependence [Bou85]. density-dependent [AK88]. Department [Nih83a]. departure [Day89]. dependence [Ale82, Bou85, Bur86]. dependent [AK88, BBC88, HPKA80]. depletion [GB88]. deposition [Cou81, KKP^{+}86, KR86, PC86]. depth [SWH84, SV89a]. derived [Bos89]. deriving [FM88]. describing [KYTS86, RJ82]. Description [BGO84, TD82]. desert [AR89, Fet81, Loe85]. design [Ade85, BN87, KS85, VAH^{+}84]. designing [RWJ89]. desirable [COS83]. detecting [Day89]. Determination [TH89, SWH84]. deterministic [HF83, SMN85]. detritivorous [Car84]. developing [GG89]. Development [COS83, DLMP87a, HS84b, LS85, RJ82, RMS89, Bla83, Car84, CSG85a, DJGR88, GL87, GPD^{+}84, JUGH83, KKP^{+}86, KSB^{+}84, Lai87, LK87, MRM81, ON85, PU88, Par86, Rub83, TD82, WGSC84, Wal87, dC81, MP83, Smi83a]. Developments [Baz86, Lho87, Mid84a, Mit87, Nih84d, Str84a, Svi89, Zuc89b]. Dfl [Baz86, Fit84, God86a, God86b, Gro81, Gro83b, Jør86b, Lho87, Mau88, Mit87, Nih84d, Str84a, Svi89, Tre89]. Dfl.180 [Jør86c]. Dfl.190 [Zuc89b]. diagrams [HHM89]. dialectics [Pal89]. die-off [GS86]. dieback [Bos86]. diel [HP83]. difference [Bre89, Log81]. different [Cas88, Hal83b, SD83]. Differential [Par82, Log81, Urc84]. difficulties [DeA88]. diffusion [BKU84, DLMP87b, DLMP87d, DLMP87e, Gro84b, RU89, SH86, Sul88, Mej83, Nih81]. dimensional [DLMP87d, DLMP87e, Nih83a, VKDS86, WMP^{+}87, Mau88]. dioxide [RA85, TSJ87]. direct [KP89]. discharge [Smi80b]. discrete [Agn82, Gre83, TD82]. disease [And83, DAA82, KH87a]. disintegration [BBD85]. Dispersal [WC81, HBB^{+}84, McC86, SF82]. Dispersion [Chy84, GHV86, Gro84b, HSA85, KB81, Par82]. dispersive [Nih83a]. display [HdJ89]. disposal [BJ89, Hal86a]. dissolved [CSG85b, OF81, TSOS86]. disstria [RSG87]. distance [GHV86]. distibution [WRSZ87]. Distributed [Hal86a, GPD^{+}84, Jac87, Sko88]. distribution [Bre89, DH87b, HPKA80, TI86, TSJ87, Tay88, TH88]. distributions [Wah81]. Disturbance [Jør86e, CEMF84, DWPO85, Gat87, GP85, KB84]. D’Itri [Jac87]. diurnal [KHMBD86, Kaw86a]. diversion [PMK85]. diversity [BK86, GM86, KB84, Nor87]. DM [BK86, Gra88b, Jør86e, Nih81, Nih85, Som84]. DM138 [Str87a]. DM139/US$51.90. [Sko86d]. DM168.00 [Svi88b]. DM178.00 [Wel88]. DM198.00 [Leg89a]. DM48.00 [Zuc88]. DM49.80 [Wul86]. DM58.00 [Jol84]. DM62.00 [Ush85]. DM62.00/ [Ush85]. DM98.00 [Pal89]. DM98.00/US$42.30 [Hof84]. do [TH88, CP88]. Documentation [Jør81c]. domain [DK83b]. domains [Ker83]. domestic [Hal86a]. dominance [Var88b]. dominated [CHK84]. Dordrecht [Jør86b]. double [Ful87]. Dowden [Jør82a]. drained [Phi85]. Drake [Gra88b]. Dramatic [KSB^{+}84]. drinking [FD86, MK86]. driven [GS86]. drone [RM87]. dry [Cou81, Var87a]. dryland [Tre89]. duck [JSC87]. ducks [AT83]. due [FP86, KBPR87, RFJ87]. during [OL88a]. dying [KSB^{+}84]. Dynamic [Cha88, SCD85, GS86, Her89, Hip83, JB87, Jør86f, Lev88, RS88, TH80, LA84b]. dynamical [Woo88]. Dynamics [Bos86, SW84, TA86, Ano87h, AFSK85, BHD89a, BHD89b, BM89, BBDK81, BK86, BGTH82, BMJ89, BS89, BWC83, Bur86, Car82, CWTH86, CM87, CF88, DHRLE89, EP89, EM84, FCC81, Fit84, FV84, FSK82, Gat83, Her88a, IY80, JBL^{+}89, KWSM89, KI86, LS87, Lon87, MSR87, Man82, MFG82, Nor85, NBF88, Nut89, Ons88, Pat88, PTS88, RJ82, RFCC89, SH85, SH86, SG85, SFS84, Sto83, Sum85, TI86, Umb89, Vla88a, Vol85, Vol86, Wol83, WH87b, WH89, YW87, RSG87, Gat87, Hof84, Mau88].

E4CHEM [MBM^{+}89]. each [Kau87]. East [Jac87, CMW84b, Cou84, IY80, KI86]. east-African [CMW84b]. eastern [Pol85, EBSW84]. eating [Rub83]. echinata [Kar81]. Ecohydrodynamics [Gro83b]. Ecol. [Ano83h]. ecologic [LAF81]. Ecological [Bla83, BG87, DWPO85, Gra88b, Gro80, Jør80d, Kre83, Leg89a, Mau87, Nih81, Sko85, Sko86e, Str87b, Wel88, WSWP85, Zuc89b, Ano80g, Ano83i, Ano88i, Ano88j, Aok88, Baz86, BR82a, BM86, Bos89, BV86, Bre89, Bro81, COS83, CMM^{+}86, CD87, DKR84, Ewe89, Fah88, Fle82, Gar84, GP85, Gro85, HHS^{+}86, Hal89a, HU86, Hul85, JB87, JG86, KK84, KSBM84, Loe87a, Loe87b, LA84b, LS86a, Meh84, Mid84a, MC86, MRBU89, Nih84d, NHTO86, ODP^{+}89, OSJ^{+}86, OL88b, PH84, Pea87, Pod82, Rec84, Ros86, SW85, SFT89, Str84a, SFO^{+}86, TA86, Tho84, Tri87, Ush84b, Ver87, Jør86e, Ano89l, Ano88g, Ano88h, Ano89h, Ano89i, Ano89j, Ano89k, Jør82c, Jør83]. ecologically [VAH^{+}84]. ecologically-based [VAH^{+}84]. Ecology [Jør82a, Nih83b, Cas88, Che89, Fle84, Gal87, Gat87, Gro85, Hal88a, Her88b, Leo83, Loe89, Luc84, Nih84b, Pat83, Ryk89, Sil82, Som82, Svi88b, Ula88, Ush81b, VBV87, Jør81d, Jør82a, Jør87b, Gra88b, Jør87d]. ecology-oriented [Som82]. Ecology/9 [Jør82a]. econometric [Joy85]. Economic [BV86, CMM^{+}86, Hue87, KOSS82, Lai87, SW85, TI85, Umb89]. Economic-ecological [BV86, CMM^{+}86]. Economics [Nor87, Cle87, CD87, GL87, Kau87, Pea87, VBV87]. Ecosystem [BR82b, Ale82, AKL84, Bai86, Bax85, Bel84, Cal88, CWTH86, CAJ^{+}87, CR80, DK83a, DK83b, Gar81, Gar88, GOMC81, GE84, JT80, Jør80a, KvdK88, KS89a, KD82, KBSM88, LS87, LS85, LJZ82, LA84a, LA84b, LC84, MW80, MS87a, OGC82, OFMS86, Pat85, Rec85, RFJ87, RC88, SB85, SAT86, SM81, Sum89, Sve85, SKV84, TCL82, Wol83, WZE86, dC84, Vol86, Svi88a]. Ecosystems [AK85, Mit87, AK88, Atk87, BS82, BS84, BS87, Bis88, Dah85, FVB81, GB81, Gen84, GM86, HPI85, Her89, Hig86a, HP86, Hig86b, IAL87, Jør86e, Ker83, KR82, KMW84, Lev88, LU82b, LSV83, Phi89, Rec85, Sko85, Sko86a, Str82a, Str83, SU87, Umb89, VS84, Vol85, WO81, Gro84c, Leg89a, SV89b]. ecotoxicity [MBM^{+}89]. edited [Nih82]. Edition [Gro84b]. Editor [BK86, Gra89, Gro80, Gro83b, Gue83, Jør80b, Jør80c, Jør81a, Jør81d, Jør86c, Jør86d, Jør87a, Jør87b, Jør87e, Leg89a, Lon86, Lon87, Mid84a, Nih84a, Pac88, Saw83, Svi88a, Ush84a]. Editorial [DP87, GR82, Jør82d, Jør82b, Jør84a, Jør86a, Jør80d, Ano80d, Ano80e, Ano81h, Ano81i, Ano81j, Ano82i, Ano82j, Ano82k, Ano83d, Ano83e, Ano83f, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano85h, Ano85i, Ano85j, Ano85k, Ano86f, Ano86g, Ano86h, Ano86i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q]. Editors [Baz86, Fit84, Gat83, Gat87, Gra88b, Gro81, Gro83c, Gro85, Hof84, Jac87, Jør82a, Jør86e, Jør86b, Jør87c, Leg84b, Leg84a, Mau87, Mau88, Mid84b, Nih84d, Sko85, Sko86a, Sko86b, Sko86d, Som84, Str84a, Svi88b, SV89b, Svi89, Tre89, Ush81b, Ush84b, Ush84c, Ush85, Wel88, Wul86, Zuc88, Zuc89b]. Edmonton [Ush85]. education [Hal89a]. Effect [GFI84, KGHM86, PTS88, Sen88, Sul88, Ågr81, AMO83, AR83, BBDK81, Bos81b, HBB^{+}84, KSB^{+}84, KP89, OBPK89, SHOP88]. effectiveness [CS85]. Effects [Bur86, DH87b, DH87a, EF82, Fet81, HCL83, KH87b, Sen89, Arp83, BRC84, BCJ82, Cas89, DK83a, GÅ89, GKS87, HS87, HP86, Jør82c, Jør83, KA84, MW80, MK86, PLRV88, RFJ87, RFCC89, RW87, Sch85, SF82, Sum89, SR84, WC81, WS80, CSG85b, CSG85c, SB85]. efficacy [SSJ89]. efficiencies [SAT86]. efficiency [Odu83, Sil82, URC88]. effort [GLLN82, SRM83]. Efraim [Jør81d]. egg [BWC83]. Einstein [YW87]. Elasmopalpus [MSR87]. electronic [Ano87i]. elevated [TSJ87]. elevation [KP89]. elimination [Ågr81]. Ellis [Gro84b, Log87]. Elsevier [Ano82h, Baz86, Gro81, Gro83b, Jør81a, Jør86c, Lho87, Mau88, Mid84a, Mit87, Nih82, Nih84d, Str84a, Svi89, Zuc89b]. emission [TI85]. emissions [HSA85]. emphasizing [Gro83a]. empirical [Zuc83]. emsy [KBSM88]. encroachment [FP86]. Energy [Her89, Pat85, Zuc83, Ano80g, Chr87, Jør81a, KE84, Lim85, Lon88a, Mau82b, Nih85, SAT86, Uch82, Vol82, Wul80, Sko86b, Str84a]. engineering [Uri80]. Engineers [Mid84b]. England [Gro83a, HBB^{+}84]. Englewood [Jør87f]. enhancement [BH84]. enigmas [Pat83]. enrichment [Dah85, IY80, OL88a]. entities [Cal88]. entrainment [PTS88, Rag84, Sum89]. Entropy [Aok87, Aok88, Aok89, Mau82b]. Environ [Hip83, Hir85, PM82]. environment [Ano85n, Böl82, Col83, HdJ89, Hue87, Kir89, LP84, MIK81, OMA^{+}86, Pod82, TH89, Wul80]. Environmental [Baz86, Lho87, Mit87, Nih84a, Str84a, Svi89, TI85, Zuc89b, Ano89r, Bec81, BRC84, BN87, DKR84, DW88, ES85, FM88, GG89, HCL83, HIB85, Jør86b, JG86, LS86b, MBM^{+}89, Mej83, Mid84a, Nih82, Nih84d, PM89, RC89, RBK81, Vid89, Gro84a, Nih84a, Zuc88]. Environmental-economic [TI85]. environments [Man82, SFO^{+}86]. environs [LA84a, PM82]. enzootic [DAA82]. EPI [DAA82]. EPI-enzootic [DAA82]. epidemiological [GHV86]. epizootic [HF83]. equation [Ful87, GMR88, Moo89, Nih83a]. equations [Fon84b, Ker83, Wah81]. equilibria [Gro83a]. equilibrium [Bos83, Bri83, Fon84b, Fon84c, KR82, LAF81, HCL83]. Equivalence [Hir85, Hul89, Hul85]. equivalent [AEO83]. Erie [Jen82]. Erlangian [MB88]. erosion [God86b, KR86, PC86, RW87]. Errata [Ano80f, Ano83g, Ano85l, Ano87o]. Erratum [Ano83h, Ano84p]. Error [Hal84, BCOG88, GOMC81, Loe87b, OGC82, WG89]. Errors [Loe87b]. Erxleben [FV84]. estimated [NO84]. estimates [Ågr81, LW82, SBL88]. Estimating [LHSS86, SO88, Uri80, WC88, CP88, IA86, Smi80a]. Estimation [GHV86, Pal88, VH82, BK84, Cer89, JJKNM81, Jør84d, Loe88, MBM^{+}89, NHTO86, ZBBLCSTL86]. estuaries [Wul86]. Estuarine [Jør80b, Wul86, Zuc88, DK83b, KK82, Mau88, SFT89, SKMD80, SM81, Sum89, Jør80b]. Estuary [HBB^{+}84, KvdK88, OF81]. Europe [KKP^{+}86]. European [BS89, WH89]. eutrophic [MF86]. eutrophicated [KI86]. Eutrophication [BDLP87, DLMP87a, Gro86, JJKNM81, JKNJ86, JKNC^{+}86, KOSS82, MGN86, Miy86, RPM86, TD82, VS84, VT87]. evaluate [KH87a, TCS89]. evaluated [SS87]. evaluating [CMLB85, WG89]. Evaluation [GFI84, KHF88, Leo83, WC85, AEO83, BRC84, DWPO85, GM86, KBPR87, Loe87b, MB84, OSJ^{+}86, Phi89, Pac88, Wul86]. evaporation [Var87a]. even [LK87]. even-aged [LK87]. event [FGS89, Gal87, LPV87]. events [Smi87]. Evolutionary [BK86]. Examination [JKNJ86]. example [BH84, DDM84, FVB81, GPD^{+}84, JCN85, Loe89, OFMS86, Rub83, ZBBLCSTL86]. examples [Wil82]. exchange [Ano83h, GÅ89, Jør81a, Las83, Lon88a, RC88, SKMD80]. exergy [Her89, Jør88]. existence [SS82]. experiences [GG89]. experiment [JG88, KI86]. Experimental [Car84, Jør88, Sci86, SS87, Smi80b]. experiments [Cas89, KS85, RPM86]. expert [RWJ89, Rit89, Ryk89]. explain [KBPR87]. explaining [Ano87h, MFG82]. exploit [Gla88]. exploitation [Agn82, SW85, SS82, SFS84, SW84, Wal85, Wal87]. exploitative [CO88, Jen87]. exploited [Lev88]. Exploiting [SF83]. Exposure [MBM^{+}89, BRC84, HCL83, PM89]. exposures [Sch85]. Extended [Hig86a]. Extinction [FP86, HdL84, Man82]. Extreme [FGS89]. Extremes [Bla83].

F [ES85, Fit84, God86a, Gro84b, Jør80b, Jør87c, Jør87d, Mau87, Sko85, Zuc89b]. facilitation [Gla88]. facilities [Nih85]. factor [Gro88, Phi89]. factorial [KS85]. factors [Bos81b, RBK81, Umb89]. fall [SF82]. farm [Tre89]. fasciculatum [SR86]. Fast [WSML89, KDC84]. fate [Hal83a, Hal83b, Hal86b, MBM^{+}89, MC86, PLRV88, PM89, Tho84]. Fates [BGO84]. feasibility [HS84a]. February [Ano81r, Ano82n, Ano83q, Ano84t, Ano87q, Ano88y, Ano89v]. Federal [Pal89, Zuc88]. Fedra [Gro84a]. feedback [Str87a]. Feeding [Vol81, CSG85c, Sjö80]. Felsot [God86a]. Ferguson [Gro83c]. ferrugineus [KWSM89]. fertilization [BHD89a]. FF [Gro83c]. FF50.00 [Leg84a]. Fibonacci [Var89a]. Field [CEMF84, FCC81, Kam81, KHMBD86, Lom86, Lom88, PC86, RPM86, VKDS86, WSWP85]. fig [Nih83b]. figs [Gro84a, Jør80a, Jør80b, Jør82c, Jør83, Nih81, Jør81c]. figures [Jør86e, Nih84b]. filtration [Ano83h, Las83]. finite [Bre89, UO84, WRSZ87]. finite-difference [Bre89]. fir [KSB^{+}84]. Fire [CHC89, GGN83, KA84, RWJ89, Ush81a]. fire-spread [GGN83]. fires [Gre83, Mal85]. First [Jac87]. Fischer [Jol84]. Fish [CSG85a, CSG85b, CSG85c, IY80, BDVC88, DTL89, DW88, FGL82, GLLN82, Jen86, MTC85, Pol85, Rag84, Ree83, SKV84]. fisheries [Bax85, Gat88, GIG81, LJZ82, LPV87, Svi88c]. Fishery [CG87, BH84, Cha86, Cha88, IHH81, KGG87, Man82]. Fit [Aiz89, Cos89, Hal85]. Fitting [LW88, WHB83]. fjord [NNP84]. FL [Gat87, Jør87a, Jør87c, Jør87e, Lon87, Ste88, Ush84d]. flat [PC86]. flatwoods [GE84]. FLEX [Tri87]. FLEX-REFLEX [Tri87]. flood [Var89b]. flooding [MTC85]. floodplain [PMK85]. Florida [EF82]. flos [Var88b]. flos-aquae [Var88b]. flow [Bos89, Bra85, HF80, Hig86a, Hir85, Lim85, Miy86, SM81, Uch82, Vol81, Vol82]. flows [AM81, SAT86, SU87]. Flug [Baz86]. fluoride [MA82]. flux [Var87a, Var87b]. fluxes [HON86]. fly [Car82, CF88, Wel88]. FOAM [BGO84]. following [KYTS86]. food [CSG85b, DeA88, JG88, Leg89b, Lim85, Par82]. Forage [RBK81]. foraging [Sen89]. Ford [Ush84b]. Forecasting [GMM88, KA84]. forest [AM87, AMB87, Bos86, BS89, BWKR87, DJGR88, DW84, EBSW84, GS86, Gro88, HK84, Jia88, Jør81a, KYTS86, KKP^{+}86, Kaw86a, Kaw86b, KE84, KP89, Mau87, Moo89, MA82, OBPK89, PMK85, RSS83, Rob89, RC88, WF88, RSG87]. forested [HL85]. forestland [AW85]. forestry [SFP^{+}89]. forests [Bon89, JBL^{+}89, KA84, KSB^{+}84, RSS83]. Foreword [GH87, RG85, RG88]. forgone [Rag84]. Form [Vol85, Ker83, Uch82].

Form-building [Vol85]. formation [DH87b]. formations [Fit84]. formulating [Ade85]. formulation [GP85, Lev88, Pat84, Str83]. formulations [KS85]. Foundations [Pea87]. four [Bri83, Nih85]. Fourth [Ano83i, Str87b]. fox [GHV86]. fraction [MRE87]. framework [Fed84, Wil82, WMP^{+}87]. France [Leg84a, Sko86c]. Francesco [Ush81b]. Francisco [Jør81d, Ush84a]. Frank [Jac87]. Free [Vol82]. Freedman [Ush85]. Frequency [DK83b, Ker83]. Frequency-domain [DK83b]. freshwater [CS85, Gro84c, IHH81, Leg89a, VS84, Mit87]. frontalis [FCC81, FWS^{+}81, WGSC84]. fruit [Car82, CF88, Wel88]. fuel [CHC89]. fuels [Gre83]. function [EF83, Ras87, ZMRB84]. Functional [Bah82, IUME85, GMR88, QF88]. functionalizing [RBK81]. functioning [Bel84]. functions [AM81, Luc88]. fungus [Mic85]. fur-seal [FSK82]. Further [HP86]. future [Ess89, Zuc83]. futures [Sko86b]. fuzzy [BR82b, Rob89].

G [BK86, Col83, God86a, Gro85, Jam87, Jør86d, Lon86, Mej83, SB85, Sko86a, Sko86d, Ste88, Svi88b]. G.D.R. [KK84]. Gårdsjön [Mau87]. Galveston [GG83]. game [SSS81, VG82]. gaming [RSS83]. gap [Moo89]. gas [Nih85, RC88]. gastropod [FDC89]. Gaylord [Baz86]. Gdańsk [Vid89]. General [Uch82, Abd80, FVV86, GIG81, Gro85, GPD^{+}84, LW88, LU82a, MSH85, PG89, RC88]. generality [JKNJ86]. generalizations [CL89]. Generalized [Vav82, KS89a, MS87a, SM82]. Generic [BS89]. genetic [BK86, LP83]. Genuchten [Nih83a]. Geographical [GS86, Ale82]. geography [Jam87, Col83]. George [Jør81b]. Georges [Bax85]. German [Jol84, KSB^{+}84]. Germany [Pal89, Zuc88]. Gilad [Jør86b]. Giorgini [Zuc89b]. Giorgio [Lho87]. gives [Hue87]. Global [Ess89, KS89a, Sko86b, Ano85n, Jør86d, LS85, MRE87, Var87b, YL87]. Gnauck [Jol84, Mit87]. Godron [Jør86e]. Goh [Ano82h]. golden [Ush81b]. Goldschmid [Sko86d]. good [Hal88b]. goodness [Cos89, Hal85]. Gossypium [GPD^{+}84]. Gotland [LJZ82]. gradient [DKR84, Ush81a]. Graf [Gro81]. grain [KWSM89, Lon88b]. grama [BHD89a, BHD89b]. graminoid [CMW84b]. graminoids [CMW84a, Cou84]. Grant [Svi88c]. graph [Rob89]. graphical [Hus82, KKG82, WSML89]. graphics [HdJ89, Hal89a]. grassland [Cou81, HPI85, PS84, SV89a, SAT86, Whi84a, Whi84b, Whi84c]. Grasslands [Fit84, LHSS86]. gray [RJBL88]. grazing [BHD89a, BHD89b, Cou84, Fet81, HP83, Mau82b, RJ88]. Great [Ano87i, BK86, Gat88, Gra89, Gro84b, Hal86a, Jac87, Jør86d, Jør87d, Leg84b, Log87, Lon86, Mid84b, Nih84b, SB85, Sko86a, Sko88, Svi88a, Svi88c, Zuc89a, WSML89, MFG82, RU89]. ground [HK84]. groundwater [CMLB85, Ste88, Urc84, Sko86d]. group [TH80]. growth [BDLP87, BKU84, BS89, Cas89, CSG85a, CSG85b, CSG85c, DTL89, DDS85, DH87b, De 83, DV86, FVV86, Fet81, GG83, GPD^{+}84, HPI85, HK84, Kar81, KDC84, Lei86, Maa88, MK86, MH86, Mau82a, Moo89, Nem87, Pat86, RS89, RA85, SR89, SWPW85, SHOP88, Smi80a, Smi87, VH82, Vla88b, Wah81, WS80, ZBBLCSTL86]. guaranteed [KS89b]. Guariso [Lho87]. Gulf [RFJ87]. Gustav [Jol84]. gypsy [BWC83, BWKR87].

H [Ano87i, Gra88b, Gro81, Gro83c, Jac87, Jør82a, Jør87a, Leg84a, Mid84b, Mit87, Sko86c, Ush84a, Ush85]. H. [Jør86e]. Haan [Mid84b]. habitat [Ano83t, FPG89, LF85, SSF^{+}88, SCD85]. habitats [Vla88b]. Hague [Fit84]. Haith [God86b]. Halfon [Jør81d]. Hall [Jør87f, Nih83b, Pac88]. Hallam [Svi88b]. Halldin [Jør81a]. halo [HdJ89]. handling [TH80]. Hard [Bec81]. Hardback [Col83, Gro83a, Gue83, Pac88]. Hardcover [Gra88b, Gra89, Pal89, Wel88, Jør87e, Leg89a, Zuc89a]. harima [OFMS86]. harima-nada [OFMS86]. Harold [Jac87]. harvest [AT83, GMM88]. Harvesting [BC82, LP84, Cha86, Cha88, CT82, Gop80, LR82, LW82, RC89, Ree82, Vla88a]. Harwood [Sko86b]. Hasses [HHM89]. having [BR82a]. Hawaii [Gra88b]. Hazard [BRC84, MC86]. hazardous [Pal89]. hb [Jør82c]. health [Jør87f, OSJ^{+}86]. heathland [DK83a]. Heathlands [Fit84]. heavy [Fon84a, NNP84]. Heidelberg [Hof84, Jør80a, Nih85, Sko85, Sko86d, Som84, Str87a, Ush81b, Ush81a, Ush85, Wul86]. held [Str84a, Ush85]. Heliothis [LP83]. Henderson [Zuc89a]. Henderson-Sellers [Zuc89a]. Herbage [SV89a]. herbivore [SR89]. herbivory [Wol83]. herd [FSK82]. Herendeen [Pat88]. Herts [Mid84b]. heterogeneity [BDVC88, HM88]. heterogeneous [FPG89, Sul88]. Heterotrophic [IUME85]. Hierarchical [Sen89]. high [BG87, Lho87]. higher [Ula88]. higher-level [Ula88]. hills [CR80]. Hiorns [Gat83]. hirsutum [GPD^{+}84]. Historical [Cle87]. historically [Hal88a]. history [BvdD80, BM89, Jen89, Mal85]. Hitchin [Mid84b]. Holland [Nih84a]. Holological [Aok89]. Holt [Ush84c]. home [Wor87]. honey [WHB83]. honeybee [DHRLE89, MR86, RM87]. hooded [FV84]. Horwood [Gro84b, Log87]. Host [Bur86, And83]. Host-parasitoid [Bur86]. host-pathogen [And83]. House [Mau87, Sko86a]. Howard [Nih84b, Nih85, BWC83]. human [Jam87, PM89, Zuc83]. humans [Jen86]. humus [PG89]. hunting [EM84]. Huson [Bac85]. Hutchinson [Jør82a]. hybrid [TSJ87]. Hydractinia [Kar81]. hydroclimate [HS84b]. hydrodynamic [Loe87c]. Hydrodynamics [Gro83b, Mau88, Jør86c, Jør86c, Gro81]. hydrogen [MA82]. hydrographic [Jør82c, Jør83]. hydrologic [GE84, RBM89, RC88, Mid84b]. hydrophobic [SI84]. hypercube [RRW84]. hypertrophic [CAJ^{+}87, VL89]. hypobromous [KEK85]. hypotheses [KBPR87]. hypothesis [MFG82]. hyrax [FVV86].

idella [EF82]. Identification [BGTH82, CH84, CHK84, EP89, TH80, Gen84]. IFIP [Nih84a]. II [BHD89a, Bos81b, CHK84, CSG85b, HHS^{+}86, Hal83b, Kaw86b, LA84b, Smi83b, SKMD80, Whi84b, Gat83]. III [CSG85c, HHS^{+}86, Hal85, Whi84c]. illus [Saw83]. illustrated [Ano82h, Jør81b, Jør81d, Jør82a, Ush81b]. illustrating [PM89]. IMA [Ano89r]. image [SFO^{+}86]. immodest [Bah82]. Impact [BWKR87, Jen82, Loe87c, Gro86, Jør86b]. impacts [KKP^{+}86, PMK85]. impingement [Rag84]. Implementation [Pal89, Sko86e]. Implications [Sko86b, CG87, Kre83, Leg87]. importance [BDVC88, Ula88]. imposing [KST^{+}88]. impounded [Loe87c]. impoundment [CAJ^{+}87]. impoundments [Gro86]. imprecise [LW82]. improve [DJGR88]. Improvement [KSW82]. improving [MRBU89]. incineration [RFJ87]. including [HS87]. Incorporation [LAF81, OGC82]. increase [Leg87]. increased [Jør88, NBF88]. Increasing [Lim85, RA85]. incubation [GHV86]. index [Ano80c, Ano81d, Ano81e, Ano81f, Ano82d, Ano82e, Ano82f, Ano82g, Ano83a, Ano83b, Ano83c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84-28, Ano85c, Ano85d, Ano85e, Ano85f, Ano86e, Ano86b, Ano86c, Ano86d, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88c, Ano88d, Ano88e, Ano88f, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, CEMF84, Day88, LW81, PH84]. India [Bis88]. indices [Lom86]. Indirect [SWH84, HP86, Sum89]. individual [CSG85b, MH86, CSG85a]. individuals [WRSZ87]. Indonesia [MS87b]. industrial [Gro84b, Hal86a, OSJ^{+}86]. industry [MIK81]. inevitability [Cas82]. inference [Loe87b]. infestans [Mic85]. infestation [BWKR87, LC84, TCL82]. Influence [Lom85, BBDK81, HHW88, Kin87]. influenced [BHD89b]. influential [Hal88a]. Information [Pod82, Maa88, OGC82, Vol81]. Information-statistical [Pod82]. inhibition [Kar81]. inhibitory [AMO83, BV81]. initial [Smi87]. Inland [IY80, KI86]. Inlet [SKMD80]. inorganic [LAF81]. input [FM88, Hig86a, Hir85, Lev88, TI85, Var88b]. input-output [Hig86a, Hir85, Lev88, TI85]. inputs [Par82]. insect [Bar80, Bar87, CWTH86, FEB89, SSJ89]. insecticide [Lon88b]. insects [Car84, Rub83]. inshore [KGG87]. instability [Rac87]. Institute [Ano87h, Gro84a]. Int [Mau88]. intakes [Rag84]. integrated [BRC84, BN87, Kir89]. Integrating [SF82, HH86]. Integration [VBV87, DLMP87e]. intelligence [DNHP89, Loe87a, Ryk89, SSF^{+}88]. intensity [EP88, Her89]. interacting [GLLN82, Sko86b]. Interaction [WH86, KHF88, KS89b, TI86]. Interactions [Vyh87, AEO83, CSG85b, EGJ^{+}86, Fle82, HdL84, Kar85, Ras87, SSF^{+}88, Leg84b]. Interactive [Fed84, HHS^{+}86, KMW84, NHTO86]. interception [Jia88, OBPK89]. intercrop [Ade85, VAH^{+}84]. intercropping [Ade85, VAH^{+}84]. Interdependence [SW85]. Interface [Gla88, BBD85]. interference [Jen87, WSWP85]. intermittent [Wul86]. internal [Kam81]. International [Ano80g, Ano87h, Gro80, Gro83b, Gro84a, Jør86c, Lon86, Vid89, Zuc89b, Ano83i, Ush85, Bis88]. interpretation [GMR88]. Interscience [Gro83a, Sko86a, Svi88c]. intertidal [FDC89]. intervals [SO88]. intraseasonal [BHD89b]. intraspecific [CSG85c]. Introduction [Jør84b, Jør84c, BV86, DLMP87c, Gro83a, PPH83, LU82a, Nih84b, Svi88b]. invariant [Hip83]. invasions [Gra88b]. invasive [RU89]. inventory [DW84, Sko88]. inverse [BJ89]. invertebrates [Kar85]. investigate [Dye81, JG88, NBF88]. investigation [Sve85]. Investigations [KBSM88]. involving [Pat86]. irrigated [AMS81, Sko86e]. irrigation [BHD89a]. irruption [Bla83]. ISBN [Ano82h, Ano87i, Ano87h, Baz86, BK86, Col83, ES85, Fit84, Gat83, Gat87, Gat88, God86a, God86b, Gra88a, Gra88b, Gra89, Gro80, Gro81, Gro83a, Gro83b, Gro83c, Gro84b, Gro84a, Gro84c, Gro85, Gue83, Hal86a, Hof84, Jac87, Jam87, Jør81d, Jør82c, Jør82a, Jør83, Jør86e, Jør86b, Jør86c, Jør86d, Jør87a, Jør87b, Jør87c, Jør87d, Jør87e, Jør87f, Koh89, Leg84b, Leg84a, Leg89a, Lho87, LU82a, Log87, Lon86, Lon87, Mau87, Mau88, Mej83, Mid84a, Mid84b, Mit87, Nih81, Nih82, Nih83b, Nih84d, Nih84a, Nih84b, Nih85, Pac88, Pal89, Saw83, SB85, Sko85, Sko86a, Sko86b, Sko86c, Sko86d, Sko88, Som84, Ste88, Str82a, Str84a, Str87a, Svi88a, Svi88b, Svi88c, SV89b, Svi89, Tre89, Ush81b, Ush81a, Ush84a, Ush84b, Ush84c, Ush84d, Ush85]. ISBN [Wel88, Wul86, Zuc88, Zuc89b, Zuc89a]. ISEM [Gro80, Jør81a]. island [LJZ82, Ter86a]. islands [McC86]. Isle [Bla83]. ISSN [Sko86b]. issue [Hal89a]. Italy [Nih84a]. Iterative [Luc88, RSKT88]. Ituri [WF88]. IUFRO [Jør81a]. IV [DH87a]. Ives [GG83]. IX [Str82a, ES85, Svi89].

J [Ano87i, ES85, God86a, God86b, Gra88b, Gro83a, Gro83b, Hof84, Jam87, Jør80a, Lon87, Mej83, Nih83a, Nih84d, SB85, Sko85, Sko86a, Sko86b, Sko86c, Sko86d, Som84, Str84a, Svi89, Tre89, Wel88]. Jacques [Gro83b, Jør86c]. Jamart [Mau88]. James [Gro83a, Gro85, Jør80c, Ush81b]. Jan [Str82a]. January [Ano83k, Ano84q, Ano88s, Ano89z, Bis88]. Japan [KI86, Ver87]. Japanese [NHTO86]. Jaques [Mau88]. Jarvis [Sko85]. Jeffers [Nih83b]. jellyfish [Leg87]. Jena [Jol84]. jeopardy [Ful87]. Jewell [Ush84c]. Joanne [Nih85]. John [Col83, Gro83a, Gro84b, Gue83, Jac87, Jør80c, Leg84b, Nih83b, Nih84b]. Johnsen [Nih82]. Johnson [Mid84b]. joint [CP88]. Jørgensen [Gro80, Mid84a, Nih82, Nih84d, Svi89]. Jr [Wul86]. Judith [Zuc88]. July [Ano80l, Ano82m, Ano83n, Ano84y, Ano86n, Ano87r, Ano88t, Ano89s]. June [Ano80h, Ano81k, Ano82q, Ano83r, Ano84x, Ano85t, Ano88w, Ano89-28, Ush85]. Junk [Fit84].

Karst [FP86]. Kasumigaura [MGN86]. Kentucky [Str84a]. Kessell [Ush81a]. key [RBK81]. Kinetic [DSSM87, MB84]. kinetics [KEK85, QF85, QF88]. Kingwell [Tre89]. Kinneret [VKS83]. Klopatek [Str84a]. knowledge [DNHP89, LSSF89]. knowledge-based [LSSF89]. Kondratyev [Koh89]. Kootenay [BGTH82]. Kremer [Jør80a]. Kruger [Sko85]. Kuantu [HH86]. Kulikov [Gro84c]. Kunreuther [Nih85]. kuvanae [BWC83].

L [God86b, Gra88a, Mid84b, Sko86b, Som84, Str87a, SV89b, Ush84d]. L. [GPD^{+}84]. Laboratory [MB84, HS84a, RSKT88]. laevis [CR80]. lagoon [DLMP87a, DLMP87d, BDLP87, DLMP87b, DLMP87c, DLMP87e]. Lake [Ano87h, BGTH82, Jen82, MFG82, EF82, EF83, FVB81, Fed84, FSTF86, GM86, Hal83a, HS84b, KK84, KBSM88, KH87b, Mau87, MB84, NHTO86, OA82, RS88, RPM86, SD83, SA86, SFT89, TH80, VL89, Var88b, Aok87, Aok89, Hal84, Hal86b, KK84, MGN86, TH80, VKS83, Mau87]. Lakes [Jac87, Aok89, BR82a, Gro81, Gro83c, KBPR87, Lij87, MF86, SI84, TSOS86, TD82, VT87, WSML89, dC81]. Land [Hal86a, AEO83, CMM^{+}86, Ess89, HHW88, HH86, Ter86b, TN83, WF88]. land-use [CMM^{+}86, Ess89, HHW88]. landscape [Sen89]. landscapes [PC86]. Lane [Hal86a]. Langmuir [BCJ82, WH86]. Lansing [Jac87]. Large [HU86, CE87, Gro83c, KH87b, RRW84, SSS81, SV89b]. Large-scale [HU86, CE87]. largemouth [DABG84]. larvae [DS85]. Lassey [Ano83h]. Latin [RRW84]. latitude [Ale82, OBPK89]. lattice [WMP^{+}87]. Lauenroth [Baz86]. Lausanne [Gro81]. law [Bar89, Mik86]. laws [Aok88, Mau82a]. Laxenburg [Ano87h, Gro84a]. layer [KD82]. leachates [Arp83, AR83]. leaching [CMLB85, Gyd84]. lead [WS80]. leaf [Cou84, KHMBD86, Rub83, SH85, SH86]. leaf-eating [Rub83]. Learning [Ewe89, Kau87]. LEC [AEO83]. Lecture [Som84, Ush81b, Ush85, Wul86, Zuc88, BK86]. length [GHV86, Sul88]. Lenz [Hof84]. Lepidoptera [LP83]. Leslie [Jør87e, CL89, FV84, Mal84]. lesser [MSR87]. level [KK82, Ula88, SB85]. levels [FD86, OFMS86, SD83]. Levin [Svi88b]. Levy [Sko86b]. Lewis [Jac87]. LG [MTC85]. LG-2 [MTC85]. Library [Jør80b]. Liège [Gro83b, Mau88, Jør86c]. Life [BM89, BvdD80, Cas89, FWS^{+}81, GPD^{+}84, Jen89, Mal85]. light [CH84, CHK84, EP88, KT89, Smi80a]. Ligneous [Fit84]. lignosellus [MSR87]. limit [JG86]. limitation [De 83]. limitations [Cal88, ODP^{+}89]. Limited [Jac87, Log87, SS82]. limits [MHB84]. lindane [Hal86b]. Lindell [Jør82c, Jør83]. Linear [Gen84, Log87, Luc84, Hip83, JUGH83, JT80, Ker83, LS86a, Mar87, Nih84c, PM82, PE81, Ras87, Wal85, LA84a, Log87]. Linearity [Pat83]. lines [Aiz89]. Linked [Mal84]. Linnerooth [Nih85]. lion [SSS81]. liquefied [Nih85]. List [Ano85m, Ano86j, Ano87p]. litter [FFMB82]. live [KD83]. lived [SR86]. living [Vla88b]. load [KH87b]. loading [BG87, BK84, IDS81]. lobsters [SHL82]. local [AFSK85, FSK82, KB84, OMA^{+}86]. locally [Ush84c]. location [TH88]. locations [TH89]. logarea [Aiz89]. logical [Hul85]. Logistic [Mau82a, Ful87, GMR88, Gop80, Mur88]. logspecies [Aiz89]. logspecies-logarea [Aiz89]. London [ES85, Gat83, Gat87, Gro84c, Gro85, Jør81d, Jør87c, LU82a, Lon87, Nih83b, Pac88, Sko86b, Sko86c, Str82a, Ush84a, Ush84b, Ush84c, Jør81b]. London/New [Gro84c, Ush84a, Ush84b]. London/Orlando [Gat87, Jør87c, Lon87]. long [And83, SR86]. long-lived [SR86]. long-term [And83]. longer [KD83]. Loop [GRS88]. loss [HM88, God86b]. losses [Kin87, Rag84, Reu80, Sum89]. Lotka [Bou85, RC89]. Lotka-volterra [Bou85]. Louisville [Str84a]. low [dC84]. lower [OFMS86]. lowlands [Phi85]. Lowrance [Sko86a]. Ltd [Mid84b, Sko86c].

M [Ano87i, Fit84, Gra88a, Gra89, Jør86e, Jør86d, Log87, Mau88, Nih83a, Pac88, Sko85, Sko86b, Sko86c, Som84, Str84a, SV89b, Svi89, Ush81b, Wel88, Wul86]. Maarel [Fit84]. Macfadyen [Ush84b]. MacMillan [Gra88a, Mej83]. macrophyte [CHK84, VT87]. Macroscopic [MIK81]. Magela [Var89b, HON86, Var88a]. magna [Wul80]. mahoganies [NO84]. maize [WS80]. major [Leg84b]. making [DDM84, Ver87]. Malacosoma [RSG87]. male [Bar82]. mallard [JSC87]. mammals [Ush84c]. man [SHL82, Ano85n]. managed [WH89]. Management [Nih84a, CG87, Fed84, HH86, IHH81, JB87, JG86, Lai87, Lho87, LPV87, Lon88b, MP83, Mid84a, MC86, Nih84d, Pal89, RHC85, RSS83, RJ88, Ryk89, SFT89, SV89b, Ter86b, Ush84c, Vid89, VG82, Wil82, Gat88, Gra88a, Nih84a, Som84, Ush81a, Wel88, Ano82h]. managers [DNHP89, Sil89]. Managing [Tya88]. Manchester [BK86]. Mangel [Wel88]. Mani [BK86]. manipulation [Lon88b]. Mann [Leg84a]. mantchurian [AFSK85]. maps [GS86]. March [Ano81l, Ano82s, Ano85q, Ano87t, Ano88v, BK86]. maricultural [GGBH84]. Marine [Jør80b, AK88, Atk87, CG83, Gal87, GIG81, GMM88, JT80, Jør80a, KEK85, Leg89a, Mau88, OFMS86, Pal88, RFJ87, Sei80, Smi80b, SV89b]. Marketed [Jac87, Sko88]. Markov [CHC89, GMM88, OSW85, SWPW85]. Marsh [Jør80b, BBD85, Har84, Jør80b]. Marsh-Estuarine [Jør80b, Jør80b]. Marshall [Baz86, Jør87b]. marshland [BBD85, OL88a]. marshland-water [BBD85]. Marxist [Kau87, Lon88a]. mass [Cal88, FEB89, ZMRB84]. mass-balance [Cal88]. Material [Lei86, Gro84b, Vol82]. materials [Chr87]. materials-energy [Chr87]. Mathematical [AFSK85, KB81, Nih81, Ost87, Rub83, SKV84, Ade85, Ano83h, Bla83, BMOW82, CS85, DLMP87a, FCC81, FSK82, Gat83, Hus84, KT89, Las83, MSR87, Nih84c, QF85, QF88, RS89, SHL82, Sei80, SFP^{+}89, SA86, VKS83, Jam87, Jør80c, Leg84a, Sko86c, Svi88b]. Mathematics [Hal86a]. matrices [Mal84]. matrix [FV84, Had81, Pat86, WH87a, WH87b, Woo88, WH89]. matter [BG87, LA84a, LA84b]. maxima [Par86]. Maximum [Odu83, Moo89]. May [Ano81p, Ano83j, Ano85p, Ano87s, Ano89y]. McFarlane [Mej83]. McGuffie [Zuc89a]. Meadows [Ano87i]. Mean [WG89, Var87b]. means [FSTF86, KBSM88]. measure [Day88, GB88]. measured [Kam81]. Measurement [Jør86d, Lon88a]. measurements [BDLP87, ES85]. measures [Bos89, Gro86]. mechanics [Nor87]. mechanism [KBPR87, VV83]. mechanistic [Cou84, DS85, KK82]. media [KB81, Sko86d]. medicine [Str82b, Str84b, Str87b]. Mediterranean [Fit84, Car82, SV89a, Sko85]. Mediterranean-type [Sko85]. mellifera [RM87]. members [Gro80]. memory [Vol86]. memphremagog [SA86]. Mendota [Aok89]. mercedis [Mur88]. mercury [HGG^{+}82]. merkusii [MS87b]. mesotrophic [LA84a, LA84b]. metabolic [Smi80b]. metabolite [Nem87]. metal [Fon84a, Fon84b, Fon84c, HON86]. metals [NNP84]. method [AM87, Bar80, Bar82, Bra85, CP88, Day89, Pow84, Rag84, SLG84, TH80].

Methodological [DV86, Ros86]. Methodology [Gro83c, FM88]. Methods [TCS89, AMB87, BM86, BV86, Har84, Jam87, OMA^{+}86, Sve85, Col83, Gro83c, Lij87]. Metridia [BM89]. Mexico [RFJ87]. MI [Jac87, Jac87]. michaelian [KDC84]. Michigan [FD86, VAH^{+}84]. microbial [Böl82, DSSMZG88, Jør87b]. microbiology [Ush84d]. microclimatic [GKS87]. microcomputer [Ewe89, SFO^{+}86]. Microcomputers [Hal89a, HHM89]. microcosm [Gar81, Sci86, SR84]. microcosms [SLG84, Smi80b]. microorganism [Bos81a, Bos81b]. mid [SL84]. mid-Atlantic [SL84]. MIDAS [Tre89]. migration [FGL82, FRCC89, RJBL88]. migrations [Pol85]. Milan [Jol84, Mit87]. Miller [SB85]. mine [NNP84]. mineralization [Bos81a, PG89]. minimal [VS84]. minimodel [LJZ82]. mining [RW87]. Minlake [RS88]. Minshall [Gro85]. mirex [Hal86b, Hal84]. Mitchell [Sko85]. mitchilli [PTS88]. mite [WH87a, WH89]. Mitsch [Nih84d, Str84a, Svi89]. mixedwood [EBSW84]. mixing [PU88]. mixtures [AEO83, ON85]. mmodellierung [Jol84]. Model [Ale82, BGO84, BV81, CMLB85, Cos89, JBL^{+}89, KKP^{+}86, KK84, KBPR87, RI85, YL87, Zuc89b, AR89, Abd80, Ade85, AH80, AMO83, AKL84, AK85, AT83, And83, AR83, AFSK85, Atk87, BHD89b, BvdD80, BCOG88, BBC88, Bax85, BR82a, BDLP87, BT82, Bla83, Bon89, BS89, BDVC88, Bri83, BBD85, BMOW82, CMM^{+}86, Car84, Cha86, CDJS84, CAJ^{+}87, CR80, CL89, CSG85a, DTL89, DAA82, Day82, DABG84, DHRLE89, DK83a, DS85, DLMP87b, DLMP87a, DLMP87c, DLMP87d, DLMP87e, DSSM87, DSSMZG88, DK83b, EP88, EGJ^{+}86, EF82, EEKM82, FVV86, FCC81, Fet81, FM88, FV84, Fon84a, FRCC89, FSK82, FFMB82, GOMC81, GG83, GM86, GIG81, GMM88, GGBH84, GPD^{+}84, Had81, HF83, Hal83a, Hal83b, Hal85, Hal88b, HSP88, HS84a, HBB^{+}84, Har84, HCR81, HIB85, HS84b]. model [HSA85, HGG^{+}82, HPKA80, HL85, IHH81, IDS81, IAL87, JUGH83, JT80, Jia88, JSC87, JKNM82, Jør86f, JKNC^{+}86, KYTS86, KOSS82, KEK85, KWSM89, KDC84, KA84, KR82, Kin87, KHF88, KW84, KS89a, KT89, KH87a, KD82, KSBM84, KBSM88, KGG87, KK82, KS89b, KSW82, Leg87, LS87, LS85, Loe85, Log81, LU82b, LA84b, LC84, LAF81, MSH85, MSR87, Mah88, MH86, ML84, MIK81, MGN86, MR86, Mic85, MM87, MC86, MF86, Nem87, Nih84c, OGC82, Ost87, PU88, PS84, PM89, Phi89, PG89, Pod82, PE81, QF85, QF88, RHC85, RSS83, RC89, Rec84, Ree83, RRW84, RU89, RS88, RBM89, Rit89, RSKT88, RMS89, RPM86, Rub83, RC88, SH86, SHL82, Sei80, SG85, SF82, SS87, SFP^{+}89, SHOP88, Smi83a, Smi80b, SA86, SFT89, SMN85, Sto81]. model [SR86, SKMD80, SM81, Sum85, TH80, TSJ87, TD82, TCL82, Tre89, Uch82, Ush84d, Var87b, Var88a, Var89a, Var89b, Vav82, Ver87, VKDS86, Vla88b, VS84, VT87, VKS83, WGSC84, WC88, WC81, WS80, WHB83, Whi84a, Whi84b, Whi84c, WF88, WC85, WZE86, WH87a, WH87b, WH89, ZBBLCSTL86, ZMRB84, RSG87]. Model-assisted [KBPR87]. Modeling [Ano83t, AMS81, Dah85, IA86, KH88, KR86, LMR84, NNP84, OA82, PLRV88, RW87, Urc84, BV86, Cal88, Fed84, Fon84b, Fon84c, Gro84a, Hul85, KB81, Kir89, Loe87a, Mid84b, Saw83, SCD85, Tho84, Tri87, Ush81a, WO81, Hal83a]. modelled [CHC89, Jen86]. Modelling [Ano83h, AK88, Arp83, Baz86, CMW84a, De 83, EBSW84, Gro80, Hal86b, HHW88, KHMBD86, Lho87, LA84a, LA84b, Mit87, Miy86, OSJ^{+}86, OFMS86, OF81, PMK85, SSJ89, SI84, SFS84, SAT86, Sil89, Str84a, Svi89, VL89, Var87a, VVK80, Zuc89b, Ano80g, Ano83i, Ano88i, Ano88j, Ano89r, AM87, AMB87, Bai86, BM89, Bis88, BG87, BN87, Che89, DWPO85, DV86, FVB81, FPG89, Gar84, Gat88, Gen84, GFI84, Gue83, Gui86, Hal89a, HPI85, HS87, HP83, HU86, JCN85, Jør80d, Jør80c, JJKNM81, KMRV89, Lev88, MS87a, Meh84, Mid84a, MHB84, MRBU89, Mur88, NUK86, Nih84d, NHTO86, Nor85, OSW85, OL88b, Pat86, PC86, RFCC89, Ros86, SSF^{+}88, Sci86, SWPW85, Sko86e, Som82, Str87b, SKV84, Wul86, Zuc89a, Ano89l, Baz86, Str84a, Svi89]. Modelling [Ano87h, Ano88g, Ano88h, Ano89h, Ano89i, Ano89j, Ano89k, God86a, Nih83b]. Models [Bar80, Bar87, DW88, LK87, Sch85, Sko86b, Ågr81, Agn82, Ano83h, Ano87i, BKU84, Bou85, COS83, CO88, Cal88, Cas88, Cer89, CH84, CHK84, CS85, DDS85, DH87b, DeA88, DH87a, DDM84, Ewe89, FWS^{+}81, FNK81, Gar81, Gar88, GB81, Gop80, Gra88a, GGN83, GS86, Gyd84, Hal83b, Hal88a, HdL84, Her88b, Hus84, JB87, Jen87, Jør81a, Jør84d, JKNJ86, JG86, Jør88, Ker83, LSSF89, LW88, Las83, LW81, Leg84a, Lho87, Lim85, Loe87b, Loe88, LS86a, Log87, Lom84, LSV83, Luc84, Maa88, MRM81, MBM^{+}89, Mau88, MB88, MS87b, MB84, Moo89, MRE87, NBF88, ODP^{+}89, Ons88, PPH83, RS89, RJ82, Ras87, RC83, RJBL88, RA85, Sen89, SRM83, Sko86c, SD83, SM82, Sto83, TA86, TI85, Tay88, TCS89, UO84, Ula88, Wah81, WG89, Woo88]. models [Wor87, YW87, Zuc83, dC84, Nih81, Lon86]. Modern [Saw83, Chr87]. Modified [PH84]. moisture [Bon89, Kin87, KHF88]. Molchanova [Gro84c]. monitoring [FSTF86, Sko86e]. monomictic [CAJ^{+}87, EF83]. Mont [Mic85]. Monte [Gro84a, OGC82, RMS89]. Mooney [Gra88b, Jør86e]. moose [Bla83, CTJ81]. Morgan [Gro83a]. morphometric [CMW84a]. mortality [BvdD80, NBF88]. Mortimer [Gro81]. Moscow [Gro84c]. most [Hal88a]. moth [BWC83, BWKR87, GKS87, SG85]. motile [WH86]. mountain [BG87, RBK81]. mouth [Hal86b]. movement [DS85, GG83, HIB85, Wor87]. movements [FPG89, FRCC89]. MRF [RBM89]. multi [Bax85, HSP88, RSKT88, Sen88, TD82]. multi- [Bax85]. multi-annual [TD82]. multi-patch [Sen88]. multi-species [HSP88, RSKT88]. multidimensional [Ree83]. Multiple [Man82, ODP^{+}89, BS82, Cos89, De 83, Pat86]. Multiplicative [MRM81]. Multiseasonal [MP83]. multispecies [BS84, Cha86, Whi84a, Whi84b, Whi84c]. multivariate [Hus82]. mussels [SHL82]. Mutualism [EEKM82, Bou85].

N [God86b, Gro84c, Jør80a, Mej83, Nih83b, RBM89]. nada [OFMS86]. National [Hal86a, Sko88]. Nations [Sko86b]. NATO [Leg89a, Saw83, Wel88]. Natural [Koh89, DNHP89, Gat87, GRS88, Gro83a, Jør87d, PU88, Ryk89, Str87a, TSOS86, Tho84, Vid89, Wil82]. Nature [Fah88]. Nauka [Gro84c]. near [Sey89, TSJ87]. near-optimal [Sey89]. nearest [WRSZ87]. need [JCN85, Lai87]. Needs [Sko88, Dah85]. neighbor [WRSZ87]. neighborhood [WC81]. nematodes [EEKM82]. Neoclassical [GL87]. Neomysis [Mur88]. NERC [Ano89r]. NERC/IMA [Ano89r]. net [LHSS86, Var87a]. Netherlands [God86a, God86b, Jør86b, Tre89, KvdK88]. Network [Her88a, Pat88, AH80]. networks [Aok88, Bra85]. Newby [Mej83]. Niagara [Hal86b]. Nicholas [Svi88a]. Nihoul [Gro83b, Jør86c, Mau88]. nippon [SFS84]. nitrate [HHW88]. Nitrification [AMO83, KGHM86]. nitrogen [BHD89a, BHD89b, Bos81a, BS89, DK83a, KT89, MW80, RSS83, SV89a, SHOP88]. Nixon [Jør80a]. NJ [Jør87f]. No [Fit84, Jør80b, Nih83a, RBM89]. Noctuidae [LP83]. non [Fon84b, Fon84c, GLLN82, GFI84, KDC84, LS86a, Nih84c, PE81]. non-equilibrium [Fon84b, Fon84c]. non-interacting [GLLN82]. non-linear [LS86a, Nih84c, PE81]. non-michaelian [KDC84]. non-target [GFI84]. nonequilibrium [Aiz89]. nonhomogeneous [CHC89]. nonlinear [Atk87, LW88, Loe88, Vla88b, ZBBLCSTL86]. Nonlinearity [GKS87]. nonparametric [CP88]. nonpoint [JUGH83, ZMRB84, Zuc89b]. North [Gra88b, Nih84a, BWKR87, SKMD80, Hof84]. North-Holland [Nih84a]. northeastern [WF88]. Northern [HON86, FRCC89, FSK82, RFCC89]. Norway [CDJS84]. note [Ano80n, Ano80o, Ano85w, Jør82b, dC81, dC84]. Notes [BK86, Som84, Ush85, Wul86, Zuc88, Ush81b]. notion [BHD89a]. November [Ano80m, Ano81m, Ano83s, Ano84w, Ano86k, Ano88u, Ano89x, Jac87, Str82b, Str84b, Str87b]. Number [KR82, Ale82, AFSK85]. numbers [RM87]. numeric [CS89]. Numerical [AV81, HP83, Kaw86a, Kaw86b, KI86, Pol85, RJBL88, DLMP87e, PU88, Pow84]. Nutrient [HON86, AR89, BK84, De 83, God86b, HF80, HCR81, HPKA80, IY80, KH88, Mah88, ODP^{+}89, OFMS86, OL88a, Reu80, Ush84d]. nutrient-biomass [Mah88]. nutrients [Sko85]. nutrition [QF85, QF88, SV89a]. NW [CG83]. Nyssa [RSG87].

O [Ush84d, Sch85]. oak [RSS83, CR80]. Oakleaf [Rub83]. objective [BS82]. objectives [BCOG88]. observational [HP83]. Observations [VKS83, DABG84, VKDS86]. observed [TSJ87]. Ocean [Gro83b, Hal86a, Jør86c, Mau88, Har85]. oceanic [Hul85]. Oceanography [Jør86c, Mau88, Leg84a]. October [Ano80i, Ano83l, Ano84s, Ano85o, Ano87u, Ano88p, Ano89u, Gro81]. Odum [Nih84b]. off [GS86, Hal86b]. Ohrid [OA82]. oil [Nih84c, RFCC89, Uri80]. oilspills [SL84]. Okefenokee [OL88a, PM82]. ökosystem [Jol84]. Okubo [Nih81]. oleae [CF88]. oligotrophic [KK84]. olive [AR89, CF88]. Olsson [Mau87]. once [Jen82]. once-through [Jen82]. one [Log81, Nih83a, Sve85]. one-dimensional [Nih83a]. O’Neill [Jør82a]. onion [CWTH86]. Ontario [Hal84, Hal86b, TH80]. ontogenetical [DTL89]. Ooencyrtus [BWC83]. oosterschelde [KvdK88]. open [DLMP87d]. operation [Ost87].

Operational [Str82a]. Operations [Wel88, KB81]. Opportunities [Sko88]. Optimal [BH84, FGL82, Gop80, Loe85, LW82, Wil82, AT83, FEB89, JB87, Ree82, Sey89, SS82, SFS84, SW84]. optimality [BS84]. optimalization [Ter86b]. Optimization [Sto81, BHD89a, BS82, Cha88, KOSS82, Sto83]. optional [AR89]. oracle [Ano87i]. orchard [WH89]. order [KST^{+}88]. organic [BG87, Gyd84, PM89, SI84, Urc84]. organism [Uch82]. organisms [Ber85, KB84, KI86, Vol82]. oriented [Fed84, Som82]. Orlando [Gat87, Jør87a, Jør87c, Jør87e, Lon87, Ste88]. oryzae [LC84, Lon88b, TCL82]. oscillations [KDC84, URC88]. other [Gro84b, Kau87, KDC84, Rag84]. otter [SF83]. Otus [Nor85, NBF88]. outbreaks [Cas82]. outdoor [RRW84]. outflow [Leg89a, SD83]. Outline [Nih83b]. output [Hig86a, Hir85, Lev88, SKMD80, TI85]. overgrowth [Kar85]. overlap [Lom86]. overview [Ano84z]. owl [Nor85, NBF88]. Oxford [Gro81, Jør86c, Lho87, Mid84a, Mit87, Nih84d, Str84a]. oxidation [Gro88, VVK80]. oxygen [CSG85b, OF81, Sum85, TSOS86, VL89]. oyster [RMS89]. ozonated [KEK85]. ozone [Kin87, KHF88].

P [Fit84, Gat87, Lon87, Mid84b, SB85]. P. [BMJ89]. PA [Jør82a]. Pacific [Ree82]. pacifica [BM89]. Page [Ste88]. Pages [Ano80j, Ano80k, Ano80h, Ano80l, Ano80m, Ano80i, Ano81o, Ano81n, Ano81r, Ano81k, Ano81l, Ano81p, Ano81m, Ano81q, Ano82p, Ano82r, Ano82o, Ano82n, Ano82q, Ano82m, Ano82s, Ano82l, Ano83m, Ano83p, Ano83q, Ano83k, Ano83r, Ano83n, Ano83j, Ano83s, Ano83l, Ano83o, Ano84r, Ano84u, Ano84t, Ano84q, Ano84x, Ano84y, Ano84w, Ano84s, Ano84v, Ano85u, Ano85v, Ano85s, Ano85t, Ano85q, Ano85p, Ano85o, Ano85r, Ano86m, Ano86n, Ano86k, Ano86l, Ano87v, Ano87q, Ano87r, Ano87t, Ano87s, Ano87u, Ano88q, Ano88z, Ano88r, Ano88y, Ano88s, Ano88w, Ano88t, Ano88v, Ano88u, Ano88p, Ano88x, Ano89-29, Ano89w, Ano89t, Ano89v, Ano89z, Ano89-28, Ano89s, Ano89y, Ano89x, Ano89u, Ano89-27]. paleoclimates [Gui86]. Pannell [Tre89]. Panonychus [WH89]. Pansystems [Bai86, Che89]. Paperback [Col83, Gro83a, Jør87d, Nih83b, Tre89, Gro84b, Lon87, Sko88]. Papers [Jør82a, Ano85g]. Parameter [BBC88, JJKNM81, Jør84d, OGC82, ZBBLCSTL86, AM87, AMB87, Atk87, Kre83, Loe88, LW82, NHTO86, RRW84, Var89a]. parameter-based [AM87, AMB87]. parameterized [Bax85]. parameters [BDLP87, FM88, GHV86, KK84, MB84, NO84, VH82, ZBBLCSTL86]. parasite [BWC83]. parasitic [EEKM82]. parasitoid [Bur86]. parasitoids [Bar87]. Paris [Gro83c, Leg84a, Lij87]. park [BT82]. part [BHD89a, Mid84a, Nih84d, Bla83]. Particle [Har85]. particles [VV83]. particular [CF88]. particulate [BG87]. partitions [Hir85]. Pasquill [Gro84b]. Passaic [AH80]. passive [WO81]. pastures [FP86]. patch [Gat87, Sen88]. patches [LF85]. patchy [Vla88b]. Paterson [Mej83]. path [Bra85, Day82, FGL82]. pathogen [And83, Lon86]. pathway [KW84]. pattern [LK87, SR89, Shi80]. patterns [KB84, KHMBD86, LU82b, SO88, WMP^{+}87, Zuc83]. Paul [Jør86b]. pb [Jør82c]. PCB [BMOW82, RFJ87]. PC’s [Bre89]. Pedley [Mej83]. pelt [AFSK85]. Penaeus [CG87, GG83]. Pennisetum [BMJ89]. perch [Jen82]. perennial [CMW84a, CMW84b, JBL^{+}89]. Perfect [IAL87]. performance [TCS89]. period [GHV86, OFMS86]. Periodic [LR82, LP84, LS86b]. permanent [SWH84]. Persistence [Gar81, God86a, HdL84, Sen88]. persistent [PM89, Urc84]. perspective [Cas88, Cle87, HU86, Jør86b]. perturbations [HF80, RC89, Var88b]. perturbed [Atk87]. pest [CWTH86, HM88, MP83, Nut89, Lon86, Wel88]. Pesticide [CMLB85, RJ82, SSJ89]. pesticides [Bar80, DSSM87, DSSMZG88, God86a, GFI84, PLRV88, SSJ89]. pests [Lon88b]. Peter [Ush84c]. pH [BDVC88]. phenol [BV81]. phenology [BMJ89]. phenomena [OL88b]. phenotypic [Kar85]. phosphorus [Ano87h, AV81, BGTH82, CAJ^{+}87, DK83a, FNK81, GE84, IDS81, JKNM82, NUK86, SD83, SA86]. phosphorus-retention [FNK81]. photo [Gro88]. photo-oxidation [Gro88]. photons [Vol85]. photosynthate [SHOP88]. photosynthesis [CH84, CHK84, EP88, OBPK89, VL89]. photosynthesis-light [CH84, CHK84]. Physical [AKL84, Hal86a]. physico [LAF81]. physico-chemical [LAF81]. physics [WRSZ87]. Physio [Tho84]. Physio-chemical [Tho84]. physiological [CMW84a, Sch85]. physiologically [SWPW85]. phyto [Mau82a]. phyto- [Mau82a]. Phytophthora [Mic85]. phytoplankton [AV81, BDLP87, EP88, Hul85, IUME85, JT80, KDC84, KI86, RSKT88, Smi80a, Tay88, WH86]. phytoplankton-zooplankton [RSKT88]. Pickering [God86b]. Pickett [Gat87]. Piedmont [BWKR87]. pine [FCC81, FWS^{+}81, GE84, WGSC84]. Pinus [MW80, MS87b]. plain [HH86, Var89b]. planci [Sey89]. plankton [BBC88, BMOW82, Pow84]. planktonic [Par82, Sci86]. Planning [TN83, CMM^{+}86, DDM84, HH86, Wul86, Jam87, Ste88]. Plant [HPI85, Wel88, And83, Day89, GPD^{+}84, HSP88, HK84, HL85, KH87a, LK87, Nem87, OSW85, Ost87, PTS88, RA85, SR89, SWPW85, SHOP88, Sum89, WC81]. plantation [NO84]. plantations [CE87, MS87b]. plants [Dah85, DH87b, KE84, Rag84, RA85, Sch85, TI86, URC88, VVK80, WSWP85]. Platt [Leg84a]. Plenum [Gro84c, Gro85, Jør87b, LU82a, Saw83]. plot [HHM89]. plotters [HHM89]. Plume [HS87, HS81, HSA85]. plumes [WSML89]. PO19 [Hal86a]. point [Bri83, JUGH83, SWH84, TSJ87, WMP^{+}87]. point-nonpoint [JUGH83]. Poissonet [Fit84]. Poland [Vid89]. polarization [JG88]. Policy [DDM84, Hal86a, Sko86b, BH84, Jør80d, Lon86]. Policy-making [DDM84]. Pollination [Jør87e]. pollutant [ZMRB84]. pollutants [SB85, SI84, Sko86d]. polluted [NNP84]. pollution [Chy84, DLMP87b, DLMP87c, God86b, Jør80c, KA84, KMRV89, Lei86, Saw83, TI85, Zuc89b]. Polunin [Svi88a]. pomonella [GKS87, SG85]. pond [SKV84, WZE86]. ponds [CSG85a, CSG85b, CSG85c]. pool [Ter86a, Wul80]. Population [ÅA80, KI86, Ons88, RSG87, Agn82, AFSK85, BM89, BMJ89, Bla83, BDVC88, BWC83, Car82, Cas89, CF88, CTJ81, DHRLE89, EP89, EM84, EEKM82, FVV86, FCC81, FV84, FSK82, GKS87, Gop80, GPD^{+}84, HdL84, IY80, IUME85, Jen89, KWSM89, LF85, LP84, LK87, Lom88, MSR87, Nor85, NBF88, Nut89, Pat86, PTS88, QF88, RFCC89, SL84, Sen88, SG85, SFS84, Shi80, Smi87, Sto83, SR86, Ter86a, Vla88a, Wah81, WH87b, Woo88, WH89, Wul80, YW87, ZBBLCSTL86, LU82a, Lon87, Ush85]. Population-dynamics [Ons88]. population-growth [Wah81]. population-toxicant [HdL84]. populations [Bar87, Cal88, CT82, DeA88, DH87a, FGS89, Gat83, GFI84, Gra89, HCL83, Leg87, RHC85, SF83, SR89, SSJ89, Smi83a, Smi83b, Van87, Vla88b, WH87a, Ano82h]. porous [KB81, Sko86d]. position [Bur86]. positive [Bou85, Str87a]. possible [KSB^{+}84]. Post [Str87a]. potencies [Vol85]. Potential [BDVC88, BRC84, Jen86, Jen89, KHMBD86, SL84]. power [Bar89, Odu83, PTS88, Rag84, Sil82, Sum89]. pp [Ano82h, Ano87i, Ano87h, Baz86, BK86, ES85, Fit84, Gat83, Gat87, Gat88, God86a, God86b, Gra88a, Gra88b, Gra89, Gro80, Gro81, Gro83a, Gro83b, Gro83c, Gro84b, Gro84a, Gro84c, Gro85, Gue83, Hal86a, Hof84, Jac87, Jam87, Jol84, Jør80a, Jør81b, Jør81c, Jør81d, Jør82c, Jør82a, Jør83, Jør86e, Jør86b, Jør86c, Jør86d, Jør87a, Jør87b, Jør87c, Jør87d, Jør87e, Jør87f, Koh89, Leg84b, Leg84a, Leg89a, Lho87, LU82a, Log87, Lon86, Lon87, Mau87, Mau88, Mej83, Mid84a, Mid84b, Mit87, Nih81, Nih82, Nih83b, Nih84d, Nih84a, Nih84b, Nih85, Pac88, Pal89, Saw83, SB85, Sko85, Sko86a, Sko86b, Sko86c, Sko86d, Sko88, Som84, Ste88, Str82a, Str84a, Str87a, Svi88a, Svi88b, Svi88c, SV89b, Svi89, Tre89, Ush81b, Ush81a, Ush84a]. pp [Ush84b, Ush84c, Ush84d, Ush85, Wel88, Wul86, Zuc88, Zuc89b, Zuc89a]. pp. [Jør80b, Jør81a, Lij87, Nih83a]. PR [Gro84a]. PR-83-28 [Gro84a]. Practical [OL88b]. practice [Lho87]. practices [HHW88]. Praha [Str82b, Str84b, Str87b]. precipitation [Arp83, AR83, KBPR87]. Predation [Leg89b, Shi80, DABG84, MDS84, QF85, QF88]. Predation-affected [Shi80]. predator [BBDK81, BC82, Had81, Kar81, LR82, Leg87, MFG82, QF88, RC89, Sey89, Sul88, Wah81, WH87b]. predators [BBDK81, Lom84, Lom85]. Predicting [RA85, Dah85, KEK85, Kin87, SD83, TSJ87]. prediction [DW84, MS87b, MF86, WG89]. predictions [AR83, DJGR88, GÅ89, KKG82]. Preface [GNIW86]. preliminary [HBB^{+}84, KEK85]. Prentice [Jør87f]. Prentice-Hall [Jør87f]. prescriptions [RWJ89]. Press [Gat83, Gat87, Gra89, Gro84c, Hal86a, Jør80b, Jør81d, Jør82c, Jør87a, Jør87c, Jør87e, Leg84a, LU82a, Lon87, Mej83, Sko88, Ste88, Str82a, SV89b, Ush84a, Ush84b, Ush84c, Ush84d]. prey [BBDK81, BC82, Had81, LR82, Leg87, Lom88, MFG82, Mur88, QF88, RC89, Shi80, Sul88, Wah81, WH87b]. prey-predator [BC82, RC89, Wah81]. primary [CMW84a, LHSS86, Loe87c, Mau82b, RC88, SBL88, VH82]. primer [Zuc89a]. Prince [Jac87]. Principles [Gal87, GL87, RS89, ES85, Jør87f, Nih82]. priori [OGC82]. priority [Hue87]. private [Ver87]. Probabilistic [Hal89b, BBD85, SRM83]. Probability [TH88, CP88]. problem [Sve85, Vla88a]. Problems [BR82a, Cal88, Nih81, CO88, GLLN82, Lei86, Ros86, Wal85, Ush84c]. Proc [Jør86c, Mau88]. Proc. [BK86]. Procavia [FVV86]. procedure [CMLB85, Cos89]. Proceeding [Ush85]. Proceedings [Gro80, Gro81, Gro83b, Jac87, Jør81a, Nih84a, Str84a]. process [Bla83, CHC89, FWS^{+}81, Fon84b, Jia88, KS85, PG89, TD82, Vol81, Vol82]. processes [CMW84a, DSSM87, Las83, ML84, Meh84, Mej83, RC88, Urc84, VVK80, Ano83h]. producers [Whi84a, Whi84b, Whi84c, Whi84b]. product [Nut89]. Production [Rag84, CMW84a, DW84, HPI85, HP83, JT80, LHSS86, Loe87c, Mau82b, MHB84, Pal88, Rac87, RW87, RBK81, RC88, SBL88, SV89a, Smi80a, VH82]. productive [AEO83]. productivity [AR89, CSG85c, DW88, JSC87]. Products [Gro88, AFSK85]. profiles [VH82]. profit [SW85]. profitable [KD83]. prognosis [JKNC^{+}86]. program [HdJ89, JUGH83, LP83]. programming [Atk87]. programs [LW88]. Projecting [AW85]. projections [Ess89]. Prolog [MRBU89]. properties [Hus84, Tay88]. proportions [Ade85]. proposal [Bah82]. protection [KBPR87, Ste88]. protocols [Pat84]. provided [Hal88a]. PRZM [CMLB85]. Ptinus [Nut89]. Publ [Mau87, Nih84a]. public [Ver87]. Publication [Ano84z, Sko86a]. published [AR83]. Publisher [Ano80n, Ano80o, Ano85w]. Publishers [Gro84c, Jac87, Mid84a, Mid84b, Nih84d, Sko86b, Sko86c]. Publishing [Gro81, Jør81c, Saw83]. Pudoc [God86a, God86b, Tre89]. pupation [Car84].

QUAL [WC85]. Qualitative [SM82, VT87, Bos81a, Bos81b, DH87a, HCL83]. qualities [LAF81]. quality [BCJ82, DDM84, Fed84, KH87b, MF86, PU88, Pat84, RC83, RS88, Som82, VKDS86, WC85]. Quebec [SA86]. Quercus [CR80]. queueing [Sjö80].

R [Ano83h, Gat83, Gro85, Jam87, Jør82a, Jør87c, Leg84b, Leg84a, Log87, Lon86, Mej83, Nih83b, SB85, Sko86a, Str84a, Tre89, Ush81b, Ush81a, Wel88, Wul86]. R. [Pat88]. R1 [WC85]. rabies [DAA82, GHV86, Lon87]. radiation [Bon89, Bro81, KP89, OBPK89, Var87a]. radioactive [BJ89, Chy84, KB81]. radiocesium [CR80]. radioecology [Gro84c]. radionuclides [HIB85, KW84]. Raghuva [BMJ89]. Rainfall [RJ88, Dye81, Jia88, Reu80, SV89a, TN83, Var88a]. rainfall-runoff [Var88a]. Ramade [Jør87d]. Random [PC86, Bra85, Man82, SH86, SO88]. randomness [Day89]. range [Loe85, Wor87]. rangeland [HSP88, MSH85, Rit89]. rapidly [Var87b]. Rapporteur [Lij87]. raptor [Smi83b]. rate [Cas89, CSG85c, DSSMZG88, EP88, Fet81, GB88, Smi80a]. Rates [HS81, Mau82b, Ågr81, Abd80, DH87b, HF80, Sul88]. ratio [AT83, FEB89]. ration [MM87]. ration-based [MM87]. Rational [FSTF86]. Raton [Ush84d]. reactions [Bos83, LAF81]. reactor [AH80]. Real [Jør87e]. reappraisal [LC84]. rearing [FEB89]. rebutttal [Odu83]. recapture [MB88]. reclamation [TN83]. reconstituted [Sci86]. reconstructing [Gui86]. records [AR83]. recovery [KB81, SL84]. recreation [RRW84]. recruitment [CG87]. recurrent [Mal85]. recursive [Gen84]. red [Kam81, KI86, WH89]. redistribution [MTC85]. Reduced [Smi83b]. reduction [IDS81]. Reductionism [Ros86]. reductions [KH87b]. reference [CF88, Lei86]. REFLEX [Tri87]. regeneration [SFP^{+}89, WF88]. regime [OA82]. regimes [Bon89]. region [CG83, SL84, Sul88, Var87a, Vid89, Zuc88]. Regional [Wal87, AW85, BN87, CMM^{+}86, CE87, Fed84, RC88, SH85, SMR86, TI85]. regression [Aiz89, KS89b, LW88, Mur88, PE81]. regression-based [PE81]. regular [SO88]. regulate [Jen89]. regulation [CTJ81]. Reidel [Jør86b]. related [Ros86]. relation [BG87, DW88, SV89a, Wul80]. relations [Ber85, Hul85]. Relationship [Cou81, NO84, OBPK89, BBD85, Day88, EP88, IUME85]. relationships [Ano83t, HM88, KSB^{+}84, SRM83]. relative [DLMP87d, Fet81]. Relaxation [KDC84]. release [Bar80, Bar82]. released [NUK86]. releases [Bar87]. relevant [BCOG88]. reliability [DLMP87e, LW81, WC88]. remote [Jør86d, KYTS86, Jør86c]. remotely [Maa88, SFO^{+}86]. remotely-sensed [Maa88]. removal [HHW88]. renewable [SW85, VG82, Wal85, Wal87, Som84]. replacement [DKR84, Rac87]. Reply [Pat88, Bac85]. report [GG89]. representative [IA86]. reproduction [AK88]. Republic [Jol84, Pal89, Zuc88]. Research [BCOG88, Hal86a, Mau87, Sko88, CWTH86, Cle87, Hal89a, Str82a, Ush84b]. reserve [SSS81]. reservoir [BG87, KOSS82, LAF81, MTC85, Sum85, WC85]. reservoirs [BR82a, Miy86, Gro83c, Lij87]. Residence [Hig86b, Her89]. residents [OMA^{+}86]. resilient [Ver87]. Resistance [HF80, Kar81, Lon88b]. resolution [Cos89, dC84]. Resource [Lai87, Ush81a, Bla83, BCJ82, CO88, DNHP89, Rac87, Ryk89, Sen89, SFP^{+}89, Som84, VBV87, VG82, Wal85, Wal87, Wil82, WSWP85, DH87a]. Resources [Hal86a, Gla88, SW85, Sko88, Tya88, Uri80, Vid89, Jør87d]. respiration [ÅA80, Jør81c, Mau82b]. response [Cas89, Gui86, Jør86e, QF88, RA85, RBM89, SLG84, ZMRB84]. responses [CMW84b, Dah85]. resprouting [SR86]. Ress [God86a]. resulting [Reu80]. Results [BM89, Gro88, HF83]. retention [FNK81]. revegetation [KYTS86]. Review [Gro84a, CS85, Jør80d, Loe89, Wor87]. reviewer [Ano85m]. reviewers [Ano87p]. revisited [CL89]. rhythmical [FDC89]. rice [Lon88b]. Richard [Jør80b]. Rieser [Zuc88]. rigour [MRBU89]. Rinaldi [Nih84a]. ring [Gui86]. riparian [HHW88]. rise [HSA85, HS87]. rise/dispersion [HSA85]. risk [FGS89, Fon84a, Jør87f, LPV87, RJ88, SW85, Pal89]. river [Hal86b, IA86, KKG82, Loe87c, PMK85, AH80, AM81, BK84, PMK85, SMR86]. Robindson [Sko86b]. Robinson [Ano87i]. Robust [Loe88]. Rock [SHL82, Wul80]. Rock-lobsters [SHL82]. rocky [RBK81]. Role [BKU84, Her88b, SR89, URC88, Ano85n, BWC83, DDM84, Jør86d, Leg89a, MRM81, Sko85, Wol83]. roles [Ons88]. roller [Rub83]. Romane [Fit84]. Rome [Nih84a]. Ronald [Jør86b]. rookery [OL88a]. Root [CMLB85, Bos81a, Bos81b, Day82]. root-microorganism [Bos81a, Bos81b]. root-zone [Day82]. rooting [SWH84]. roots [Chr87]. Ross [Jør82a]. roughness [HS87]. Royale [Bla83]. Ruesink [God86a]. rule [SFT89]. rule-based [SFT89]. run [LP83]. runoff [Var88a, ZMRB84]. rusty [KWSM89].

S [BK86, Gat87, God86a, Gro80, Jac87, Jør80a, Jør80c, Jør87a, Jør87f, Lon87, Mid84a, Nih82, Nih84d, Nih84a, Sko88, Ste88, Svi88b, Svi89, Tre89, Ush81a, Ush84d]. S.W [KvdK88]. salinity [Tya88]. SALMO [BR82a]. salmon [Ree82]. Salt [MFG82, AMS81, Har84]. saltmarsh [CM87]. sample [FP86]. San [Ano82h, Jør81d, Ush84a]. sand [CR80]. Santee [PMK85]. satellite [BN87]. saturation [BBDK81]. savanna [FFMB82]. saving [Hue87]. scale [CE87, DJGR88, HU86]. scales [Man82]. scaling [SLG84]. scenario [Hue87]. Schmidt [Fit84]. school [DW88]. Schuda [Ush81b]. Science [Fit84, Hal86a, Jør80b, Mid84a, Nih84d, Nih82]. Sciences [Hal86a, Leg89a, Wel88, Ano89r, Svi88c]. Scientific [Gro81, DNHP89]. Scolytidae [DS85, WGSC84]. SCOPE [Gue83, Jør86d, Leg84b, JG86, SB85]. Scotian [HP83]. screech [Nor85, NBF88]. Sea [Hal86a, KI86, AMO83, Hof84, Jør86c, MM87, SF83, SKMD80, IY80, Par86, Pod82, Pol85]. seabird [SL84]. seal [FV84, FRCC89, FSK82]. seals [RFCC89]. search [KSB^{+}84, Var89a]. Seasonal [RM87, FRCC89, JBL^{+}89, SA86, Var87a, WH87a, Kaw86b]. seasonality [CG87]. Seattle [Str82a]. seawater [KGHM86]. seaweed [MM87]. seaweed-sea [MM87]. sector [Ver87]. sediment [JKNM82, KD82, MB84]. sedimentation [Har85, Lij87]. sediments [NUK86]. Seed [McC86, SF82]. Selected [SV89b]. Selection [JB87, Mur88, Var89a, Zuc89b]. selective [CT82, Ree82]. self [KBPR87]. self-protection [KBPR87]. Sellers [Zuc89a]. semantic [CS89]. semi [Loe85, SWPW85]. semi-arid [SWPW85]. semi-desert [Loe85]. seminar [Bis88]. sensed [Maa88, SFO^{+}86]. sensing [Jør86c, KYTS86, Jør86d]. Sensitivity [Bac85, Bos89, HIB85, MSH85, Bax85, DJGR88, DK83b, GOMC81, GB88, GKS87, Hus82, Hus84, Ker83, KSW82, LS85, Mah88, MRM81, Rec84, RRW84, SM81, Var88b]. separability [Luc84]. September [Ano81q, Ano82l, Ano83o, Ano84v, Ano85r, Ano86l, Ano88x, Ano89-27, Gro80, Jør81a, Nih84a, Vid89]. Series [Jør86c, Jør87c, Leg89a, Lon86, Mau88, SV89b, Wel88]. services [OSJ^{+}86]. sessile [Kar85]. set [BR82b, FM88]. Seto [IY80, KI86]. sets [BM86]. settlement [Zuc83]. several [Bos81b, Hal88a]. sewage [Ost87]. Sex [AT83, CT82, Ree82, FEB89]. Sex-selective [CT82, Ree82]. shade [ON85]. shading [KP89, TI86]. shallow [KSBM84, KH87b, TD82, VL89]. shape [KP89]. Shapes [Gre83, GGN83]. shear [HS87]. Sheehan [SB85]. Sheehy [ES85]. shelf [HP83, HPKA80, Jør86c]. Sherman [SV89b]. shocks [Koh89]. Short [LP83, OFMS86, Sch85]. short-period [OFMS86]. show [Jør88]. shrimp [CG87, GG83, GMM88, GGBH84, KGG87]. shrub [Mal85, SMN85, SR86]. shrubs [Fet81]. Shugart [Jør82a]. Sidney [Ush84c]. Sign [LU82b]. sign-stable [LU82b]. Significance [GÅ89]. SILVA [KA84]. Simple [Ågr81, Jen87, AEO83, Ano83h, Arp83, Bra85, Day82, JKNM82, Las83, Sci86, Smi80b, Var87b, Var88a, VV83, Wah81]. simplification [RI85]. Simplified [Mah88, DDM84]. simulate [Ewe89]. Simulated [BHD89a, EM84, Nut89, Gre83, Sen89]. Simulating [CTJ81, MTC85, Par86, SH85, Sum89, SR84, Pow84, SSF^{+}88]. Simulation [Bel84, CG87, CF88, CMW84b, DKR84, Had81, Jen89, KS85, KMRV89, KP89, MW80, Mal85, RFJ87, RFCC89, Reu80, Smi83a, Smi83b, SSS81, Uch82, WMP^{+}87, Aiz89, And83, BHD89b, BM89, BMJ89, Bos86, BS89, BWKR87, CG83, CWTH86, CM87, Cou84, CE87, DAA82, DHRLE89, DK83a, DS85, DK83b, Fed84, Fet81, FRCC89, Gal87, GG83, GRS88, GKS87, GE84, GIG81, HHS^{+}86, Hal84, HS84a, IHH81, JG88, Jør80a, Jør80b, KB81, Kam81, Kar81, KB84, KWSM89, KW84, KMW84, KH87a, KBSM88, Kre83, LHSS86, Lom85, Lom88, Lon88b, LSV83, LAF81, MRM81, Mal84, Mic85, Moo89, NBF88, Ons88, PS84, Pol85, RHC85, RS89, RJ82, Rec84, RS88, Rit89, RSKT88, SF83, SV89a, SG85, SCD85, SMN85, Str87b, SKMD80, Sum85, Svi88c, TCS89]. simulation [Var88b, WS80, Whi84a, Whi84b, Whi84c, WZE86, RSG87, Str82b, Str84b, Gro84a, Jol84, Jør80b, Mit87, God86b, Jør81c]. Simulations [BWC83, SH86, CH84, Hal89b, KD82, KK82, WSML89]. simulator [CS89]. simultaneously [Kam81]. simultaneously-measured [Kam81]. single [HSA85]. singular [Sen88]. Sisy [Str82b, Str84b, Ano84-27, Str87b]. site [BJ89, Day88, Pat84]. site-specific [Pat84]. siting [Nih85]. Sitophilus [LC84, Lon88b, TCL82]. situ [BDLP87]. Size [TI86, BBC88, CSG85b, CSG85c, DH87b, DW88, Jen89, Pal88, WRSZ87]. size-dependent [BBC88]. SKEBUB [Bax85]. Skogerboe [Baz86]. Skowronski [Som84]. Skreslet [Leg89a]. slicks [Nih84c]. slow [KDC84]. slow-fast [KDC84]. sludge [ML84, VV83]. small [Mid84b, Smi83a, Smi83b, WC85]. Smethurst [Jør81b]. Smith [Gro84b, Ush84d]. social [Ano87i]. Society [Gro80, Mid84b, Saw83, Zuc89b]. Socio [Umb89]. Socio-economic [Umb89]. sociobioeconomic [KGG87]. soft [Bec81, Wul86]. softcover [BK86]. software [GG89, HHS^{+}86]. Soil [Gro84c, Arp83, AR83, Bon89, Bos81a, Bos81b, Bos83, Day82, DSSM87, DSSMZG88, GB88, GÅ89, God86b, Gyd84, HL85, KR86, KE84, MBM^{+}89, Phi89, PG89, Reu80, SV89a, SS87, SLG84, SHOP88, Sko86d, God86a, Sko88, Ush84d]. soil-plant-water [HL85]. soils [AMS81, KKP^{+}86, RJ82]. solar [Bon89, Bro81, Var87b, WZE86]. solar-algae [WZE86]. solisianus [EGJ^{+}86]. solute [Day82, Nih83a, RBM89, SS87]. solution [Ree82, SBL88]. solutions [Bre89, Ker83, Nih83a]. Some [BCJ82, GLLN82, BKU84, CL89, GG89, KD82, NO84, Sve85]. sonches [KMW84, KBSM88]. Sons [Gro84b, Jac87, Jør80c, Leg84b]. Soule [Gra89]. source [JUGH83, ZMRB84, Zuc89b]. sources [Loe87b, TSJ87, Gro84b]. South [Jør80b, Gro86, PMK85]. Southern [CDJS84, FCC81, FWS^{+}81, WGSC84, Kam81, VAH^{+}84]. soybean [KHF88]. soybeans [VAH^{+}84]. Spartina [MHB84]. Spatial [HM88, LF85, BDVC88, EGJ^{+}86, LK87, PC86, SH86, Shi80, SCD85, SO88, SA86, TCS89, WF88, WSWP85]. spatially [Sul88]. spatiotemporal [SH85]. Special [Jør87c, Hal89a, LR82, Lei86]. speciation [Fon84a, Fon84b, Fon84c]. species [Abd80, Agn82, Ale82, BvdD80, Bri83, Cha88, DKR84, Gla88, HSP88, HK84, KB81, KR82, ON85, RU89, RSKT88, Smi83b, Ter86a]. specific [Bar89, Pat84, Smi80a]. spectrum [Pal88]. Spiller [Zuc88]. spills [RFCC89, WSML89]. spline [FSTF86]. spread [CHC89, GGN83, HS81, RU89]. spreading [Nih84c]. Spreadsheets [Bre89]. Springer [BK86, Gra88b, Hof84, Jør80a, Jør86e, Leg89a, Nih81, Nih85, Pal89, Sko85, Sko86d, Som84, Str87a, Svi88b, Ush81b, Ush81a, Ush85, Wel88, Wul86, Zuc88]. Springer-Verlag [Hof84, Jør80a, Nih81, Nih85, Som84, Str87a, Ush81a]. springfed [HHW88]. spruce [BS89, Cas82, KSB^{+}84]. spruce-fir [KSB^{+}84]. Spurious [Gar84]. squared [WG89]. squirrels [AFSK85]. Stability [Agn82, GRS88, Phi85, Rac87, Bos81a, Bos81b, LU82b, Lom85, Rec85, Sve85, TA86]. stabilization [Gop80]. stable [KR82, LU82b]. stack [HSA85]. stages [BvdD80]. Stand [MH86, ON85, NO84, OBPK89]. standards [Pat84]. stands [AM87, AMB87, RBK81]. Starfield [Gra88a]. State [Baz86, Ano83i, Aok88, DLMP87c, Phi89, Gro80]. State-of-the-art [Baz86, Ano83i, Gro80]. states [BR82a, SL84]. static [PM82]. stations [FSTF86]. Statistical [BDLP87, FNK81, LS86b, Pod82, UO84, WRSZ87]. statistics [YW87]. status [Bos81a]. Steady [DLMP87c, Aok88]. Steady-state [DLMP87c]. Stechlin [KK84]. Stephen [Ush81a]. Stepwise [RSKT88]. sterile [Bar80, Bar82, Bar87]. sterling [Jam87, Svi88c]. Stevan [Lij87]. Stig [Leg89a]. Stinner [Sko86a]. Stochastic [AM81, Fle82, Fle87, BT82, Cer89, DABG84, EGJ^{+}86, Gar88, GG83, Gyd84, JB87, JT80, KYTS86, Kar85, Kre83, NBF88, Smi83a, SMN85, Sve85, TH89, Wil82]. stochastic-deterministic [SMN85]. stochasticity [AK85]. stock [Jen82]. Stockholm [Mau87]. stocking [BH84, CSG85c, Loe85, Sen89]. stocks [GLLN82]. stomatal [URC88]. storage [CE87]. stored [KWSM89, LC84, Lon88b, Nut89, TCL82]. stored-product [Nut89]. Storgama [CDJS84]. Storms [VKS83]. stormwater [HH86]. Straskraba [Jol84, Svi89, Mit87]. Strategic [Lon86]. Strategies [DeA88, BS87, HK84, JB87, Lai87, LR82, Mal85, SSS81]. strategy [Rac87, SS82]. stratification [HS84b, Tay88]. stratified [Miy86]. Straus [Wul80]. stream [CHK84, GOMC81, HHW88, HGG^{+}82, OGC82, TH89, Gro85]. streamflow [Dye81]. streams [PU88, TH88]. streamwater [CDJS84]. strength [Day88]. stress [Kin87, KHF88]. Strobeck [Ush85]. Stroudsberg [Jør82a]. Structural [Jør86f, BS87, Jør88, Rec85, Vol85, Vol86]. Structure [EF83, BCOG88, Hal83a, Hal83b, Hal85, Mik86, OBPK89, QF88, SKMD80, Ter86a, Van87, Wul80]. structured [FGS89, Vla88a, Vla88b]. structures [CS89, CL89, Hal83b, OMA^{+}86]. Structuring [BM86]. Studies [Gra88b, Nih82, Nih83b, Sko85, Wul86, Zuc88, Bro81, Jør86e]. Study [DSSMZG88, Gro84b, NUK86, Sci86, AK88, Aok89, BWKR87, CMW84b, HS84a, HH86, HL85, Kar81, Lho87, Lom85, Lom88, Lon88b, Nor85, NBF88, RBM89, SV89a, TSOS86, Var88b, Ver87, Wul86]. Stumm [Gro83a]. sturgeon [BH84]. Subject [Ano84-28, BHD89a, MDS84]. submodel [KD82]. Submodels [BK84, Böl82]. subsequent [SFP^{+}89]. substance [Hal83a, Jør84d]. substances [Tho84, Urc84, Van87]. substrate [BV81]. substratum [KB84]. substratum-bound [KB84]. subsystem [SKMD80]. subtrates [VV83]. succession [BWKR87, DKR84, EBSW84, LA84b, Mal84, Rob89, SCD85, SMN85]. sufficiency [HS84a, dC84]. sulfur [Cou81]. sulphate [CDJS84]. sulphur [TSJ87]. summary [BRC84]. sun [KP89]. Sündermann [Hof84]. supercomputer [Ons88]. supply [AR89, MK86]. support [Hal88a]. surface [Chy84, HS87, MBM^{+}89, Sko86d, Var87b]. survival [BDVC88, GG83, LF85]. Sussex [Col83, Gro83a, Hal86a, Nih84b]. sustainable [GL87]. swamp [PM82]. swards [BHD89a, BHD89b]. swarming [MR86]. Sweden [Jør81a, LJZ82]. Swedish [Mau87]. Swen [Jør81a]. switching [MDS84]. Switzerland [Gro81]. Sydney [Ush84a]. Symposia [SV89b]. Symposium [Str82b, Ano80g, Gro81, Str84a, Str84b, Str87b]. Synergism [KK82]. Synthesis [GM86, Wul86]. System [GP85, Pat84, Ade85, AV81, BBDK81, Bos81a, Bos81b, Bri83, EM84, Had81, HON86, Jen82, KMW84, KBSM88, KS89b, MDS84, NNP84, Pod82, Rac87, RWJ89, Rit89, SH85, Sci86, Sen88, Sto81, SW84, SMR86, SFO^{+}86, Sul88, Ter86b, UO84, VAH^{+}84, Vyh87, WG89, Tre89]. Systematics [Jør87c, Jør87c]. Systems [Ano87h, Bos86, CG83, Col83, Gro84a, Lon86, Baz86, Bec81, BCJ82, CWTH86, CH84, CHK84, Fon84b, Fon84c, GGBH84, HF80, Hip83, HU86, Jør80b, Jør81d, LSSF89, LR82, LU82a, Log87, Mar87, Nih84a, NHTO86, OSJ^{+}86, Pat86, PPH83, Ryk89, SS87, Str82b, Str84b, Str87b, SR84, Tho84, Umb89, Urc84, Vid89, Wel88, Mej83, Str87a, Jør82a, Nih84b, Svi88c].

T [Gat87, Jør81c, Jør82c, Jør83, Jør87a, Leg84a, Mid84b, Nih84b, Sko85, Som84, Svi88b, Ush84a]. table [Cas89, GPD^{+}84]. tables [Nih83b]. tactical [Lon86]. tailed [EM84]. tailings [NNP84]. Taipei [HH86]. Taiwan [HH86]. taken [Rub83]. Tamar [HBB^{+}84]. target [GFI84]. taxonomy [Bah82]. Technical [Mid84b, Nih83a]. technique [Dye81, FSTF86, KKG82, LS86b, Var89a]. techniques [Bax85, Fon84a, Gui86, Loe87a]. technology [Nih82]. tectus [Nut89]. Telford [Jør81b]. TELOC [WS80]. temperate [dC81]. Temperature [Lon88b, CSG85b, IUME85, Kaw86a, Kaw86b, KKG82, PLRV88, Smi80a, Var87b, Böl82]. Temporal [Var88b, BDVC88]. tent [RSG87]. term [And83]. terms [Vol81]. terrestrial [Ano83t, AK88, Jør86d, Var87b, Wol83]. territoriality [Lom85]. test [Car84, CEMF84, SD83]. Testing [HS84a, KSBM84, Gro85, Hal85]. Tests [Lom86, KD82]. Texas [CG87, GG83, KGG87]. Th [Nih83a]. Their [Sko86b, Hal88a, Leg84b]. theoretic [VG82]. Theoretical [Ågr81, FDC89, JUGH83, MHB84, ÅA80, CO88, Hal88a, Jia88, Phi85, Smi80a, Ush81b, Jør81d]. theoretically [Ree82]. theories [Fah88, Leo83]. Theory [AM87, Cas88, GB88, Lon88a, Zuc83, BR82b, Fle82, Fle84, Fle87, Fon84b, Gat83, GP85, Loe85, Loe89, MP83, Ons88, Pat84, QF85, QF88, Rob89, Sil82, Sjö80, Svi88a, Wal85, WSWP85, CH84, Gro85]. there [FEB89, Hal83a, Hal83b, Hal85]. thermal [Bon89, OA82]. thermocline [dC81]. Thermodynamic [MS87a]. Thermodynamics [Meh84]. Third [Str84b]. Thomas [Jør81b]. Thornley [Sko86c]. threats [CMLB85]. Three [RI85, VKDS86, DLMP87e, KD82, KK82, Mau88]. Three-dimensional [VKDS86, DLMP87e, Mau88]. three-layer [KD82]. three-trophic [KK82]. threshold [HF83]. Tidal [DLMP87e, DLMP87c]. tide [Kam81, KI86]. time [Abd80, Agn82, FP86, GPD^{+}84, Her89, Hig86b, Hip83, HPKA80, Jen86, Ker83, LK87, Man82, OSW85, Pat86, SWPW85, TD82, Kir89]. time-dependent [HPKA80]. time-invariant [Hip83]. time-varying [Abd80]. TIME-ZERO [Kir89]. times [Man82]. Tokyo [Jør86c, Mit87, Nih84d, Str84a, Str87a, Ush85, Wul86]. tolerances [ON85]. tomatoes [VAH^{+}84]. Tomczak [Wul86]. Tomlinson [Jør86b]. tool [Jør88]. tools [HHS^{+}86, Leo83]. top [Hue87]. Toronto [Jør80c, Ush84a]. Total [SU87, TI85]. toxic [BRC84, Fon84c, Hal83a, Jør84d, KB81, Tho84, Urc84]. toxicant [DH87a, HdL84]. toxicants [DH87a, HCL83, SR84]. toxins [HBB^{+}84]. trace [HON86]. trade [Joy85]. traits [CMW84a]. transboundary [Zuc88]. transfer [AM81, Jen86, KE84, Ras87]. Transferring [DNHP89]. transient [HCR81, Smi80b]. transition [WH87a, WH87b, Woo88, WH89]. transpiration [Jør81c, URC88]. transport [AMS81, KR86, KW84, Nih83a, Nih84c, NNP84, PLRV88, Pow84, RBM89, SS87, Som82, Urc84, VKDS86, WO81]. transport-oriented [Som82]. Travis [Str87a]. treatment [BDLP87, DV86, Jør81b, Ost87, VVK80, Vav82]. tree [DDS85, FWS^{+}81, Gui86, KP89, LPV87]. tree-ring [Gui86]. trees [AR89, MH86]. trends [Cle87]. trials [SSJ89]. Trihalomethane [FD86]. tritium [WHB83]. Troester [God86a]. trophic [BR82a, Her88a, KK82, KS89b, Luc88, OFMS86, Pat88]. tropical [BS89, PS84, Var87a, Var88a, Var89b]. true [DLMP87c]. Tubbs [God86b]. tumor [DV86]. tupelo [RSG87]. turbulent [WO81]. Turkey [CR80]. Turnover [Van87]. Two [MS87a, Abd80, Agn82, BvdD80, Cha88, DLMP87d, GHV86, GLLN82, HK84, KD83, KK82, Lom84, SS87, Smi83b, SD83, WH87b, WMP^{+}87, ZBBLCSTL86]. two-dimensional [DLMP87d, WMP^{+}87]. two-predator [WH87b]. two-prey [WH87b]. two-species [Abd80, Agn82, Cha88]. type [JBL^{+}89, Sko85]. typhoides [BMJ89].

U [Col83, Jac87, Jør80c, Jør87a, Jør87f, Lon87, Sko85, Sko88, Ste88, Ush81a, Ush84d]. U.S [Gro81, Jør81d, Nih81, Nih83a, Ush81b]. U.S.A [Gat87, Nih84d]. U.S.A. [AW85, GGBH84]. U.S.E.D [HS84b]. Ulanowicz [Leg84a]. ulmi [WH89]. Uncertainty [FVB81, Atk87, BBC88, DJGR88, Fon84a, Gro84a, Har84, Kre83, TN83, WC88]. undiscovered [Uri80]. unequal [Bar82, Lon88a]. UNESCO [Gro83c, Leg84a, Lij87]. Ungulate [Sto83, Sto81, SW84]. Unifying [Sko86a]. unionized [CSG85b]. unique [Bri83]. unit [ZMRB84]. unit-mass [ZMRB84]. United [Sko86b, SL84]. uniting [CMW84a]. univariate [Bax85]. University [Gra89, Jør80b, Jør82c, Sko86b, Str82a, Ush85]. unsaturated [PLRV88, Sko86d]. upland [RSS83]. upon [JKNC^{+}86, QF85, QF88]. upper [AM81]. Uppsala [Jør81a]. uptake [Jen86, MH86, MA82, PM89, WHB83]. upwelling [CG83]. uranium [KB81]. urchin [MM87]. US$11.50 [Sko88]. US$13.60 [Som84]. US$19.10 [Nih85]. US$24.00. [Ush85]. US$35.00 [Jør87e]. US$40.50. [Sko85]. US$41.00 [SV89b]. US$44.00 [Mau87]. US$45.00 [Jør87b]. US$45.00. [Ush84b]. US$45.00/£ [Ste88]. US$49.00/softcover [Jør87e]. US$53.25/Dfl.125.00 [Nih82]. US$55.00 [Gro85]. US$70.00. [Ano87h]. US$79.50/US$90.00 [Ush84d]. USA [Hal86a]. usage [KMW84]. Use [Hul85, Jør88, KH87a, Lom86, Maa88, Mur88, YW87, BT82, CMM^{+}86, Cou84, Ess89, HHW88, HH86, Hul89, KK84, MRBU89, Sen89, TCL82, Tya88, URC88, WF88]. used [Hal88a, Moo89, MRE87]. Usher [Pac88]. Using [DJGR88, BR82b, DNHP89, GMM88, Had81, Hal89b, KS85, Kam81, LSSF89, OL88b, OA82, Ras87, RSKT88, WC88, WC85]. USPR [HSA85]. Utah [MFG82]. utilising [BV81]. utilization [Nem87, SMR86].

V [Baz86, Gro83c, Gro84c, Jør82a]. vacinity [KB81]. validate [MRE87]. Validation [GB81, JKNC^{+}86, KW84, Hal89b, SKMD80, Whi84a]. value [Lon88a]. values [GÅ89]. VanderZwang [Zuc88]. variability [GKS87, SV89b]. variable [HS87]. variables [JBL^{+}89, LSV83, WO81]. variance [LS86b]. variation [Kar85, SA86, Var87a]. variations [Ågr81, Gui86, KEK85, Kaw86a, Kaw86b, OFMS86, PC86]. various [BR82a]. varying [Abd80, GPD^{+}84, ON85]. VEB [Jol84]. Vegetation [Fit84, CEMF84, HHW88, Jør86d, McC86, SW84, Fit84]. vegetation-ungulate [SW84]. velocities [DLMP87c, SF82]. Venice [BDLP87, DLMP87b, DLMP87a, DLMP87c, DLMP87d, DLMP87e]. Verification [HSA85, KT89, KK84, RJ82]. Verifying [AR83]. Verlag [Hof84, Jol84, Jør80a, Nih81, Nih85, Som84, Str87a, Ush81b, Ush81a]. vertebrate [GFI84]. vertebrates [Ano83t]. vertical [Tay88, VH82]. vessels [FGL82]. VI [Ush84a, JG86]. Viable [Gra89]. vicuña [RHC85]. viewpoint [Aok89]. vii [BK86, Nih84d, SB85]. viii [Wul86]. Vincent [Som84]. virescens [LP83]. Voinov [Ano87h]. Vol [Gro83c, Jør87c, Lho87, Mit87, Saw83, Sko86b, Str84a, Ush81b, Ush84a, Ano86e, Ano88g, Ano88h, Ano88i, Ano88j, Ano89h, Ano89i, Ano89j, Ano89l, Ano89k, Jør87a, Jør87b, Ush84b]. vole [Lom88]. volterra [Bou85, Mar87, RC89, SM82]. Volume [Jør87c]. vulpine [DAA82].

W [Col83, Fit84, Gat83, Gat88, God86a, Gro81, Hof84, Jør80a, Jør80b, Nih83a, Nih84d, Ste88, Str84a, Str87a, Svi88c, Svi89, Wul86]. Wageningen [God86a, God86b, Jør81c, Tre89]. walleye [Jen89]. warm [CAJ^{+}87, EF83]. Washington [Hal86a, Sko88, Str82a]. waste [BJ89, Jør82c, Jør83, Pal89, RFJ87, VVK80]. wastes [Hal86a]. Water [KE84, Som82, AR89, AMO83, Böl82, BBD85, BCJ82, CM87, Chy84, Cou84, Fed84, FD86, GÅ89, Gro83c, HL85, IA86, Jør80c, Jør81b, Jør81a, KEK85, KHMBD86, KSBM84, KKG82, KH87b, LSV83, MK86, MBM^{+}89, MF86, PU88, PM82, Pat84, Rag84, RC83, RS88, SMR86, Tho84, Tya88, URC88, Var88a, Var89b, VKDS86, WC85, RSG87, Jør82c, Jør83, Lho87]. Water-quality [Som82]. water-quality-transport [VKDS86]. water-use [URC88]. waterhyacinth [LMR84]. waters [HPKA80, IUME85, KI86, KMRV89, RJBL88, Gro83a]. Watershed [LSSF89, Fed84, KMRV89, TD82]. watersheds [Mid84b]. wave [Kam81, Var87a]. Wayne [Gro85]. weather [OA82]. web [JG88]. webs [DeA88, Leg89b, Lim85]. weevil [Lon88b]. Weibull [NO84]. weight [Ber85, Van87]. weight-structure [Van87]. Welch [Jør82c, Jør83]. Werner [Gro83a]. West [Gro83a, Hal86a, Nih84b]. western [KA84, Jen82]. Westview [SV89b]. wet [Var87a]. wet-dry [Var87a]. wetland [CS85, SCD85, Svi89]. Wetlands [KH88, Jac87]. WG [Nih84a]. whale [RJBL88]. wheat [KWSM89, LC84, TCL82]. White [Gat87, BH84, EM84, EF82]. white-tailed [EM84]. Whittington [Lho87]. Whole [OSW85, Var88b]. whole-lake [Var88b]. Whole-plant [OSW85]. whose [Hal88b]. wide [MC86]. wild [Ush84c]. wildfire [MW80]. Wildlife [Ano83t, CT82, GFI84, Gra88a, SRM83, Svi88c, Lon87, Pac88]. Wiley [Ano87i, Col83, Gat88, Gro83a, Gro84b, Gue83, Hal86a, Jac87, Jam87, Jør80c, Jør86d, Jør87d, Koh89, Leg84b, Lon86, Nih84b, SB85, Sko86a, Sko88, Svi88a, Svi88c, Zuc89a]. Wiley-Interscience [Gro83a, Sko86a, Svi88c]. William [Baz86]. willow [KE84]. Wilson [Col83, Jam87]. wilting [SWH84]. wind [HS87, SF82]. windborne [Gro84b]. window [Car84]. Wisconsin [WC85]. Wispelaere [Saw83]. Wit [Jør81c, LW88]. Within [FWS^{+}81, HdJ89, KP89]. Within-tree [FWS^{+}81, KP89]. without [KT89]. wolf [Sto81]. wolf-ungulate [Sto81]. Wolves [CTJ81]. woodland [HGG^{+}82]. woodlands [SWPW85]. Woodward [ES85]. Woodwell [Jør86d]. word [OMA^{+}86]. Working [Nih84a]. workshop [Jør81a]. worldwide [Jør81b]. worth [YL87]. Wynne [Pal89].

x [Sko86a, Gra88a, Gro81, Gro84c, Gue83, Jam87, Sko86a, Svi88c, LU82a, Ush84d]. xi [Leg89a, Mej83, Sko86c]. xii [Gro83b, Jør86c, Lon86, Mau88, Sko86b, Som84, Str87a, Svi88b, Ush81b]. xiv [Saw83, Svi88c, Ush84c]. xv [LU82a, Svi88a, Ush81a, Zuc89a]. xvi [Jør86e, Sko86d]. xvii [Gra88b, Lon87]. xviii [Nih85]. xxi [Leg84b]. xxvi [SV89b].

Ya [Koh89]. Yaron [Sko86d]. years [KD83]. yeast [ZBBLCSTL86]. yellow [Jen82]. Yield [MS87b, HM88]. yield-pest [HM88]. Yodo [SMR86]. York [Gra88a, Gro81, Gro84c, Gro85, Hof84, Jør80a, Jør80c, Jør81d, Jør86c, Jør87b, Koh89, Lho87, LU82a, Mid84a, Mit87, Nih83b, Nih84d, Nih84a, Nih85, Saw83, Som84, Str84a, Str87a, Svi88b, Svi88c, Ush81b, Ush81a, Ush84a, Ush84b, Ush84c, Ush85, Wel88, Wul86]. York/Chichester [Svi88c]. York/London [Gro85, Ush84c]. York/Tokyo [Jør86c, Mit87, Nih84d, Str87a, Ush85, Wul86]. York/Toronto/Sydney/San [Ush84a]. Yugoslavia [OA82].

Zaire [WF88]. Zala [BK84]. Zea [WS80]. ZERO [Kir89]. Ziegler [Ush81b]. Zimmermann [WGSC84]. Zingales [Zuc89b]. Znamensky [Gro83c]. Zone [CMLB85, Day82, PLRV88, PU88, Sko86d]. Zooplankton [Sjö80, AV81, HP83, Mau82a, Mik86, RSKT88, Vol81, Vyh87].

References

Aagren:1980:PRT

[ÅA80] Göran I. Ågren and Björn Axelsson. Population respiration: a theoretical approach. Ecological Modelling, 11(1):39–54, October 1980. CODEN ECMODT. ISSN 0304-3800 (print),

              1872-7026 (electronic). URL http://www.sciencedirect.
              com/science/article/pii/0304380080900708.

Abdelkader:1980:GTS

[Abd80] Mostafa A. Abdelkader. A general two-species competition model with time-varying rates. Ecological Modelling, 10(1):31– 45, June 1980. CODEN ECMODT. ISSN 0304-3800 (print),

              1872-7026 (electronic). URL http://www.sciencedirect.
              com/science/article/pii/0304380080900770.

Adetiloye:1985:MMF

[Ade85] P. O. Adetiloye. A mathematical model for formulating intercrop proportions for intercropping systems’ design. Ecological Modelling, 27(1–2):81–93, March 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

              URL http://www.sciencedirect.com/science/article/
              pii/0304380085900262.

Adetiloye:1983:LEC

[AEO83] P. O. Adetiloye, F. O. C. Ezedinma, and B. N. Okigbo. A land equivalent coefficient (LEC) concept for the evaluation of competitive and productive interactions in simple to complex crop mixtures. Ecological Modelling, 19(1):27–39, May 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380083900686.

Ashikhmina:1985:MMD

[AFSK85] E. V. Ashikhmina, E. Ya. Frisman, E. I. Skaletskaya, and A. N. Kulikov. Mathematical model for dynamics of the number of pelt products from the local population of mantchurian squirrels. Ecological Modelling, 30(1–2):145–156, October 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900407.

Agnew:1982:SET

[Agn82] T. T. Agnew. Stability and exploitation in two-species discrete time population models with delay. Ecological Modelling, 15(3):235–249, April 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008290028X.

Aagren:1981:SEM

[Ågr81] Göran I. Ågren. Simple elimination models: Theoretical estimates of the effect of variations in elimination rates. Ecological Modelling, 12(4):281–295, May 1981. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380081900442.

Ahlert:1980:RNM

[AH80] Robert C. Ahlert and Shing-Fu Hsueh. A reactor network model of the Passaic River. Ecological Modelling, 10(1):47– 61, June 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900782.

Aizen:1989:FLL

[Aiz89] Marcelo A. Aizen. Fit of logspecies-logarea regression lines to nonequilibrium archipelagos: a simulation approach. Ecological Modelling, 47(3–4):265–273, September 15, 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900045.

Alekseev:1985:ESM

[AK85] V. V. Alekseev and A. N. Kornilovsky. Ecosystems stochasticity model. Ecological Modelling, 28(3):217–229, June 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900845.

Antonelli:1988:MDD

[AK88] P. L. Antonelli and N. D. Kazarinoff. Modelling densitydependent aggregation and reproduction in certain terrestrial and marine ecosystems: a comparative study. Ecological Modelling, 41(3–4):219–227, June 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900294.

Alekseev:1984:PMA

[AKL84] V. V. Alekseev, A. O. Kokorin, and M. Ya. Lyamin. Physical model of the aquatic ecosystem. Ecological Modelling, 25(1–3):31–46, October 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084900310.

Alekseev:1982:MNS

[Ale82] Vyacheslav Viktorovich Alekseev. Model of the number of species in an ecosystem in dependence on geographical latitude. Ecological Modelling, 17(2):107–112, October 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900473.

Ahlert:1981:SAT

[AM81] R. C. Ahlert and B. M. Mehta. Stochastic analyses and transfer functions for flows of the upper Delaware River.

Ecological Modelling, 14(1–2):59–78, November 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900144.

Arp:1987:PBMa

[AM87] Paul A. Arp and Timothy P. McGrath. A parameterbased method for modelling biomass accumulations in forest stands: Theory. Ecological Modelling, 36(1–2):29–48, April 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900561.

Arp:1987:PBMb

[AMB87] Paul A. Arp, Timothy P. McGrath, and Judith A. Beck. A parameter-based methods for modelling biomass accumulations in forest stands: an application. Ecological Modelling, 36(1–2):49–64, April 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900573.

Akai:1983:NMI

[AMO83] Dairo Akai, Osamu Miki, and Shinichiro Ohgaki. Nitrification model with an inhibitory effect of sea water. Ecological Modelling, 19(3):189–198, July 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900534.

Ayars:1981:MST

[AMS81] James E. Ayars, David B. McWhorter, and Gaylord V. Skogerboe. Modeling salt transport in irrigated soils. Ecological Modelling, 11(4):265–290, February 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380081900624.

Anderson:1983:LTM

[And83] Max G. Anderson. A long-term model for plant disease hostpathogen simulation. Ecological Modelling, 20(2–3):201–222, November 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900078.

Anonymous:1980:Aa

[Ano80a] Anonymous. Announcement. Ecological Modelling, 10(1):65, June 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900800.

Anonymous:1980:Ab

[Ano80b] Anonymous. Announcement. Ecological Modelling, 11(3):231, December 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900861.

Anonymous:1980:AI

[Ano80c] Anonymous. Author index. Ecological Modelling, 10(3–4):265, August 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900630.

Anonymous:1980:EBa

[Ano80d] Anonymous. Editorial Board. Ecological Modelling, 10(1): ifc, June 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900745.

Anonymous:1980:EBb

[Ano80e] Anonymous. Editorial Board. Ecological Modelling, 11(1):iii, October 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900678.

Anonymous:1980:E

[Ano80f] Anonymous. Errata. Ecological Modelling, 11(1):71, October 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900733.

Anonymous:1980:ISE

[Ano80g] Anonymous. International symposium on energy and ecological modelling. Ecological Modelling, 11(2):155, November 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380080900927.

Anonymous:1980:PJa

[Ano80h] Anonymous. Pages 1–65 (June 1980). Ecological Modelling, 10 (1):??, June 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1980:PO

[Ano80i] Anonymous. Pages 1–78 (October 1980). Ecological Modelling, 11(1):??, October 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1980:PA

[Ano80j] Anonymous. Pages 139–265 (August 1980). Ecological Modelling, 10(3–4):??, August 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1980:PD

[Ano80k] Anonymous. Pages 157–231 (December 1980). Ecological Modelling, 11(3):??, December 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1980:PJb

[Ano80l] Anonymous. Pages 67–137 (July 1980). Ecological Modelling, 10(2):??, July 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1980:PNb

[Ano80m] Anonymous. Pages 79–155 (November 1980). Ecological Modelling, 11(2):??, November 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1980:PNa

[Ano80n] Anonymous. Publisher’s note. Ecological Modelling, 10(3–4): 139, August 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900563.

Anonymous:1980:PNc

[Ano80o] Anonymous. Publisher’s note. Ecological Modelling, 11(2):155, November 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900939.

Anonymous:1981:Ab

[Ano81a] Anonymous. Announcement. Ecological Modelling, 13(4):315, September 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900363.

Anonymous:1981:Ac

[Ano81b] Anonymous. Announcement. Ecological Modelling, 14(1– 2):147, November 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380081900193.

Anonymous:1981:Aa

[Ano81c] Anonymous. Announcements. Ecological Modelling, 12(4):301– 302, May 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900466.

Anonymous:1981:AIa

[Ano81d] Anonymous. Author index. Ecological Modelling, 11(4): 299–300, February 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380081900673.

Anonymous:1981:AIb

[Ano81e] Anonymous. Author index. Ecological Modelling, 12(4):303– 304, May 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900478.

Anonymous:1981:AIc

[Ano81f] Anonymous. Author index. Ecological Modelling, 13(4): 317–318, September 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380081900375.

Anonymous:1981:C

[Ano81g] Anonymous. Corrigendum. Ecological Modelling, 13(3):221, August 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900533.

Anonymous:1981:EBa

[Ano81h] Anonymous. Editorial Board. Ecological Modelling, 12(1–2): iii, March 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008190020X.

Anonymous:1981:EBb

[Ano81i] Anonymous. Editorial Board. Ecological Modelling, 13(1–2): iii, June 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008190003X.

Anonymous:1981:EBc

[Ano81j] Anonymous. Editorial Board. Ecological Modelling, 14(1–2):iii, November 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900107.

Anonymous:1981:PJ

[Ano81k] Anonymous. Pages 1–129 (June 1981). Ecological Modelling, 13(1–2):??, June 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1981:PMa

[Ano81l] Anonymous. Pages 1–134 (March 1981). Ecological Modelling, 12(1–2):??, March 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1981:PN

[Ano81m] Anonymous. Pages 1–147 (November 1981). Ecological Modelling, 14(1–2):??, November 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1981:PAb

[Ano81n] Anonymous. Pages 131–221 (August 1981). Ecological Modelling, 13(3):??, August 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1981:PAa

[Ano81o] Anonymous. Pages 135—211 (April 1981). Ecological Modelling, 12(3):??, April 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1981:PMb

[Ano81p] Anonymous. Pages 213–304 (May 1981). Ecological Modelling, 12(4):??, May 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1981:PS

[Ano81q] Anonymous. Pages 223–318 (September 1981). Ecological Modelling, 13(4):??, September 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1981:PF

[Ano81r] Anonymous. Pages 233–300 (February 1981). Ecological Modelling, 11(4):??, February 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1982:Aa

[Ano82a] Anonymous. Announcement. Ecological Modelling, 15(2):189, March 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900631.

Anonymous:1982:Ab

[Ano82b] Anonymous. Announcement. Ecological Modelling, 15(4):367, June 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900904.

Anonymous:1982:Ac

[Ano82c] Anonymous. Announcement. Ecological Modelling, 17(3– 4):322, December 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900394.

Anonymous:1982:AIa

[Ano82d] Anonymous. Author index. Ecological Modelling, 14(3–4): 301–302, January 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900254.

Anonymous:1982:AIb

[Ano82e] Anonymous. Author index. Ecological Modelling, 15(4):369– 370, June 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900916.

Anonymous:1982:AIc

[Ano82f] Anonymous. Author index. Ecological Modelling, 16(2–4):299– 300, August 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900138.

Anonymous:1982:AId

[Ano82g] Anonymous. Author index. Ecological Modelling, 17(3– 4):323–324, December 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900400.

Anonymous:1982:MAB

[Ano82h] Anonymous. Management and analysis of biological populations: Bean-San Goh. Elsevier, Amsterdam, 1980, 290 pp., illustrated, US $75.50/Dfl. 155, ISBN 0-444-41793-1. Ecological Modelling, 15(2):186–187, March 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008290062X.

Anonymous:1982:EBa

[Ano82i] Anonymous. Editorial Board. Ecological Modelling, 15(1):iii, February 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900643.

Anonymous:1982:EBb

[Ano82j] Anonymous. Editorial Board. Ecological Modelling, 16(1): iii, July 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900709.

Anonymous:1982:EBc

[Ano82k] Anonymous. Editorial Board. Ecological Modelling, 17(1):iii, September 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900825.

Anonymous:1982:PS

[Ano82l] Anonymous. Pages 1–51 (September 1982). Ecological Modelling, 17(1):??, September 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1982:PJb

[Ano82m] Anonymous. Pages 1–80 (July 1982). Ecological Modelling, 16 (1):??, July 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1982:PF

[Ano82n] Anonymous. Pages 1–86 (February 1982). Ecological Modelling, 15(1):??, February 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1982:PD

[Ano82o] Anonymous. Pages 183–324 (December 1982). Ecological Modelling, 17(3–4):??, December 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1982:PAa

[Ano82p] Anonymous. Pages 191–285 (April 1982). Ecological Modelling, 15(3):??, April 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1982:PJa

[Ano82q] Anonymous. Pages 287–370 (June 1982). Ecological Modelling, 15(4):??, June 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1982:PAb

[Ano82r] Anonymous. Pages 81–300 (August 1982). Ecological Modelling, 16(2–4):??, August 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1982:PM

[Ano82s] Anonymous. Pages 87–189 (March 1982). Ecological Modelling, 15(2):??, March 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:AIa

[Ano83a] Anonymous. Author index. Ecological Modelling, 18(3–4):309– 310, April 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900224.

Anonymous:1983:AIb

[Ano83b] Anonymous. Author index. Ecological Modelling, 19(4): 309–310, September 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380083900467.

Anonymous:1983:AIc

[Ano83c] Anonymous. Author index. Ecological Modelling, 20(4): 323–324, December 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380083900406.

Anonymous:1983:EBa

[Ano83d] Anonymous. Editorial Board. Ecological Modelling, 18(1):iii, January 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900716.

Anonymous:1983:EBb

[Ano83e] Anonymous. Editorial Board. Ecological Modelling, 19(1): iii, May 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900662.

Anonymous:1983:EBc

[Ano83f] Anonymous. Editorial Board. Ecological Modelling, 20(1):iii, October 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900273.

Anonymous:1983:E

[Ano83g] Anonymous. Errata. Ecological Modelling, 19(3):235–237, July 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900650.

Anonymous:1983:EEM

[Ano83h] Anonymous. Erratum: Ecol. Modelling: Lassey, K. R., 1982. Conceptually simple mathematical models of filtration (and exchange) processes. 18: 1–26. Ecological Modelling, 19(1):71, May 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900704. See [Las83].

Anonymous:1983:FIC

[Ano83i] Anonymous. Fourth international conference on state-of-the-art in ecological modelling. Ecological Modelling, 20(1):83–84, October 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900339.

Anonymous:1983:PM

[Ano83j] Anonymous. Pages 1–71 (May 1983). Ecological Modelling, 19 (1):??, May 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PJa

[Ano83k] Anonymous. Pages 1–72 (January 1983). Ecological Modelling, 18(1):??, January 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PO

[Ano83l] Anonymous. Pages 1–84 (October 1983). Ecological Modelling, 20(1):??, October 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PA

[Ano83m] Anonymous. Pages 155–310 (April 1983). Ecological Modelling, 18(3–4):??, April 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PJc

[Ano83n] Anonymous. Pages 163–237 (July 1983). Ecological Modelling, 19(3):??, July 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PS

[Ano83o] Anonymous. Pages 239–310 (September 1983). Ecological Modelling, 19(4):??, September 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PD

[Ano83p] Anonymous. Pages 243–324 (December 1983). Ecological Modelling, 20(4):??, December 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PF

[Ano83q] Anonymous. Pages 73–153 (February 1983). Ecological Modelling, 18(2):??, February 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PJb

[Ano83r] Anonymous. Pages 73–161 (June 1983). Ecological Modelling, 19(2):??, June 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:PN

[Ano83s] Anonymous. Pages 85–241 (November 1983). Ecological Modelling, 20(2–3):??, November 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1983:WMH

[Ano83t] Anonymous. Wildlife 2000: Modeling habitat relationships of terrestrial vertebrates. Ecological Modelling, 20(1):84, October 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900340.

Anonymous:1984:Aa

[Ano84a] Anonymous. Announcement. Ecological Modelling, 23(3):275, June 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901108.

Anonymous:1984:Ab

[Ano84b] Anonymous. Announcement. Ecological Modelling, 23(4):349, July 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901327.

Anonymous:1984:Ac

[Ano84c] Anonymous. Announcement. Ecological Modelling, 26(3– 4):327, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438008490084X.

Anonymous:1984:AIa

[Ano84d] Anonymous. Author index. Ecological Modelling, 21(4):339– 340, March 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008490067X.

Anonymous:1984:AIb

[Ano84e] Anonymous. Author index. Ecological Modelling, 22(1–4):341– 342, April 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008490019X.

Anonymous:1984:AIc

[Ano84f] Anonymous. Author index. Ecological Modelling, 23(4):351– 352, July 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901339.

Anonymous:1984:AId

[Ano84g] Anonymous. Author index. Ecological Modelling, 24(3– 4):309–310, September 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084900498.

Anonymous:1984:AIe

[Ano84h] Anonymous. Author index. Ecological Modelling, 25(4): 169, 171–206, November 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084901121.

Anonymous:1984:AIf

[Ano84i] Anonymous. Author index. Ecological Modelling, 26(3–4): 328–330, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084900851.

Anonymous:1984:EBa

[Ano84j] Anonymous. Editorial Board. Ecological Modelling, 21(1–2):iii, January 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900206.

Anonymous:1984:EBb

[Ano84k] Anonymous. Editorial Board. Ecological Modelling, 22(1–4): iii, April 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900012.

Anonymous:1984:EBc

[Ano84l] Anonymous. Editorial Board. Ecological Modelling, 23(1–2): iii, May 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901145.

Anonymous:1984:EBd

[Ano84m] Anonymous. Editorial Board. Ecological Modelling, 24(1–2): iii, August 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900504.

Anonymous:1984:EBe

[Ano84n] Anonymous. Editorial Board. Ecological Modelling, 25(1–3):v, October 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900292.

Anonymous:1984:EBf

[Ano84o] Anonymous. Editorial Board. Ecological Modelling, 26(1–2):iii, December 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900863.

Anonymous:1984:E

[Ano84p] Anonymous. Erratum. Ecological Modelling, 26(3–4):325, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900838.

Anonymous:1984:PJa

[Ano84q] Anonymous. Pages 1–147 (January 1984). Ecological Modelling, 21(1–2):??, January 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:PA

[Ano84r] Anonymous. Pages 1–157 (August 1984). Ecological Modelling, 24(1–2):??, August 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:POa

[Ano84s] Anonymous. Pages 1–165 (October 1984). Ecological Modelling, 25(1–3):??, October 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:PF

[Ano84t] Anonymous. Pages 149–207 (February 1984). Ecological Modelling, 21(3):??, February 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:PD

[Ano84u] Anonymous. Pages 155–330 (15 December 1984). Ecological Modelling, 26(3–4):??, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:PS

[Ano84v] Anonymous. Pages 159–310 (September 1984). Ecological Modelling, 24(3–4):??, September 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:PN

[Ano84w] Anonymous. Pages 167–254 (November 1984). Ecological Modelling, 25(4):??, November 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:PJb

[Ano84x] Anonymous. Pages 191–275 (June 1984). Ecological Modelling, 23(3):??, June 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:PJc

[Ano84y] Anonymous. Pages 277–352 (July 1984). Ecological Modelling, 23(4):??, July 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:POb

[Ano84z] Anonymous. Publication overview. Ecological Modelling, 25 (4):167, November 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438008490111X.

Anonymous:1984:S

[Ano84-27] Anonymous. (SISY). Ecological Modelling, 26(1–2):??, December 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1984:SI

[Ano84-28] Anonymous. Subject index. Ecological Modelling, 25(4): 207, 209–254, November 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084901133.

Anonymous:1985:Aa

[Ano85a] Anonymous. Announcements. Ecological Modelling, 27(1–2): 161, March 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900316.

Anonymous:1985:Ab

[Ano85b] Anonymous. Announcements. Ecological Modelling, 30(3– 4):319, December 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380085900754.

Anonymous:1985:AIa

[Ano85c] Anonymous. Author index. Ecological Modelling, 27(3–4):315– 316, April 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900110.

Anonymous:1985:AIb

[Ano85d] Anonymous. Author index. Ecological Modelling, 28(4):323– 324, August 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900808.

Anonymous:1985:AIc

[Ano85e] Anonymous. Author index. Ecological Modelling, 29(1– 4):423–424, September 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380085900651.

Anonymous:1985:AId

[Ano85f] Anonymous. Author index. Ecological Modelling, 30(3– 4):321–322, December 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380085900766.

Anonymous:1985:CP

[Ano85g] Anonymous. Call for papers. Ecological Modelling, 28(3): 241, June 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900912.

Anonymous:1985:EBa

[Ano85h] Anonymous. Editorial Board. Ecological Modelling, 27(1–2): iii, March 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900201.

Anonymous:1985:EBb

[Ano85i] Anonymous. Editorial Board. Ecological Modelling, 28(1–2): iii, May 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900122.

Anonymous:1985:EBc

[Ano85j] Anonymous. Editorial Board. Ecological Modelling, 29(1–4):iii, September 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900420.

Anonymous:1985:EBd

[Ano85k] Anonymous. Editorial Board. Ecological Modelling, 30(1–2):iii, October 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900328.

Anonymous:1985:E

[Ano85l] Anonymous. Errata. Ecological Modelling, 30(3–4):317–318, December 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900742.

Anonymous:1985:LR

[Ano85m] Anonymous. List of reviewer. Ecological Modelling, 29(1– 4):421, September 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438008590064X.

Anonymous:1985:MRC

[Ano85n] Anonymous. Man’s role in changing the global environment. Ecological Modelling, 27(3–4):313, April 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900109.

Anonymous:1985:PO

[Ano85o] Anonymous. Pages 1–161 (October 1985). Ecological Modelling, 30(1–2):??, October 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1985:PMb

[Ano85p] Anonymous. Pages 1–164 (May 1985). Ecological Modelling, 28(1–2):??, May 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1985:PMa

[Ano85q] Anonymous. Pages 1–168 (March 1985). Ecological Modelling, 27(1–2):??, March 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1985:PS

[Ano85r] Anonymous. Pages 1–424 (September 1985). Ecological Modelling, 29(1–4):??, September 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1985:PD

[Ano85s] Anonymous. Pages 163–322 (December 1985). Ecological Modelling, 30(3–4):??, December 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1985:PJ

[Ano85t] Anonymous. Pages 165–241 (June 1985). Ecological Modelling, 28(3):??, June 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1985:PAa

[Ano85u] Anonymous. Pages 169–316 (April 1985). Ecological Modelling, 27(3–4):??, April 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1985:PAb

[Ano85v] Anonymous. Pages 243–324 (August 1985). Ecological Modelling, 28(4):??, August 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1985:PN

[Ano85w] Anonymous. Publisher’s note. Ecological Modelling, 30(1–2): 1, October 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008590033X.

Anonymous:1986:A

[Ano86a] Anonymous. Announcement. Ecological Modelling, 34(1– 2):141, November 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380086900888.

Anonymous:1986:AIa

[Ano86b] Anonymous. Author index. Ecological Modelling, 31(1–4):371– 373, May 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900761.

Anonymous:1986:AIb

[Ano86c] Anonymous. Author index. Ecological Modelling, 32(4):323– 325, July 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008690102X.

Anonymous:1986:AIc

[Ano86d] Anonymous. Author index. Ecological Modelling, 34(3– 4):299–300, December 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380086900104.

Anonymous:1986:AIV

[Ano86e] Anonymous. Author index (vol. 33). Ecological Modelling, 33(2–4):341–342, October 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900487.

Anonymous:1986:EBa

[Ano86f] Anonymous. Editorial Board. Ecological Modelling, 31(1–4): iii, May 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900499.

Anonymous:1986:EBb

[Ano86g] Anonymous. Editorial Board. Ecological Modelling, 32(1–3): iii, June 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900116.

Anonymous:1986:EBc

[Ano86h] Anonymous. Editorial Board. Ecological Modelling, 33(1):iii, September 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086901031.

Anonymous:1986:EBd

[Ano86i] Anonymous. Editorial Board. Ecological Modelling, 34(1–2):iii, November 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900773.

Anonymous:1986:LC

[Ano86j] Anonymous. List of contributors. Ecological Modelling, 32 (1–3):241, June 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380086900311.

Anonymous:1986:PN

[Ano86k] Anonymous. Pages 1–141 (November 1986). Ecological Modelling, 34(1–2):??, November 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1986:PS

[Ano86l] Anonymous. Pages 1–80 (September 1986). Ecological Modelling, 33(1):??, September 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1986:PD

[Ano86m] Anonymous. Pages 143–300 (December 1986). Ecological Modelling, 34(3–4):??, December 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1986:PJ

[Ano86n] Anonymous. Pages 243–325 (July 1986). Ecological Modelling, 32(4):??, July 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1987:Ab

[Ano87a] Anonymous. Announcement. Ecological Modelling, 37(3–4): 328, July 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008790038X.

Anonymous:1987:Aa

[Ano87b] Anonymous. Announcements. Ecological Modelling, 35(3– 4):315–317, March 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380087901207.

Anonymous:1987:AIa

[Ano87c] Anonymous. Author index. Ecological Modelling, 35(3–4):319– 320, March 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087901219.

Anonymous:1987:AIb

[Ano87d] Anonymous. Author index. Ecological Modelling, 36(3–4):323– 324, May 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900780.

Anonymous:1987:AIc

[Ano87e] Anonymous. Author index. Ecological Modelling, 37(3–4):329– 330, July 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900391.

Anonymous:1987:AId

[Ano87f] Anonymous. Author index. Ecological Modelling, 38(3–4): 319–320, October 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380087901086.

Anonymous:1987:AIe

[Ano87g] Anonymous. Author index. Ecological Modelling, 39(3– 4):325–326, December 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438008790007X.

Anonymous:1987:MEP

[Ano87h] Anonymous. Modelling and explaining the phosphorus dynamics of Lake Balaton, 1976–1979: A. Voinov. International Institute for Applied Systems Analysis, Laxenburg, Austria, 1985. 59 pp., US$70.00. ISBN 3-7045-0075-5. Ecological Modelling, 35(3–4):312–313, March 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087901190.

Anonymous:1987:EOC

[Ano87i] Anonymous. The electronic oracle: Computer models and social decisions: D. H. Meadows and J. M. Robinson. Wiley, Chichester, Great Britain, 1985. 438 pp., £24.95. ISBN 0-471-90558-5. Ecological Modelling, 36(3–4):320–321, May 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900779.

Anonymous:1987:EBa

[Ano87j] Anonymous. Editorial Board. Ecological Modelling, 35(1–2):iii, February 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900871.

Anonymous:1987:EBb

[Ano87k] Anonymous. Editorial Board. Ecological Modelling, 36(1–2): iii, April 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900524.

Anonymous:1987:EBc

[Ano87l] Anonymous. Editorial Board. Ecological Modelling, 37(1–2): iii, June 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900792.

Anonymous:1987:EBd

[Ano87m] Anonymous. Editorial Board. Ecological Modelling, 38(1–2):iii, September 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900408.

Anonymous:1987:EBe

[Ano87n] Anonymous. Editorial Board. Ecological Modelling, 39(1–2):iii, November 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900081.

Anonymous:1987:E

[Ano87o] Anonymous. Errata. Ecological Modelling, 37(3–4):327, July 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900378.

Anonymous:1987:LR

[Ano87p] Anonymous. List of reviewers. Ecological Modelling, 36(1–2): 153, April 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900640.

Anonymous:1987:PF

[Ano87q] Anonymous. Pages 1–156 (February 1987). Ecological Modelling, 35(1–2):??, February 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1987:PJ

[Ano87r] Anonymous. Pages 131–330 (July 1987). Ecological Modelling, 37(3–4):??, July 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1987:PMb

[Ano87s] Anonymous. Pages 155–324 (May 1987). Ecological Modelling, 36(3–4):??, May 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1987:PMa

[Ano87t] Anonymous. Pages 157–320 (March 1987). Ecological Modelling, 35(3–4):??, March 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1987:PO

[Ano87u] Anonymous. Pages 191–320 (October 1987). Ecological Modelling, 38(3–4):??, October 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1987:PD

[Ano87v] Anonymous. Pages 201–326 (December 1987). Ecological Modelling, 39(3–4):??, December 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:Aa

[Ano88a] Anonymous. Announcement. Ecological Modelling, 42(1):87, July 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900968.

Anonymous:1988:Ab

[Ano88b] Anonymous. Announcements. Ecological Modelling, 42(3– 4):309, September 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380088900658.

Anonymous:1988:AIa

[Ano88c] Anonymous. Author index. Ecological Modelling, 40(3–4):311– 312, March 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900221.

Anonymous:1988:AIb

[Ano88d] Anonymous. Author index. Ecological Modelling, 41(3–4):329– 331, June 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900373.

Anonymous:1988:AIc

[Ano88e] Anonymous. Author index. Ecological Modelling, 42(3– 4):311–312, September 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438008890066X.

Anonymous:1988:AId

[Ano88f] Anonymous. Author index. Ecological Modelling, 43(3– 4):321–322, November 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380088900130.

Anonymous:1988:CEMa

[Ano88g] Anonymous. Contents of Ecological Modelling, vol. 40. Ecological Modelling, 40(3–4):313–314, March 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900233.

Anonymous:1988:CEMb

[Ano88h] Anonymous. Contents of Ecological Modelling, vol. 41. Ecological Modelling, 41(3–4):331–332, June 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900385.

Anonymous:1988:CEMc

[Ano88i] Anonymous. Contents of ecological modelling, vol. 42. Ecological Modelling, 42(3–4):313–314, September 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900671.

Anonymous:1988:CEMd

[Ano88j] Anonymous. Contents of ecological modelling, vol. 43. Ecological Modelling, 43(3–4):323–324, November 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900142.

Anonymous:1988:EBa

[Ano88k] Anonymous. Editorial Board. Ecological Modelling, 40(1):iii, January 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008890097X.

Anonymous:1988:EBb

[Ano88l] Anonymous. Editorial Board. Ecological Modelling, 41(1–2): iii, April 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900397.

Anonymous:1988:EBc

[Ano88m] Anonymous. Editorial Board. Ecological Modelling, 42(1): iii, July 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900889.

Anonymous:1988:EBd

[Ano88n] Anonymous. Editorial Board. Ecological Modelling, 43(1–2):iii, October 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900683.

Anonymous:1988:EBe

[Ano88o] Anonymous. Editorial Board. Ecological Modelling, 44(1–2):iii, December 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900798.

Anonymous:1988:PO

[Ano88p] Anonymous. Pages 1–136 (October 1988). Ecological Modelling, 43(1–2):??, October 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PAa

[Ano88q] Anonymous. Pages 1–156 (April 1988). Ecological Modelling, 41(1–2):??, April 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PD

[Ano88r] Anonymous. Pages 1–164 (December 1988). Ecological Modelling, 44(1–2):??, December 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PJa

[Ano88s] Anonymous. Pages 1–84 (January 1988). Ecological Modelling, 40(1):??, January 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PJc

[Ano88t] Anonymous. Pages 1–87 (July 1988). Ecological Modelling, 42 (1):??, July 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PN

[Ano88u] Anonymous. Pages 137–324 (November 1988). Ecological Modelling, 43(3–4):??, November 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PM

[Ano88v] Anonymous. Pages 155–314 (March 1988). Ecological Modelling, 40(3–4):??, March 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PJb

[Ano88w] Anonymous. Pages 157–332 (June 1988). Ecological Modelling, 41(3–4):??, June 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PS

[Ano88x] Anonymous. Pages 165–314 (September 1988). Ecological Modelling, 42(3–4):??, September 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PF

[Ano88y] Anonymous. Pages 85–153 (February 1988). Ecological Modelling, 40(2):??, February 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1988:PAb

[Ano88z] Anonymous. Pages 89–164 (August 1988). Ecological Modelling, 42(2):??, August 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:Aa

[Ano89a] Anonymous. Announcement. Ecological Modelling, 45(1):73, February 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089901038.

Anonymous:1989:Ab

[Ano89b] Anonymous. Announcements. Ecological Modelling, 45(2):163, April 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900951.

Anonymous:1989:AIa

[Ano89c] Anonymous. Author index. Ecological Modelling, 44(3–4): 317–318, January 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380089900379.

Anonymous:1989:AIb

[Ano89d] Anonymous. Author index. Ecological Modelling, 45(4):317– 318, June 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900781.

Anonymous:1989:AIc

[Ano89e] Anonymous. Author index. Ecological Modelling, 46(3–4):311– 312, August 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900252.

Anonymous:1989:AId

[Ano89f] Anonymous. Author index. Ecological Modelling, 47(3–4): 319–320, September 15, 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380089900124.

Anonymous:1989:AIe

[Ano89g] Anonymous. Author index. Ecological Modelling, 48(3– 4):315–316, November 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380089900549.

Anonymous:1989:CEMa

[Ano89h] Anonymous. Contents of Ecological Modelling, vol. 44. Ecological Modelling, 44(3–4):319–320, January 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900380.

Anonymous:1989:CEMb

[Ano89i] Anonymous. Contents of Ecological Modelling, vol. 45. Ecological Modelling, 45(4):319–320, June 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900793.

Anonymous:1989:CEMc

[Ano89j] Anonymous. Contents of Ecological Modelling, vol. 46. Ecological Modelling, 46(3–4):313–314, August 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900264.

Anonymous:1989:CEMe

[Ano89k] Anonymous. Contents of Ecological Modelling, vol. 48. Ecological Modelling, 48(3–4):317, November 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900550.

Anonymous:1989:CEMd

[Ano89l] Anonymous. Contents of ecological modelling, vol. 47. Ecological Modelling, 47(3–4):321–322, September 15, 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900136.

Anonymous:1989:EBa

[Ano89m] Anonymous. Editorial Board. Ecological Modelling, 45(1):iii, February 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900963.

Anonymous:1989:EBb

[Ano89n] Anonymous. Editorial Board. Ecological Modelling, 46(1–2): iii, July 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900653.

Anonymous:1989:EBc

[Ano89o] Anonymous. Editorial Board. Ecological Modelling, 47(1–2):iii, September 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008990104X.

Anonymous:1989:EBd

[Ano89p] Anonymous. Editorial Board. Ecological Modelling, 48(1–2):iii, October 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900562.

Anonymous:1989:EBe

[Ano89q] Anonymous. Editorial Board. Ecological Modelling, 49(1–2):iii, December 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900392.

Anonymous:1989:NIC

[Ano89r] Anonymous. NERC/IMA conference on computer modelling in the environmental sciences. Ecological Modelling, 47(3–4):317, September 15, 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900112.

Anonymous:1989:PJc

[Ano89s] Anonymous. Pages 1–133 (July 1989). Ecological Modelling, 46(1–2):??, July 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PD

[Ano89t] Anonymous. Pages 1–152 (December 1989). Ecological Modelling, 49(1–2):??, December 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PO

[Ano89u] Anonymous. Pages 1–157 (October 1989). Ecological Modelling, 48(1–2):??, October 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PF

[Ano89v] Anonymous. Pages 1–80 (February 1989). Ecological Modelling, 45(1):??, February 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PAb

[Ano89w] Anonymous. Pages 135–314 (August 1989). Ecological Modelling, 46(3–4):??, August 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PN

[Ano89x] Anonymous. Pages 159–317 (November 1989). Ecological Modelling, 48(3–4):??, November 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PM

[Ano89y] Anonymous. Pages 165–242 (May 1989). Ecological Modelling, 45(3):??, May 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PJa

[Ano89z] Anonymous. Pages 165–320 (January 1989). Ecological Modelling, 44(3–4):??, January 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PS

[Ano89-27] Anonymous. Pages 199–322 (15 September 1989). Ecological Modelling, 47(3–4):??, September 15, 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PJb

[Ano89-28] Anonymous. Pages 243–320 (June 1989). Ecological Modelling, 45(4):??, June 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1989:PAa

[Ano89-29] Anonymous. Pages 81–163 (April 1989). Ecological Modelling, 45(2):??, April 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Aoki:1987:EBL

[Aok87] Ichiro Aoki. Entropy balance in Lake Biwa. Ecological Modelling, 37(3–4):235–248, July 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900275.

Aoki:1988:ELE

[Aok88] Ichiro Aoki. Entropy laws in ecological networks at steady state. Ecological Modelling, 42(3–4):289–303, September 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900622.

Aoki:1989:HSL

[Aok89] Ichiro Aoki. Holological study of lakes from an entropy viewpoint — Lake Mendota. Ecological Modelling, 45(2):81–93, April 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900859.

Arp:1983:VEA

[AR83] Paul A. Arp and Seepersaad Ramnarine. Verifying the effect of acid precipitation on soil leachates: a comparison between published records and model predictions. Ecological Modelling, 19(2):119–138, June 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008390025X.

Abdel-Razik:1989:MPO

[AR89] M. Abdel-Razik. A model of the productivity of olive trees under optional water and nutrient supply in desert conditions. Ecological Modelling, 45(3):179–204, May 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900811.

Arp:1983:MEA

[Arp83] P. A. Arp. Modelling the effects of acid precipitation on soil leachates: a simple approach. Ecological Modelling, 19(2):105– 117, June 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900248.

Alexander:1983:SRO

[AT83] William C. Alexander and Robert J. Taylor. Sex ratio and optimal harvest of canvasback ducks, a model. Ecological Modelling, 19(4):285–298, September 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900443.

Atkinson:1987:NPA

[Atk87] C. Allen Atkinson. A nonlinear programming approach to the analysis of perturbed marine ecosystems under model parameter uncertainty. Ecological Modelling, 35(1–2):1–28, February 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900883.

Arnold:1981:NBZ

[AV81] E. M. Arnold and D. A. Voss. Numerical behavior of a zooplankton, phytoplankton and phosphorus system. Ecological Modelling, 13(3):183–193, August 1981. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008190051X.

Alig:1985:PRA

[AW85] Ralph J. Alig and James G. Wyant. Projecting regional area changes in forestland cover in the U.S.A. Ecological Modelling, 29(1–4):27–34, September 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900456.

Bacon:1985:SCR

[Bac85] P. J. Bacon. Sensitivity coefficients — a reply to huson. Ecological Modelling, 27(1–2):153–160, March 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900304.

Bahr:1982:FTI

[Bah82] L. M. Bahr. Functional taxonomy: an immodest proposal. Ecological Modelling, 15(3):211–233, April 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900278.

Baillian:1986:PAN

[Bai86] Li Baillian. Pansystems analysis: a new approach to ecosystem modelling. Ecological Modelling, 32(1–3):227–236, June 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900293.

Barclay:1980:MSI

[Bar80] Hugh J. Barclay. Models for the sterile insect release method with the concurrent release of pesticides. Ecological Modelling, 11(3):167–177, December 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900824.

Barclay:1982:SRM

[Bar82] Hugh J. Barclay. The sterile release method with unequal male competitive ability. Ecological Modelling, 15(3):251–263, April 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900291.

Barclay:1987:MSI

[Bar87] Hugh J. Barclay. Models of sterile insect releases for populations under attack by parasitoids. Ecological Modelling, 36(3–4):155–169, May 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380087900652.

Barreto:1989:PLC

[Bar89] Luís Soares Barreto. The ‘3/2 power law’: a comment on the specific constancy of K. Ecological Modelling, 45(3):237– 242, May 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900847.

Bax:1985:AMU

[Bax85] Nicholas J. Bax. Application of multi- and univariate techniques of sensitivity analysis to SKEBUB, a biomass-based fisheries ecosystem model, parameterized to Georges Bank. Ecological Modelling, 29(1–4):353–382, September 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900614.

Bazykin:1986:AES

[Baz86] A. D. Bazykin. Analysis of ecological systems: State-of-the-art in ecological modelling: William K. Lauenroth, Gaylord V. Skogerboe and Marshall Flug (Editors). Developments in Environmental Modelling, 5. Elsevier, Amsterdam, 1983. 992 pp., Dfl. 320.00. ISBN 0-444-42179-3. Ecological Modelling, 34(1–2):137–139, November 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900864.

Bartell:1988:PUB

[BBC88] Steven M. Bartell, Antionette L. Brenkert, and Stephen R. Carpenter. Parameter uncertainty and the behavior of a sizedependent plankton model. Ecological Modelling, 40(2):85–95, February 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088901044.

Browder:1985:PMR

[BBD85] Joan A. Browder, Henry A. Bartley, and Kevin S. Davis. A probabilistic model of the relationship between marshlandwater interface and marsh disintegration. Ecological Modelling, 29(1–4):245–260, September 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900559.

Bazykin:1981:IPS

[BBDK81] A. D. Bazykin, F. S. Berezovskaya, G. A. Denisov, and Yu. A. Kuznetzov. The influence of predator saturation effect and competition among predators on predator–prey system dynamics. Ecological Modelling, 14(1–2):39–57, November 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900132.

Beddington:1982:HPP

[BC82] J. R. Beddington and J. G. Cooke. Harvesting from a preypredator complex. Ecological Modelling, 14(3–4):155–177, January 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900163.

Buranathanitt:1982:SEL

[BCJ82] Tavicha Buranathanitt, David J. Cockrell, and Peter H. John. Some effects of Langmuir circulation on the quality of water resource systems. Ecological Modelling, 15(1):49–74, February 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900679.

Bartell:1988:AER

[BCOG88] S. M. Bartell, W. G. Cale, R. V. O’Neill, and R. H. Gardner. Aggregation error: Research objectives and relevant model structure. Ecological Modelling, 41(3–4):157–168, June 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900245.

Bertonati:1987:EMV

[BDLP87] Mauro Bertonati, Camillo Dejak, Ileana Mazzei Lalatta, and Giovanni Pecenik. Eutrophication model of the Venice Lagoon: Statistical treatment of ‘in situ’ measurements of phytoplankton growth parameters. Ecological Modelling, 37(1–2):103– 130, June 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008790086X.

Breck:1988:PIS

[BDVC88] James E. Breck, Donald L. DeAngelis, Webb Van Winkle, and Sigurd W. Christensen. Potential importance of spatial and temporal heterogeneity in pH, Al, and Ca in allowing survival of a fish population: a model demonstration. Ecological Modelling, 41(1–2):1–16, April 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900403.

Beck:1981:HSE

[Bec81] M. B. Beck. Hard or soft environmental systems? Ecological Modelling, 11(4):233–251, February 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380081900600.

Belyaev:1984:SFC

[Bel84] Valery I. Belyaev. Simulation of functioning of a complex ecosystem. Ecological Modelling, 26(1–2):9–15, December 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900887.

Bergner:1985:RBB

[Ber85] Per-Erik E. Bergner. On relations between bioaccumulation and weight of organisms. Ecological Modelling, 27(3–4):207– 220, April 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900031.

Blake:1987:EMH

[BG87] G. Blake and S. Gentil. Ecological modelling of a high mountain reservoir in relation to particulate organic matter loading. Ecological Modelling, 35(3–4):227–247, March 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087901141.

Bartell:1984:FAM

[BGO84] S. M. Bartell, R. H. Gardner, and R. V. O’Neill. The fates of aromatics model (FOAM): Description, application, and analysis. Ecological Modelling, 22(1–4):109–121, April 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900103.

Beghelli:1982:IPD

[BGTH82] S. Beghelli, R. Guidorzi, F. Terragni, and E. Halfon. Identification of phosphorus dynamics in Kootenay Lake, Canada. Ecological Modelling, 17(1):11–32, September 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900849.

Botsford:1984:OFP

[BH84] Louis W. Botsford and Roderick C. Hobbs. Optimal fishery policy with artificial enhancement through stocking: California’s white sturgeon as an example. Ecological Modelling, 23(4):293– 312, July 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901261.

Bachelet:1989:SCN

[BHD89a] D. M. Bachelet, H. W. Hunt, and J. K. Detling. Simulated carbon and nitrogen dynamics in blue grama swards subject to above- and belowground grazing, irrigation and fertilization part. II. The grazing optimization notion. Ecological Modelling, 48(1–2):83–99, October 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900616.

Bachelet:1989:SMI

[BHD89b] Dominique Bachelet, H. W. Hunt, and J. K. Detling. A simulation model of intraseasonal carbon and nitrogen dynamics of blue grama swards as influenced by above- and belowground grazing. Ecological Modelling, 44(3–4):231–252, January 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008990032X.

Biswas:1988:ISE

[Bis88] Asit K. Biswas. International seminar on ecosystems modelling, 18–22 January 1988, New Delhi, India. Ecological Modelling, 42(2):163–164, August 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088901159.

Bobba:1989:AIA

[BJ89] A. G. Bobba and S. R. Joshi. Application of an inverse approach to a Canadian radioactive waste disposal site. Ecological Modelling, 46(3–4):195–211, August 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900173.

Bolla:1984:SNL

[BK84] Marianna Bolla and Tibor Kutas. Submodels for the nutrient loading estimation on River Zala. Ecological Modelling, 26(1–2):115–143, December 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008490098X.

Bazykin:1986:EDG

[BK86] A. D. Bazykin and A. S. Kondrashov. Evolutionary dynamics of genetic diversity: Proc., Manchester, Great Britain, 29–30 March 1983. G. S. Mani (Editor). Lecture Notes in Biomathematics, 53. Springer, Berlin, 1984. vii + 312 pp., DM 44.00 softcover. ISBN 3-540-12903-0. Ecological Modelling, 34(1–2):139–141, November 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900876.

Bhargava:1984:RDS

[BKU84] S. C. Bhargava, Karmeshu, and T. E. Unny. Role of diffusion in some growth models. Ecological Modelling, 24(1–2):1– 8, August 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900516.

Blackwell:1983:ERC

[Bla83] Ann Lowes Blackwell. Extremes in the resource consumption process: a mathematical model of the irruption of the moose population on Isle Royale. Part 1 — ecological background and model development. Ecological Modelling, 20(1):47–69, October 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900315.

Bolter:1986:SED

[BM86] Manfred Bölter and Michael Meyer. Structuring of ecological data sets by methods of correlation and cluster analysis. Ecological Modelling, 32(1–3):1–13, June 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900128.

Batchelder:1989:LHP

[BM89] Harold P. Batchelder and Charles B. Miller. Life history and population dynamics of Metridia pacifica: Results from simulation modelling. Ecological Modelling, 48(1–2):113–136, October 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008990063X.

Bernardi:1989:RCS

[BMJ89] M. Bernardi, J. W. Mishoe, and J. W. Jones. Raghuva: a computer simulation of Raghuva albipunctella population dynamics, and Pennisetum americanum and P. typhoides phenology. Ecological Modelling, 44(3–4):275–298, January 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900343.

Brown:1982:MMP

[BMOW82] Mark P. Brown, John J. A. McLaughlin, Joseph M. O’Connor, and Kevin Wyman. A mathematical model of PCB bioaccumulation in plankton. Ecological Modelling, 15(1):29–47, February 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900667.

Brouwer:1987:SDI

[BN87] Floor Brouwer and Peter Nijkamp. A satellite design for integrated regional environmental modelling. Ecological Modelling, 35(1–2):137–148, February 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900950.

Bolter:1982:SBW

[Böl82] Manfred Bölter. Submodels of a brackish water environment. I. Temperature and microbial activity. Ecological Modelling, 17(3–4):311–318, December 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900370.

Bonan:1989:CMS

[Bon89] Gordon B. Bonan. A computer model of the solar radiation, soil moisture, and soil thermal regimes in boreal forests. Ecological Modelling, 45(4):275–306, June 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900768.

Bosatta:1981:QASa

[Bos81a] Ernesto Bosatta. A qualitative analysis of the stability of the root-microorganism soil system I. Carbon–nitrogen status and nitrogen mineralization. Ecological Modelling, 13(4):223–236, September 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900296.

Bosatta:1981:QASb

[Bos81b] Ernesto Bosatta. A qualitative analysis of the stability of the root-microorganism soil system II. Combined effect of several factors. Ecological Modelling, 13(4):237–245, September 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900302.

Bosatta:1983:AAC

[Bos83] Ernesto Bosatta. An alternative approach to calculating chemical equilibrium in soil reactions. Ecological Modelling, 20(2–3):165–173, November 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900054.

Bossel:1986:DFD

[Bos86] Hartmut Bossel. Dynamics of forest dieback: Systems analysis and simulation. Ecological Modelling, 34(3–4):259–288, December 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900086.

Bosserman:1989:SCM

[Bos89] Robert W. Bosserman. Sensitivity of cycling measures derived from ecological flow analysis. Ecological Modelling, 48(1–2):45– 64, October 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900598.

Boucher:1985:LVM

[Bou85] Douglas H. Boucher. Lotka-volterra models of mutualism and positive density-dependence. Ecological Modelling, 27(3– 4):251–270, April 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380085900067.

Benndorf:1982:PAE

[BR82a] Jürgen Benndorf and Frieder Recknagel. Problems of application of the ecological model SALMO to lakes and reservoirs having various trophic states. Ecological Modelling, 17(2):129–145, October 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900503.

Bosserman:1982:EAU

[BR82b] Robert W. Bosserman and Rammohan K. Ragade. Ecosystem analysis using fuzzy set theory. Ecological Modelling, 16(2–4):191–208, August 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900084.

Braner:1985:SRP

[Bra85] Moshe Braner. A simple random path method for the analysis of flow networks. Ecological Modelling, 28(3):165–180, June 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008590081X.

Bro-Rasmussen:1984:HAS

[BRC84] Finn Bro-Rasmussen and Kim Christiansen. Hazard assessment — a summary of analysis and integrated evaluation of exposure and potential effects from toxic environmental chemicals. Ecological Modelling, 22(1–4):67–84, April 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900073.

Brewer:1989:SPF

[Bre89] John W. Brewer. Spreadsheets, PC’s and the finitedifference solutions for ecological distribution. Ecological Modelling, 47(1–2):65–83, September 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089901105.

Brittain:1983:MSF

[Bri83] E. G. Brittain. A model system of four species with a unique equilibrium point. Ecological Modelling, 19(3):199–211, July 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380083900546.

Brock:1981:CSR

[Bro81] Thomas D. Brock. Calculating solar radiation for ecological studies. Ecological Modelling, 14(1–2):1–19, November 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900119.

Bakule:1982:MOO

[BS82]  Lubomír Bakule and Milan Straškraba.  On multiple

objective optimization in aquatic ecosystems. Ecological Modelling, 17(2):75–82, October 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900448.

Bakule:1984:OME

[BS84] Lubomír Bakule and Milan Straškraba. On optimality in multispecies ecosystems. Ecological Modelling, 26(1–2):33–39, December 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900917.

Bakule:1987:SCS

[BS87] Lubomír Bakule and Milan Straškraba. On structural control strategies in aquatic ecosystems. Ecological Modelling, 39(1–2):171–180, November 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900196.

Bossel:1989:GSM

[BS89] Hartmut Bossel and Heiner Schäfer. Generic simulation model of forest growth, carbon and nitrogen dynamics, and application to tropical acacia and European spruce. Ecological Modelling, 48(3–4):221–265, November 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900501.

Bertuglia:1982:SMU

[BT82] Cristoforo Sergio Bertuglia and Roberto Tadei. A stochastic model for the use of a country park. Ecological Modelling, 15(2):87–106, March 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900552.

Burclay:1986:HPD

[Bur86] Hugh J. Burclay. Host-parasitoid dynamics: Effects of the position of density dependence. Ecological Modelling, 32(4):291– 299, July 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900931.

Borighem:1981:MCU

[BV81] G. Borighem and J. Vereecken. Model of a chemostat utilising phenol as inhibitory substrate. Ecological Modelling, 12(4):231– 243, May 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900405.

Braat:1986:EEM

[BV86] Leon C. Braat and Wal F. J. Van Lierop. Economicecological modeling: an introduction to methods and applications. Ecological Modelling, 31(1–4):33–44, May 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900530.

Barclay:1980:MST

[BvdD80] Hugh J. Barclay and Pauline van den Driessche. A model for a species with two life history stages and added mortality. Ecological Modelling, 11(3):157–166, December 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900812.

Brown:1983:SRE

[BWC83] M. W. Brown, F. M. Williams, and E. Alan Cameron. Simulations on the role of the egg parasite, Ooencyrtus kuvanae (Howard), in the population dynamics of the gypsy moth. Ecological Modelling, 18(3–4):253–268, April 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900170.

Byrne:1987:IGM

[BWKR87] Stephen V. Byrne, Michael M. Wehrle, Michael A. Keller, and James F. Reynolds. Impact of gypsy moth infestation on forest succession in the North Carolina Piedmont: a simulation study. Ecological Modelling, 35(1–2):63–84, February 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900913.

Cochrane:1987:EMP

[CAJ^{+}87] K. L. Cochrane, P. J. Ashton, A. C. Jarvis, A. J. Twinch, and T. Zohary. An ecosystem model of phosphorus cycling in a warm monomictic, hypertrophic impoundment. Ecological Modelling, 37(3–4):207–233, July 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900263.

Cale:1988:CPE

[Cal88] William G. Cale. Characterizing populations as entities in ecosystem models: Problems and limitations of massbalance modeling. Ecological Modelling, 42(2):89–102, August 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088901093.

Carey:1982:DPD

[Car82] James R. Carey. Demography and population dynamics of the Mediterranean fruit fly. Ecological Modelling, 16(2–4):125–150, August 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900059.

Carpenter:1984:ETP

[Car84] Stephen R. Carpenter. Experimental test of the pupation window model for development of detritivorous insects. Ecological Modelling, 23(3):257–264, June 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084901042.

Casti:1982:CCI

[Cas82] J. Casti. Catastrophes, control and the inevitability of spruce budworm outbreaks. Ecological Modelling, 14(3–4):293–300, January 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900242.

Caswell:1988:TME

[Cas88] Hal Caswell. Theory and models in ecology: a different perspective. Ecological Modelling, 43(1–2):33–44, October 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900713.

Caswell:1989:ALT

[Cas89] Hal Caswell. Analysis of life table response experiments. I. Decomposition of effects on population growth rate. Ecological Modelling, 46(3–4):221–237, August 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900197.

Costanza:1987:TEE

[CD87] Robert Costanza and Herman E. Daly. Toward an ecological economics. Ecological Modelling, 38(1–2):1–7, September 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008790041X.

Christophersen:1984:MSS

[CDJS84] Nils Christophersen, Leif H. Dymbe, Merete Johannssen, and Hans M. Seip. A model for sulphate in streamwater at Storgama, Southern Norway. Ecological Modelling, 21(1–2):35– 61, January 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900243.

Cropper:1987:RCS

[CE87] Wendel P. Cropper and Katherine Carter Ewel. A regional carbon storage simulation for large-scale biomass plantations. Ecological Modelling, 36(3–4):171–180, May 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900664.

Clark:1984:FTV

[CEMF84] T. P. Clark, D. L. Euler, J. A. McDonnell, and S. E. Fast. Field test of a vegetation disturbance index. Ecological Modelling, 23(1–2):91–100, May 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084901200.

Cerrito:1989:DEA

[Cer89] P. B. Cerrito. Density estimation applied to stochastic abundance models. Ecological Modelling, 45(3):221–236, May 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900835.

Comins:1988:SFF

[CF88] Hugh N. Comins and Brian S. Fletcher. Simulation of fruit fly population dynamics, with particular reference to the olive fruit fly, Dacus oleae. Ecological Modelling, 40(3–4):213–231, March 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900191.

Carothers:1983:SAS

[CG83] Paul E. Carothers and W. E. Grant. Systems analysis and simulation of the NW African marine upwelling region. Ecological Modelling, 19(2):73–103, June 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900236.

Carothers:1987:FMI

[CG87] Paul E. Carothers and William E. Grant. Fishery management implications of recruitment seasonality: Simulation of the Texas fishery for the brown shrimp, Penaeus aztecus. Ecological Modelling, 36(3–4):239–268, May 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900706.

Cosby:1984:IPLa

[CH84] B. J. Cosby and G. M. Hornberger. Identification of photosynthesis-light models for aquatic systems I. Theory and simulations. Ecological Modelling, 23(1–2):1–24, May 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084901169.

Chaudhuri:1986:BMH

[Cha86] Kripasindhu Chaudhuri. A bioeconomic model of harvesting a multispecies fishery. Ecological Modelling, 32(4):267– 279, July 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900918.

Chaudhuri:1988:DOC

[Cha88] Kripasindhu Chaudhuri. Dynamic optimization of combined harvesting of a two-species fishery. Ecological Modelling, 41 (1–2):17–25, April 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380088900415.

Catchpole:1989:FST

[CHC89] E. A. Catchpole, T. J. Hatton, and W. R. Catchpole. Fire spread through nonhomogeneous fuel modelled as a Markov process. Ecological Modelling, 48(1–2):101–112, October 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900628.

Cheng:1989:PME

[Che89] Zhibin Cheng. Pansystems modelling in ecology. Ecological Modelling, 47(3–4):275–281, September 15, 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900057.

Cosby:1984:IPLb

[CHK84] B. J. Cosby, G. M. Hornberger, and M. G. Kelly. Identification of photosynthesis-light models for aquatic systems II. Application to a macrophyte dominated stream. Ecological Modelling, 23(1–2):25–51, May 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084901170.

Christensen:1987:CRM

[Chr87] Paul P. Christensen. Classical roots for a modern materialsenergy analysis. Ecological Modelling, 38(1–2):75–89, September 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900457.

Chytil:1984:DRP

[Chy84] Ivo Chytil. Dispersion of radioactive pollution in surface water. Ecological Modelling, 26(1–2):145–153, December 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900991.

Csetenyi:1989:LMR

[CL89] Arthur I. Csetenyi and Dmitrii O. Logofet. Leslie model revisited: some generalizations to block structures. Ecological Modelling, 48(3–4):277–290, November 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900525.

Cleveland:1987:BEH

[Cle87]  Cutler J. Cleveland.  Biophysical economics: Histori-

cal perspective and current research trends. Ecological Modelling, 38(1–2):47–73, September 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900445.

Childers:1987:SSW

[CM87] Daniel L. Childers and Henry N. McKellar. A simulation of saltmarsh water column dynamics. Ecological Modelling, 36(3–4):211–238, May 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008790069X.

Carsel:1985:PRZ

[CMLB85] Robert F. Carsel, Lee A. Mulkey, Matthew N. Lorber, and Leland B. Baskin. The pesticide root zone model (PRZM): a procedure for evaluating pesticide leaching threats to groundwater. Ecological Modelling, 30(1–2):49–69, October 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900365.

Camara:1986:EEM

[CMM^{+}86] António S. Câmara, António P. Mano, M. Graça Martinho, M. Paula Marques, João F. Nunes, Teresa C. Lopes, and António Cabeleira. An economic-ecological model for regional land-use planning. Ecological Modelling, 31(1–4):293–302, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900700.

Coughenour:1984:MPP

[CMW84a] M. B. Coughenour, S. J. McNaughton, and L. L. Wallace. Modelling primary production of perennial graminoids 3$ uniting physiological processes and morphometric traits. Ecological Modelling, 23(1–2):101–134, May 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084901212.

Coughenour:1984:SSE

[CMW84b] M. B. Coughenour, S. J. McNaughton, and L. L. Wallace. Simulation study of east-African perennial graminoid responses to defoliation. Ecological Modelling, 26(3–4):177–201, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900693.

Cale:1988:ACP

[CO88] W. G. Cale and R. V. O’Neill. Aggregation and consistency problems in theoretical models of exploitative resource competition. Ecological Modelling, 40(2):97–109, February 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088901056.

Colorni:1983:GES

[Col83] A. Colorni. Geography and the environment: Systems analytical methods: A. G. Wilson. John Wiley, W. Sussex, U. K., 1981, Hardback £15.50, US $35.50, ISBN: 0-471-27956-0; Paperback £7.50, ISBN: 0-471-27957-9. Ecological Modelling, 19(3):221– 222, July 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008390056X.

Cale:1983:DAD

[COS83] W. G. Cale, R. V. O’Neill, and H. H. Shugart. Development and application of desirable ecological models.

Ecological Modelling, 18(3–4):171–186, April 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008390011X.

Costanza:1989:MGF

[Cos89] Robert Costanza. Model goodness of fit: a multiple resolution procedure. Ecological Modelling, 47(3–4):199–215, September 15, 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008990001X.

Coughenour:1981:RDD

[Cou81] M. B. Coughenour. Relationship of SO_{2} dry deposition to a grassland sulfur cycle. Ecological Modelling, 13(1–2):1–16, June 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900041.

Coughenour:1984:MSA

[Cou84] M. B. Coughenour. A mechanistic simulation analysis of water use, leaf angles, and grazing in East African graminoids. Ecological Modelling, 26(3–4):203–230, December 15, 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008490070X.

Chen:1988:NME

[CP88] Keh-Wei Chen and Alex S. Papadopoulos. A nonparametric method for estimating the joint probability density of BOD and DO. Ecological Modelling, 41(3–4):183–191, June 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900269.

Croom:1980:MRC

[CR80] John M. Croom and Harvey L. Ragsdale. A model of radiocesium cycling in a sand hills–Turkey Oak (Quercus laevis) ecosystem. Ecological Modelling, 11(1):55–65, October 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008090071X.

Costanza:1985:AAE

[CS85] Robert Costanza and Fred H. Sklar. Articulation, accuracy and effectiveness of mathematical models: a review of freshwater wetland applications. Ecological Modelling, 27(1–2):45–68, March 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900249.

Conley:1989:DSD

[CS89] Walt Conley and Uttam Sengupta. A demographic simulator with deeply coupled semantic and numeric data structures. Ecological Modelling, 46(1–2):35–56, July 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900689.

Cuenco:1985:FBGa

[CSG85a] Michael L. Cuenco, Robert R. Stickney, and William E. Grant. Fish bioenergetics and growth in aquaculture ponds: I. Individual fish model development. Ecological Modelling, 27(3–4):169– 190, April 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900018.

Cuenco:1985:FBGb

[CSG85b] Michael L. Cuenco, Robert R. Stickney, and William E. Grant. Fish bioenergetics and growth in aquaculture ponds: II. Effects of interactions among, size, temperature, dissolved oxygen, unionized ammonia and food on growth of individual fish. Ecological Modelling, 27(3–4):191–206, April 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008590002X.

Cuenco:1985:FBGc

[CSG85c] Michael L. Cuenco, Robert R. Stickney, and William E. Grant. Fish bioenergetics and growth in aquaculture ponds: III. Effects of intraspecific competition, stocking rate, stocking size and feeding rate on fish productivity. Ecological Modelling, 28(1–2):73–95, May 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900146.

Clark:1982:SSH

[CT82] Colin W. Clark and David E. Tait. Sex-selective harvesting of wildlife populations. Ecological Modelling, 14(3–4):251–260, January 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900217.

Crete:1981:SCR

[CTJ81] Michel Crête, R. J. Taylor, and P. A. Jordan. Simulating conditions for the regulation of a moose population by wolves. Ecological Modelling, 12(4):245–252, May 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900417.

Carruthers:1986:SAR

[CWTH86] R. I. Carruthers, G. H. Whitfield, R. L. Tummala, and D. L. Haynes. A systems approach to research and simulation of insect pest dynamics in the onion agro-ecosystem. Ecological Modelling, 33(2–4):101–121, October 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900359.

David:1982:CSM

[DAA82] J. M. David, L. Andral, and M. Artois. Computer simulation model of the EPI-enzootic disease of vulpine rabies. Ecological Modelling, 15(2):107–125, March 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900564.

DeAngelis:1984:SPM

[DABG84] D. L. DeAngelis, S. M. Adams, J. E. Breck, and L. J. Gross. A stochastic predation model: Application to largemouth bass observations. Ecological Modelling, 24(1–2):25–41, August 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008490053X.

Dahlman:1985:MNP

[Dah85] Roger C. Dahlman. Modeling needs for predicting responses to CO_{2} enrichment: plants, communities and ecosystems.

Ecological Modelling, 29(1–4):77–106, September 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900481.

Dayananda:1982:SMP

[Day82] P. W. A. Dayananda. A simple model for the path of a solute in root-zone of soil. Ecological Modelling, 16(2–4):81–84, August 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900011.

Dayong:1988:IMS

[Day88] Zhang Dayong. An index to measure the strength of relationship between community and site. Ecological Modelling, 40(2):145–153, February 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380088901081.

Dayong:1989:MDD

[Day89] Zhang Dayong. A method of detecting departure from randomness in plant communities. Ecological Modelling, 46(3–4):261– 267, August 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900215.

deCaprariis:1981:NDT

[dC81] Pascal de Caprariis. A note on the development of the thermocline in temperate lakes. Ecological Modelling, 12(4):213– 219, May 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900387.

deCaprariis:1984:NSL

[dC84] Pascal de Caprariis. A note on the sufficiency of low resolution ecosystem models. Ecological Modelling, 21(3):199–207, February 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901376.

Dennis:1984:PMR

[DDM84] Robin L. Dennis, Mary W. Downton, and Paulette Middleton. Policy-making and the role of simplified models: an air quality planning example. Ecological Modelling, 25(1–3):1–30, October 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900309.

Dale:1985:CTG

[DDS85] V. H. Dale, T. W. Doyle, and H. H. Shugart. A comparison of tree growth models. Ecological Modelling, 29(1–4):145–169, September 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900511.

DeGroot:1983:MMN

[De 83] W. T. De Groot. Modelling the multiple nutrient limitation of algal growth. Ecological Modelling, 18(2):99–119, February 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380083900492.

DeAngelis:1988:SDA

[DeA88] D. L. DeAngelis. Strategies and difficulties of applying models to aquatic populations and food webs. Ecological Modelling, 43(1–2):57–73, October 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900737.

DeLuna:1987:ETP

[DH87a] J. T. De Luna and T. G. Hallam. Effects of toxicants on populations: a qualitative approach IV. Resource–consumer– toxicant models. Ecological Modelling, 35(3–4):249–273, March 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087901153.

DeAngelis:1987:EGR

[DH87b] D. L. DeAngelis and M. A. Huston. Effects of growth rates in models of size distribution formation in plants and animals. Ecological Modelling, 36(1–2):119–137, April 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900627.

DeGrandi-Hoffman:1989:BHP

[DHRLE89] G. DeGrandi-Hoffman, S. A. Roth, G. L. Loper, and E. H. Erickson. BEEPOP: a honeybee population dynamics simulation model. Ecological Modelling, 45(2):133–150, April 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900884.

Dale:1988:USU

[DJGR88] V. H. Dale, H. I. Jager, R. H. Gardner, and A. E. Rosen. Using sensitivity and uncertainty analyses to improve predictions of broad-scale forest development. Ecological Modelling, 42(3–4):165–178, September 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900555.

DeJong:1983:SME

[DK83a] Tom J. De Jong and Peter G. L. Klinkhamer. A simulation model for the effects of burning on the phosphorus and nitrogen cycle of a heathland ecosystem. Ecological Modelling, 19(4):263–284, September 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900431.

Dwyer:1983:FDS

[DK83b] Robert L. Dwyer and James N. Kremer. Frequency-domain sensitivity analyses of an estuarine ecosystem simulation model. Ecological Modelling, 18(1):35–54, January 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900741.

Dostalkova:1984:SSR

[DKR84] Iva Dostálková, Pavel Kindlmann, and Marcel Rejmánek. Simulation of species replacement on environmental gradient in the course of ecological succession. Ecological Modelling, 26(1–2):45–50, December 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900930.

Dejak:1987:DME

[DLMP87a] Camillo Dejak, Ileana Mazzei Lalatta, Letizia Meregalli, and Giovanni Pecenik. Development of a mathematical eutrophication model of the lagoon of Venice. Ecological Modelling, 37(1–2):1–20, June 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900810.

Dejak:1987:ADP

[DLMP87b] Camillo Dejak, Ileana Mazzei Lalatta, Ettore Messina, and Giovanni Pecenik. An advection-diffusion pollution model of the Lagoon of Venice. Ecological Modelling, 37(1–2):47–57, June 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900834.

Dejak:1987:SSA

[DLMP87c] Camillo Dejak, Ileana Mazzei Lalatta, Ettore Messina, and Giovanni Pecenik. Steady-state achievement by introduction of true tidal velocities in a pollution model of the Venice Lagoon. Ecological Modelling, 37(1–2):59–79, June 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900846.

Dejak:1987:TDD

[DLMP87d] Camillo Dejak, Ileana Mazzei Lalatta, Ettore Messina, and Giovanni Pecenik. A two-dimensional diffusion model of the Venice lagoon and relative open boundary conditions. Ecological Modelling, 37(1–2):21–45, June 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900822.

Dejak:1987:TTD

[DLMP87e] Camillo Dejak, Ileana Mazzei Lalatta, Marina Molin, and Giovanni Pecenik. Tidal three-dimensional diffusion in a model of the Lagoon of Venice and reliability conditions for its numerical integration. Ecological Modelling, 37(1–2):81–101, June 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900858.

Davis:1989:TSK

[DNHP89] J. R. Davis, P. M. Nanninga, J. R. L. Hoare, and A. J. Press. Transferring scientific knowledge to natural resource managers using artificial intelligence concepts. Ecological Modelling, 46(1–2):73–89, July 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900707.

Dejak:1987:E

[DP87] Camillo Dejak and Giovanni Pecenik. Editorial. Ecological Modelling, 37(1–2):vii–viii, June 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900809.

DeJong:1985:MSM

[DS85] Mart C. M. De Jong and Hannu Saarenmaa. A mechanistic simulation model for the movement and competition of bark beetle larvae (coleoptera, scolytidae). Ecological Modelling, 27(1–2):109–138, March 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900286.

Duo-Sen:1987:KMD

[DSSM87] Liu Duo-Sen and Zhang Shui-Ming. Kinetic model for degradative processes of pesticides in soil. Ecological Modelling, 37(3–4):131–138, July 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900214.

Duo-Sen:1988:SRM

[DSSMZG88] Liu Duo-Sen, Zhang Shui-Ming, and Li Zhen-Gao. Study on rate model of microbial degradation of pesticides in soil. Ecological Modelling, 41(1–2):75–84, April 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900464.

Dabrowski:1989:BMA

[DTL89] K. Dabrowski, F. Takashima, and Y. K. Law. Bioenergetic model for the analysis of the ontogenetical aspects of coregonid fish growth. Ecological Modelling, 44(3–4):195–208, January 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008990029X.

Duchting:1986:MAM

[DV86] W. Düchting and T. Vogelsaenger. Methodological aspects of modelling tumor growth and treatment. Ecological Modelling, 32(1–3):191–197, June 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900268.

Downing:1984:PFP

[DW84] John A. Downing and Laura A. Weber. The prediction of forest production from inventory and climatic data. Ecological Modelling, 23(3):227–241, June 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084901029.

Duffy:1988:MFS

[DW88] David Cameron Duffy and Christian Wissel. Models of fish school size in relation to environmental productivity. Ecological Modelling, 40(3–4):201–211, March 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008890018X.

DeAngelis:1985:EMD

[DWPO85] D. L. DeAngelis, J. C. Waterhouse, W. M. Post, and R. V. O’Neill. Ecological modelling and disturbance evaluation. Ecological Modelling, 29(1–4):399–419, September 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900638.

Dyer:1981:TIC

[Dye81] T. G. J. Dyer. On a technique to investigate the covariation between streamflow and rainfall in afforested catchments. Ecological Modelling, 13(3):149–157, August 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900491.

El-Bayoumi:1984:MSE

[EBSW84] Mohamed A. El-Bayoumi, Herman H. Shugart, and Ross W. Wein. Modelling succession of Eastern Canadian mixedwood forest. Ecological Modelling, 21(3):175–198, February 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084901364.

Ewing:1982:MAP

[EEKM82] M. S. Ewing, S. A. Ewing, M. S. Keener, and R. J. Mulholland. Mutualism among parasitic nematodes: a population model. Ecological Modelling, 15(4):353–366, June 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900898.

Ewel:1982:EWA

[EF82] Katherine Carter Ewel and Thomas D. Fontaine. Effects of white Amur (Ctenopharyngodon idella) on a Florida lake: a model. Ecological Modelling, 16(2–4):251–273, August 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900114.

Ewel:1983:SFW

[EF83] Katherine Carter Ewel and Thomas D. Fontaine. Structure and function of a warm monomictic lake. Ecological Modelling, 19(2):139–161, June 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900261.

Eston:1986:CBC

[EGJ^{+}86] Verena R. Eston, Antonio Galves, Claudia M. Jacobi, Rémi Langevin, and Nelson I. Tanaka. Chthamalus bisinuatus (Cirripedia) and Brachidontes solisianus (Bivalvia) spatial interactions: a stochastic model. Ecological Modelling, 34(1–2):99–113, November 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900815.

Euler:1984:SPD

[EM84] David Euler and Michael M. J. Morris. Simulated population dynamics of white-tailed deer in an any-deer hunting system. Ecological Modelling, 24(3–4):281–292, September 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900462.

Eilers:1988:MRB

[EP88] P. H. C. Eilers and J. C. H. Peeters. A model for the relationship between light intensity and the rate of photosynthesis in phytoplankton. Ecological Modelling, 42(3–4):199–215, September 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900579.

Emlen:1989:APD

[EP89] John M. Emlen and Ellen K. Pikitch. Animal population dynamics: Identification of critical components. Ecological Modelling, 44(3–4):253–273, January 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900331.

El-Shaarawi:1985:PME

[ES85] A. H. El-Shaarawi. Principles and measurements in environmental biology: F. I. Woodward and J. E. Sheehy. Butterworths, London, 1983. ix + 263 pp., £32.50. ISBN 0-408-10637-9. Ecological Modelling, 28(3):236, June 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900882.

Esser:1989:GLU

[Ess89] G. Esser. Global land-use changes from 1860 to 1980 and future projections to 2500. Ecological Modelling, 44(3–4):307–316, January 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900367.

Ewel:1989:LSE

[Ewe89] Katherine C. Ewel. Learning to simulate ecological models on a microcomputer. Ecological Modelling, 47(1–2):7–17, September 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089901063.

Fahrig:1988:NET

[Fah88] Lenore Fahrig. Nature of ecological theories. Ecological Modelling, 43(1–2):129–132, October 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900774.

Feldman:1981:MMF

[FCC81] Richard M. Feldman, Guy L. Curry, and Robert N. Coulson. A mathematical model of field population dynamics of the southern pine beetle, Dendroctonus frontalis. Ecological Modelling, 13(4):261–281, September 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900326.

Furlong:1986:TLC

[FD86] Eileen A.-N. Furlong and Frank M. D’Itri. Trihalomethane levels in chlorinated Michigan drinking water. Ecological Modelling, 32(1–3):215–225, June 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900281.

Focardi:1989:TAR

[FDC89] Stefano Focardi, Jean Louis Deneubourg, and Guido Chelazzi. Theoretical analysis of rhythmical clustering in an intertidal gastropod. Ecological Modelling, 44(3–4):177–194, January 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900288.

Fitz-Earle:1989:TOS

[FEB89] Malcolm Fitz-Earle and Hugh J. Barclay. Is there an optimal sex ratio for insect mass rearing? Ecological Modelling, 45(3):205–220, May 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900823.

Fedra:1984:IWQ

[Fed84] Kurt Fedra. Interactive water quality simulation in a regional framework: a management oriented approach to lake and watershed modeling. Ecological Modelling, 21(4):209–232, March 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900607.

Fetcher:1981:EGC

[Fet81] Ned Fetcher. Effects of grazing on cold desert shrubs: a simulation model based on relative growth rate. Ecological Modelling, 13(1–2):49–86, June 1981. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380081900077.

Furniss:1982:MSL

[FFMB82] P. R. Furniss, P. Ferrar, J. W. Morris, and J. J. Bezuidenhout. A model of savanna litter decomposition. Ecological Modelling, 17(1):33–51, September 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900850.

Frelek:1982:OAV

[FGL82] B. Frelek, M. Gatto, and A. Locatelli. Optimal allocation of vessels along a fish migration path. Ecological Modelling, 14(3–4):229–250, January 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900205.

Ferson:1989:EER

[FGS89] Scott Ferson, Lev Ginzburg, and Abraham Silvers. Extreme event risk analysis for age-structured populations. Ecological Modelling, 47(1–2):175–187, September 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089901166.

Fitter:1984:VDG

[Fit84] A. H. Fitter. Vegetation dynamics in Grasslands, Heathlands and Mediterranean Ligneous formations: P. Poissonet, F. Romane, M. P. Austin, E. van der Maarel and W. Schmidt (Editors). Advances in Vegetation Science No. 4, W. Junk, The Hague, 1981, x + 286 pp., Dfl. 195.00 ISBN 90-6193-636-5. Ecological Modelling, 23(3):267–268, June 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084901066.

Fleishman:1982:STE

[Fle82] Benzion Semionovich Fleishman. Stochastic theory of ecological interactions. Ecological Modelling, 17(2):65–73, October 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900436.

Fleishman:1984:CTA

[Fle84] Benzion Semionovich Fleishman. Contribution to the theory of adaptation with application to ecology. Ecological Modelling, 26(1–2):21–31, December 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900905.

Fleishman:1987:STC

[Fle87] Benzion Semionovich Fleishman. Stochastic theory of community control. Ecological Modelling, 39(1–2):121–159, November 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900172.

Fields:1988:MDM

[FM88] D. E. Fields and C. W. Miller. A methodology for deriving model input parameters from a set of environmental data. Ecological Modelling, 40(3–4):155–159, March 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900154.

Frisk:1981:CSP

[FNK81] T. Frisk, J. S. Niemi, and K. A. I. Kinnunen. Comparison of statistical phosphorus-retention models. Ecological Modelling, 12(1–2):11–27, March 1981. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380081900223.

Fontaine:1984:ARU

[Fon84a] Thomas D. Fontaine. Application of risk and uncertainty analysis techniques to a heavy metal speciation model. Ecological Modelling, 22(1–4):101–108, April 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900097.

Fontaine:1984:NEAa

[Fon84b] Thomas D. Fontaine. A non-equilibrium approach to modeling metal speciation in acid, aquatic systems: theory and process equations. Ecological Modelling, 21(4):287–313, March 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900656.

Fontaine:1984:NEAb

[Fon84c] Thomas D. Fontaine. A non-equilibrium approach to modeling toxic metal speciation in acid, aquatic systems. Ecological Modelling, 22(1–4):85–100, April 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900085.

Favretto:1986:ETS

[FP86] D. Favretto and L. Poldini. Extinction time of a sample of karst pastures due to bush encroachment. Ecological Modelling, 33(2–4):85–88, October 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900335.

Folse:1989:AMA

[FPG89] L. Joseph Folse, Jane M. Packard, and William E. Grant. AI modelling of animal movements in a heterogeneous habitat. Ecological Modelling, 46(1–2):57–72, July 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900690.

French:1989:SMS

[FRCC89] Deborah P. French, Mark Reed, John Calambokidis, and James C. Cubbage. A simulation model of seasonal migration and daily movements of the northern fur seal. Ecological Modelling, 48(3–4):193–219, November 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900495.

Frisman:1982:MMP

[FSK82] E. Ya. Frisman, E. I. Skaletskaya, and A. E. Kuzin. A mathematical model of the population dynamics of a local Northern fur-seal herd. Ecological Modelling, 16(2–4):151–172, August 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900060.

Fujii:1986:RAM

[FSTF86] Shigeo Fujii, Isao Somiya, Hiroshi Tsuno, and Masahiro Fujiwara. Rational allocation of monitoring stations in a lake by means of the spline technique. Ecological Modelling, 32(1–3): 43–57, June 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900153.

Fulda:1987:LED

[Ful87] Joseph S. Fulda. The logistic equation and double jeopardy. Ecological Modelling, 36(3–4):315–316, May 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900731.

Flipse:1984:ALM

[FV84] Eric Flipse and Ed J. M. Veling. An application of the Leslie matrix model to the population dynamics of the hooded seal, Crystophora cristata Erxleben. Ecological Modelling, 24(1–2):43–59, August 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900541.

Fedra:1981:UAE

[FVB81] K. Fedra, G. Van Straten, and M. B. Beck. Uncertainty and arbitrariness in ecosystems modelling: a lake modelling example. Ecological Modelling, 13(1–2):87–110, June 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900089.

Fairall:1986:GMP

[FVV86] N. Fairall, P. J. Vermeulen, and M. Van Der Merwe. A general model of population growth in the hyrax Procavia capensis. Ecological Modelling, 34(1–2):115–132, November 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900827.

Feldman:1981:WTL

[FWS^{+}81] Richard M. Feldman, Terence L. Wagner, Peter J. H. Sharpe, James A. Gagne, and Robert N. Coulson. Within-tree life process models of the southern pine beetle, Dendroctonus frontalis. Ecological Modelling, 13(4):247–259, September 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900314.

Gobran:1989:SCK

[GÅ89] George R. Gobran and Göran I. Ågren. Significance of changes in K_{c} values for Ca–Al exchange and its effects on soil and water acidification predictions. Ecological Modelling, 44(3–4):165–175, January 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900276.

Galperin:1987:PCA

[Gal87] Michael Vladimirovich Gal’perin. Principles of computeraided event simulation in marine ecology. Ecological Modelling, 39(1–2):101–119, November 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900160.

Gard:1981:PEM

[Gar81] Thomas C. Gard. Persistence for ecosystem microcosm models. Ecological Modelling, 12(4):221–229, May 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900399.

Garsd:1984:SCE

[Gar84] Armando Garsd. Spurious correlation in ecological modelling. Ecological Modelling, 23(3):191–201, June 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084901005.

Gard:1988:ASE

[Gar88] Thomas C. Gard. Aggregation in stochastic ecosystem models. Ecological Modelling, 44(1–2):153–164, December 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900877.

Gatto:1983:MTD

[Gat83] Marino Gatto. The mathematical theory of the dynamics of biological populations II: R. W. Hiorns and D. Cooke (Editors). Academic Press, London, 1981, 327 pp., £14.50, US $35.00, ISBN: 0-12-348780-3. Ecological Modelling, 19(3):232– 234, July 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900637.

Gatto:1987:END

[Gat87] Marino Gatto. The ecology of natural disturbance and patch dynamics: S. T. A. Pickett and P. S. White (Editors). Academic Press, London/Orlando, FL, U.S.A., 1985. 472 pp. ISBN 0-12-554520-7. Ecological Modelling, 37(3–4):322–324, July 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900354.

Gatto:1988:BMF

[Gat88] Marino Gatto. Bioeconomic modelling and fisheries management: Colin W. Clark. Wiley, Chichester, Great Britain, 1985. 291 pp., £52.00. ISBN 0-471-87394-2. Ecological Modelling, 42(2):161–162, August 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088901147.

Gentil:1981:VCE

[GB81] S. Gentil and G. Blake. Validation of complex ecosystems models. Ecological Modelling, 14(1–2):21–38, November 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900120.

Gobran:1988:CDR

[GB88] George R. Gobran and Ernesto Bosatta. Cation depletion rate as a measure of soil sensitivity to acidic decomposition: Theory. Ecological Modelling, 40(1):25–36, January 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088901007.

Golkin:1984:CSC

[GE84] Kenneth R. Golkin and Katherine Carter Ewel. A computer simulation of the carbon, phosphorus, and hydrologic cycles of a pine flatwoods ecosystem. Ecological Modelling, 24(1–2):113– 136, August 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900577.

Gentil:1984:LRI

[Gen84] S. Gentil. Linear and recursive identification for ecosystems modelling. Ecological Modelling, 21(1–2):21–33, January 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900231.

Grant:1984:EVP

[GFI84] W. E. Grant, S. O. Fraser, and K. G. Isakson. Effect of vertebrate pesticides on non-target wildlife populations: Evaluation through modelling. Ecological Modelling, 21(1–2):85–108, January 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900267.

George:1983:SSM

[GG83] Lee C. George and W. E. Grant. A stochastic simulation model of brown shrimp (Penaeus aztecus Ives) growth, movement, and survival in Galveston Bay, Texas. Ecological Modelling, 19(1):41–70, May 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900698.

Gatto:1989:RSR

[GG89] Marino Gatto and Giorgio Guariso. A report on some recent experiences in developing environmental software. Ecological Modelling, 47(1–2):19–32, September 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089901075.

Griffin:1984:BMS

[GGBH84] W. L. Griffin, W. E. Grant, R. W. Brick, and J. S. Hanson. A bioeconomic model of shrimp maricultural systems in the U.S.A. Ecological Modelling, 25(1–3):47–68, October 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900322.

Green:1983:FSA

[GGN83] D. G. Green, A. M. Gill, and I. R. Noble. Fire shapes and the adequacy of fire-spread models. Ecological Modelling, 20(1):33– 45, October 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900303.

Grant:1987:F

[GH87] William E. Grant and C. G. Hoogendyk. Foreword. Ecological Modelling, 36(1–2):3, April 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900536.

Garnerin:1986:ETE

[GHV86] Ph. Garnerin, S. Hazout, and A.-J. Valleron. Estimation of two epidemiological parameters of fox rabies: the length of incubation period and the dispersion distance of cubs. Ecological Modelling, 33(2–4):123–135, October 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900360.

Grant:1981:GBS

[GIG81] W. E. Grant, K. G. Isakson, and W. L. Griffin. A general bioeconomic simulation model for annual-crop marine fisheries. Ecological Modelling, 13(3):195–219, August 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900521.

Gold:1987:NEM

[GKS87] Harvey J. Gold, William L. Kendall, and Phillip L. Shaffer. Nonlinearity and the effects of microclimatic variability on a codling moth population (Cydia pomonella): a sensitivity simulation. Ecological Modelling, 37(3–4):139–154, July 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900226.

Goodland:1987:NEP

[GL87] Robert Goodland and George Ledec. Neoclassical economics and principles of sustainable development. Ecological Modelling, 38(1–2):19–46, September 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900433.

Glasser:1988:IFA

[Gla88] John W. Glasser. Interface (and facilitation) among species that exploit alternative resources. Ecological Modelling, 40 (2):111–129, February 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380088901068.

Gatto:1982:SPE

[GLLN82] M. Gatto, A. Locatelli, E. Laniado, and M. Nuske. Some problems of effort allocation on two non-interacting fish stocks. Ecological Modelling, 14(3–4):193–211, January 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900187.

Goda:1986:SAC

[GM86] Takeshi Goda and Yuzuru Matsuoka. Synthesis and analysis of a comprehensive lake model — with the evaluation of diversity of ecosystems. Ecological Modelling, 31(1–4):11– 32, May 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900529.

Grant:1988:FCH

[GMM88] W. E. Grant, J. H. Matis, and W. Miller. Forecasting commercial harvest of marine shrimp using a Markov chain model. Ecological Modelling, 43(3–4):183–193, November 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900038.

Gatto:1988:FIL

[GMR88] Marino Gatto, Simona Muratori, and Sergio Rinaldi. A functional interpretation of the logistic equation. Ecological Modelling, 42(2):155–159, August 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088901135.

Goda:1986:P

[GNIW86] Takeshi Goda, Masaaki Naito, Saburo Ikeda, and Masataka Watanabe. Preface. Ecological Modelling, 31(1–4):vii–viii, May 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900505.

Goda:1986:MPP

[God86a] Takeshi Goda. Modelling of the persistence of pesticides applied to the soil: S. J. Troester, F. A. Ress, A. S. Felsot and W. G. Ruesink. Pudoc, Wageningen, The Netherlands, 1984. 149 pp., Dfl. 42.00. ISBN 90-220-0844-4. Ecological Modelling, 32(4):313–314, July 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900967.

Goda:1986:SPS

[God86b] Takeshi Goda. Simulation of pollution by soil erosion and soil nutrient loss: D. A. Haith, L. J. Tubbs and N. B. Pickering. Pudoc, Wageningen, The Netherlands, 1984. 77 pp., Dfl. 25.00. ISBN 90-220-0842-8. Ecological Modelling, 32(4):314– 315, July 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900979.

Gardner:1981:CSA

[GOMC81] R. H. Gardner, R. V. O’Neill, J. B. Mankin, and J. H. Carney. A comparison of sensitivity analysis and error analysis based on a stream ecosystem model. Ecological Modelling, 12(3):173– 190, April 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900569.

Gopalsamy:1980:OSH

[Gop80] K. Gopalsamy. Optimal stabilization and harvesting in logistic population models. Ecological Modelling, 11(1):67–69, October 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900721.

Gerritsen:1985:STF

[GP85] Jeroen Gerritsen and Bernard C. Patten. System theory formulation of ecological disturbance. Ecological Modelling, 29(1–4):383–397, September 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900626.

Gutierrez:1984:GDD

[GPD^{+}84] A. P. Gutierrez, M. A. Pizzamiglio, W. J. Dos Santos, R. Tennyson, and A. M. Villacorta. A general distributed delay time varying life table plant population model: Cotton (Gossypium hirsutum L.) growth and development as an example. Ecological Modelling, 26(3–4):231–249, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900711.

Gatto:1982:E

[GR82] Marino Gatto and Sergio Rinaldi. Editorial. Ecological Modelling, 14(3–4):151, January 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008290014X.

Grant:1988:BMC

[Gra88a] William E. Grant. Building models for conservation and wildlife management: A. M. Starfield and A. L. Bleloch, MacMillan, New York, 1986. 253 pp., US $34.95. ISBN 0-02-948040-X. Ecological Modelling, 41(3–4):325–326, June 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008890035X.

Grant:1988:EBI

[Gra88b] William E. Grant. Ecology of biological invasions of North America and Hawaii: H. A. Mooney and J. A. Drake (Editors). Ecological Studies, 58. Springer, Berlin, 1986. xvii + 321 pp., Hardcover, DM 148.00. ISBN 3-540-96289-1. Ecological Modelling, 43(3–4):315–317, November 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900117.

Grant:1989:VPC

[Gra89] William E. Grant. Viable populations for conservation: M. E. Soule (Editor). Cambridge University Press, Cambridge, Great Britain, 1987. Hardcover, 189 pp., £25.00/US$44.50 ISBN 0-521-33390-3. Ecological Modelling, 45(2):158–159, April 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900938.

Green:1983:SSF

[Gre83] David G. Green. Shapes of simulated fires in discrete fuels. Ecological Modelling, 20(1):21–32, October 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900297.

Gromiec:1980:SAE

[Gro80] M. J. Gromiec. State-of-the-Art in Ecological Modelling, Proceedings of the Conference on Ecological Modelling, Copenhagen, Denmark, 28 August–2 September 1978: S. E. Jørgensen (Editor). International Society for Ecological Modelling, 1979, 891 pp. US $48.00 (ISEM members), ISBN 87-87257-17-3. Ecological Modelling, 11(2):151–153, November 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900915.

Gromiec:1981:HLP

[Gro81] M. J. Gromiec. Hydrodynamics of lakes. Proceedings of a symposium, 12–13 October, 1978, Lausanne, Switzerland: W. H. Graf and C. H. Mortimer (Editors). Elsevier Scientific Publishing Company, Amsterdam — Oxford — New York, 1979, 360 pp., Dfl. 125.00, U.S. $61.00, ISBN 0-444-41827-X. Ecological Modelling, 11(4):295–297, February 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900661.

Gromiec:1983:ACI

[Gro83a] M. J. Gromiec. Aquatic chemistry, an introduction emphasizing chemical equilibria in natural waters: Werner Stumm and James J. Morgan. Wiley-Interscience, John Wiley, West Sussex, England, 1981, 780 pp., Hardback £35.25, ISBN: 0-471-04831-3; Paperback £18.65, US $31.65. Ecological Modelling, 19(3):227–230, July 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900613.

Gromiec:1983:EJC

[Gro83b] M. J. Gromiec. Ecohydrodynamics: Jacques C. J. Nihoul (Editor), Proceedings of the 12th International Liège Colloquium on Ocean Hydrodynamics. Elsevier, Amsterdam, 1981, xii + 360 pp., US $58.50, Dfl. 120.-, ISBN: 0-444-41969-1. Ecological Modelling, 19(3):230–232, July 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900625.

Gromiec:1983:MCW

[Gro83c] M. J. Gromiec. Methods of computation of the water balance of large lakes and reservoirs: Vol. I, Methodology. H. I. Ferguson and V. A. Znamensky (Editors). UN-ESCO, Paris, 1981, 120 pp., FF 32.00, ISBN: 92-3-101906-6. Ecological Modelling, 19(3):226–227, July 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900601.

Gromiec:1984:BRE

[Gro84a] M. J. Gromiec. Book review: Environmental modeling under uncertainty: Monte Carlo simulation: K. Fedra. PR-83-28, International Institute for Applied Systems Analysis, Laxenburg, Austria, 1983. 78 pp., 18 figs. ISBN 3-7045-0061-5. Ecological Modelling, 26(3–4):315–316, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900760.

Gromiec:1984:ADS

[Gro84b] M. J. Gromiec. Atmospheric diffusion: Study of the dispersion of windborne material from industrial and other sources: 3rd Edition. F. Pasquill and F. B. Smith. John Wiley and Sons, Chichester, Great Britain, 1983. 437 pp., £35.00; available in paperback from Ellis Horwood, Great Britain, £12.95/US$22.40. ISBN 0-85312-426-4. Ecological Modelling, 26(3–4):314–315, December 15, 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900759.

Gromiec:1984:CRS

[Gro84c] M. J. Gromiec. Continental radioecology: Soil and freshwater ecosystems: N. V. Kulikov and I. V. Molchanova. Plenum Press, London/New York and Nauka Publishers, Moscow, 1982. 174 pp. ISBN 0-306-40494-X. Ecological Modelling, 26(3–4):318–320, December 15, 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900784.

Gromiec:1985:SEA

[Gro85] M. J. Gromiec. Stream ecology: applications and testing of general ecological theory: James R. Barnes and G. Wayne Minshall (Editors). Plenum, New York/London, 1983. 399 pp., US$55.00. ISBN 0-306-41460-0. Ecological Modelling, 28(3):236–238, June 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900894.

Grobler:1986:AIE

[Gro86] D. C. Grobler. Assessment of the impact of eutrophication control measures on South African impoundments. Ecological Modelling, 31(1–4):237–247, May 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900669.

Grossmann:1988:PPO

[Gro88] Wolf-Dieter Grossmann. Products of photo-oxidation as a decisive factor of the new forest decline? — results and considerations. Ecological Modelling, 41(3–4):281–305, June 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900336.

Giavelli:1988:SNC

[GRS88] Giovanni Giavelli, Orazio Rossi, and Enzo Siri. Stability of natural communities: Loop analysis and computer simulation approach. Ecological Modelling, 40(2):131–143, February 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008890107X.

Grossman:1986:GMF

[GS86] W. D. Grossman and J. Schaller. Geographical maps on forest die-off, driven by dynamic models. Ecological Modelling, 31(1–4):341–353, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900748.

Guettinger:1983:CCM

[Gue83] Herbert Guettinger. Carbon cycle modelling: (SCOPE 16). Bert Bolin (Editor). John Wiley, 1981, 390 pp., Hardback £18.95, ISBN: 0-471-10051-X. Ecological Modelling, 18(3– 4):305–307, April 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380083900200.

Guiot:1986:ATM

[Gui86] J. Guiot. ARMA techniques for modelling tree-ring response to climate and for reconstructing variations of paleoclimates. Ecological Modelling, 33(2–4):149–171, October 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900384.

Gydesen:1984:SAM

[Gyd84] Helge Gydesen. A stochastic approach to models for the leaching of organic chemicals in soil. Ecological Modelling, 24(3–4):191–205, September 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900401.

Hadjibiros:1981:SPP

[Had81] Kimon Hadjibiros. Simulation of a predator–prey system using a matrix model. Ecological Modelling, 12(1–2):45–67, March 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900247.

Halfon:1983:TBMa

[Hal83a] E. Halfon. Is there a best model structure? I. Modeling the fate of a toxic substance in a lake. Ecological Modelling, 20(2–3):135–152, November 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900030.

Halfon:1983:TBMb

[Hal83b] E. Halfon. Is there a best model structure? II. Comparing the model structures of different fate models. Ecological Modelling, 20(2–3):153–163, November 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900042.

Halfon:1984:EAS

[Hal84] E. Halfon. Error analysis and simulation of Mirex behaviour in Lake Ontario. Ecological Modelling, 22(1–4):213–252, April 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900152.

Halfon:1985:TBM

[Hal85] Efraim Halfon. Is there a best model structure? III. Testing the goodness of fit. Ecological Modelling, 27(1–2):15–23, March 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900225.

Halfon:1986:DID

[Hal86a] Efraim Halfon. Disposal of industrial and domestic wastes: Land and Sea Alternatives by the Board on Ocean Science and the Policy Commission on Physical Sciences, Mathematics, and Resources National Research Council. National Academy Press, Washington, DC, USA. Distributed by Wiley, Baffins Lane, Chichester, West Sussex PO19 1UD, Great Britain. 210 pp., £19.00. ISBN 0-309-03484-1. Ecological Modelling, 32(4):315–317, July 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900980.

Halfon:1986:MFM

[Hal86b] Efraim Halfon. Modelling the fate of mirex and lindane in Lake Ontario, off the Niagara river mouth. Ecological Modelling, 33(1):13–33, September 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086901055.

Hall:1988:ASH

[Hal88a] Charles A. S. Hall. An assessment of several of the historically most influential theoretical models used in ecology and of the data provided in their support. Ecological Modelling, 43(1–2):5–31, October 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900701.

Hall:1988:WCG

[Hal88b] Charles A. S. Hall. What constitutes a good model and by whose criteria? Ecological Modelling, 43(1–2):125–127, October 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900762.

Halfon:1989:MEM

[Hal89a] Efraim Halfon. Microcomputers in ecological modelling: a special issue dedicated to research, education and computer graphics. Ecological Modelling, 47(1–2):3–6, September 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089901051.

Halfon:1989:PVC

[Hal89b] Efraim Halfon. Probabilistic validation of computer simulations using the bootstrap. Ecological Modelling, 46(3–4):213– 219, August 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900185.

Harrison:1984:AUS

[Har84] Gary W. Harrison. Analysis of uncertainty in a salt marsh carbon cycling model: a comparison of methods. Ecological Modelling, 24(3–4):207–220, September 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900413.

Hargrave:1985:PSO

[Har85] B. T. Hargrave. Particle sedimentation in the ocean. Ecological Modelling, 30(3–4):229–246, December 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900699.

Harris:1984:PMD

[HBB^{+}84] J. R. W. Harris, A. J. Bale, B. L. Bayne, R. F. C. Mantoura, A. W. Morris, L. A. Nelson, P. J. Radford, R. J. Uncles, S. A. Weston, and J. Widdows. A preliminary model of the dispersal and biological effect of toxins in the Tamar Estuary, England. Ecological Modelling, 22(1–4):253–284, April 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900164.

Hallam:1983:ETP

[HCL83] T. G. Hallam, C. E. Clark, and R. R. Lassiter. Effects of toxicants on populations: a qualitative approach I. Equilibrium environmental exposure. Ecological Modelling, 18(3–4):291– 304, April 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900194.

Harwell:1981:ATC

[HCR81] Mark A. Harwell, Wendell P. Cropper, and Harvey L. Ragsdale. Analyses of transient characteristics of a nutrient cycling model. Ecological Modelling, 12(1–2):105–131, March 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900272.

Halfon:1989:CPD

[HdJ89] Efraim Halfon and David de Jong. A computer program to display animations within the computer graphics halo environment. Ecological Modelling, 47(1–2):153–160, September 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089901142.

Hallam:1984:EPM

[HdL84] T. G. Hallam and J. T. de Luna. Extinction and persistence in models of population-toxicant interactions. Ecological Modelling, 22(1–4):13–20, April 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900048.

Herendeen:1988:NTD

[Her88a] Robert A. Herendeen. Network trophic dynamics. Ecological Modelling, 42(1):75–78, July 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900932.

Herendeen:1988:RME

[Her88b] Robert A. Herendeen. Role of models in ecology. Ecological Modelling, 43(1–2):133–136, October 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900786.

Herendeen:1989:EIR

[Her89] Robert Herendeen. Energy intensity, residence time, exergy, and ascendency in dynamic ecosystems. Ecological Modelling, 48(1–2):19–44, October 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900586.

Harrison:1980:RNC

[HF80] Gary W. Harrison and Stephan Fekete. Resistance of nutrient cycling systems to perturbations of the flow rates. Ecological Modelling, 10(3–4):227–241, August 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900617.

Hahn:1983:DMA

[HF83] Brian D. Hahn and Peter R. Furniss. A deterministic model of an anthrax epizootic: threshold results. Ecological Modelling, 20(2–3):233–241, November 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900091.

Herrick:1982:MMC

[HGG^{+}82] Christopher J. Herrick, Erik D. Goodman, Charles A. Guthrie, Richard H. Blythe, Glenn A. Hendrix, Richard L. Smith, and Jim E. Galloway. A model of mercury contamination in a woodland stream. Ecological Modelling, 15(1):1–28, February 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900655.

Huang:1986:ACC

[HH86] Shu-Li Huang and Mei-Chuen Huang. Applied carrying capacity concept for integrating stormwater management and land use planning, a case study: the Kuantu plain of Taipei, Taiwan. Ecological Modelling, 33(1):35–58, September 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086901067.

Halfon:1989:APH

[HHM89] Efrain Halfon, Jo-Ann Hodson, and Karon Miles. An algorithm to plot Hasses diagrams on microcomputers and Cal-Comp plotters. Ecological Modelling, 47(1–2):189–197, September 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089901178.

Hakamata:1986:IST

[HHS^{+}86] Tomoyuki Hakamata, Syota Hirosaki, Yasuo Sekine, Yuzuru Suzuki, and Sumio Kato. Interactive software tools, BGS-II and BGS-III, for ecological simulation. Ecological Modelling, 32(1–3):71–84, June 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900177.

Hearne:1988:MNR

[HHW88] John W. Hearne and Clive Howard-Williams. Modelling nitrate removal by riparian vegetation in a springfed stream: the influence of land-use practices. Ecological Modelling, 42 (3–4):179–198, September 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380088900567.

Helton:1985:SAA

[HIB85] Jon C. Helton, Ronald L. Iman, and Jack B. Brown. Sensitivity analysis of the asymptotic behavior of a model for the environmental movement of radionuclides. Ecological Modelling, 28(4):243–278, August 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900778.

Higashi:1986:EIO

[Hig86a] Masahiko Higashi. Extended input-output flow analysis of ecosystems. Ecological Modelling, 32(1–3):137–147, June 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900220.

Higashi:1986:RTC

[Hig86b] Masahiko Higashi. Residence time in constant compartmental ecosystems. Ecological Modelling, 32(4):243–250, July 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008690089X.

Hippe:1983:EAL

[Hip83] Peter W. Hippe. Environ analysis of linear compartmental systems: the dynamic, time-invariant case. Ecological Modelling, 19(1):1–26, May 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900674.

Hirata:1985:EBI

[Hir85] Hironori Hirata. Equivalence between input-output analysis and environ analysis as concerns flow partitions. Ecological Modelling, 30(1–2):3–12, October 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900341.

Holm:1984:CGS

[HK84] Juhani Holm and Seppo Kellomäki. Comparison of growth strategies of two competing plant species in forest ground cover. Ecological Modelling, 23(1–2):135–150, May 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084901224.

Huillet:1985:SPW

[HL85] Thierry Huillet and Jacques Lauga. A soil-plant-water model with a case study in a forested catchment. Ecological Modelling, 27(3–4):235–250, April 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900055.

Hughes:1988:SHY

[HM88] G. Hughes and R. G. McKinlay. Spatial heterogeneity in yield-pest relationships for crop loss assessment. Ecological Modelling, 41(1–2):67–73, April 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900452.

Hofmann:1984:NSD

[Hof84] Eileen E. Hofmann. North sea dynamics: J. Sündermann and W. Lenz (Editors). Springer-Verlag, Berlin/Heidelberg/ New York, 1983. 693 pp., Cloth DM98.00/US$42.30. ISBN 3-540-12013-0. Ecological Modelling, 24(3–4):305–307, September 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900486.

Hart:1986:NTM

[HON86] B. T. Hart, E. M. Ottaway, and B. N. Noller. Nutrient and trace metal fluxes in the Magela Creek system, Northern Australia. Ecological Modelling, 31(1–4):249–265, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900670.

Herman:1983:NMD

[HP83] Alex W. Herman and Trevor Platt. Numerical modelling of diel carbon production and zooplankton grazing on the Scotian shelf based on observational data. Ecological Modelling, 18(1):55–72, January 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900753.

Higashi:1986:FAA

[HP86] Masahiko Higashi and Bernard C. Patten. Further aspects of the analysis of indirect effects in ecosystems. Ecological Modelling, 31(1–4):69–77, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900566.

Hanson:1985:PGP

[HPI85] J. D. Hanson, W. J. Parton, and G. S. Innis. Plant growth and production of grassland ecosystems: a comparison of modelling approaches. Ecological Modelling, 29(1–4):131–144, September 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008590050X.

Hofmann:1980:TDM

[HPKA80] Eileen E. Hofmann, Leonard J. Pietrafesa, John M. Klinck, and Larry P. Atkinson. A time-dependent model of nutrient distribution in continental shelf waters. Ecological Modelling, 10(3–4):193–214, August 1980. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380080900599.

Henderson-Sellers:1981:RSB

[HS81] B. Henderson-Sellers. Rates of spread for a bent-over chimney plume. Ecological Modelling, 11(4):253–263, February 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900612.

Harmsen:1984:TMF

[HS84a] Rudolf Harmsen and Bonnie Sibbald. Testing model feasibility and sufficiency through laboratory simulation: a case study. Ecological Modelling, 21(3):161–174, February 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084901352.

Henderson-Sellers:1984:DAU

[HS84b] B. Henderson-Sellers. Development and application of “U.S.E.D.”: a hydroclimate lake stratification model. Ecological Modelling, 21(4):233–246, March 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900619.

Henderson-Sellers:1987:PRM

[HS87] B. Henderson-Sellers. Plume rise modelling: the effects of including a wind shear and a variable surface roughness. Ecological Modelling, 37(3–4):269–286, July 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900299.

Henderson-Sellers:1985:VPR

[HSA85] B. Henderson-Sellers and S. E. Allen. Verification of the plume rise/dispersion model USPR: plume rise for single stack emissions. Ecological Modelling, 30(3–4):209–227, December 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900687.

Hanson:1988:MSM

[HSP88] J. D. Hanson, J. W. Skiles, and W. J. Parton. A multispecies model for rangeland plant communities. Ecological Modelling, 44(1–2):89–123, December 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900841.

Hirata:1986:LSS

[HU86] Hironori Hirata and Robert E. Ulanowicz. Large-scale systems perspective on ecological modelling and analysis. Ecological Modelling, 31(1–4):79–104, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900578.

Hueting:1987:ESG

[Hue87] Roefie Hueting. An economic scenario that gives top priority to saving the environment. Ecological Modelling, 38(1–2):123–140, September 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900482.

Hulburt:1985:ULE

[Hul85] Edward M. Hulburt. Use of logical equivalence in modeling ecological relations of oceanic phytoplankton. Ecological Modelling, 27(1–2):25–43, March 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900237.

Hulburt:1989:EU

[Hul89] Edward M. Hulburt. Equivalence and its use. Ecological Modelling, 45(1):1–26, February 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900975.

Huson:1982:GAM

[Hus82] L. W. Huson. A graphical aid to multivariate sensitivity analysis. Ecological Modelling, 16(2–4):91–98, August 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900035.

Huson:1984:DPC

[Hus84] L. W. Huson. Definition and properties of a coefficient of sensitivity for mathematical models. Ecological Modelling, 21(3):149–159, February 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084901340.

Ichikawa:1986:MER

[IA86] Arata Ichikawa and Koji Amano. Modeling for estimating representative water conditions in a river basin. Ecological Modelling, 32(1–3):199–213, June 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008690027X.

Iwasa:1987:AME

[IAL87] Yoh Iwasa, Viggo Andreasen, and Simon Levin. Aggregation in model ecosystems. I. Perfect aggregation. Ecological Modelling, 37(3–4):287–302, July 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900305.

Istvan:1981:CMP

[IDS81] Bogardi Istvan, Lucien Duckstein, and Ferenc Szidarovszky. A control model for phosphorus loading reduction under certainty. Ecological Modelling, 12(1–2):83–103, March 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900260.

Innis:1981:SMM

[IHH81] George S. Innis, David F. Hanson, and James W. Haefner. A simulation model of management alternatives in a freshwater fishery. Ecological Modelling, 12(4):267–280, May 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900430.

Iriberri:1985:HBA

[IUME85] J. Iriberri, A. Undurraga, A. Muela, and L. Egea. Heterotrophic bacterial activity in coastal waters: Functional relationship of temperature and phytoplankton population. Ecological Modelling, 28(1–2):113–120, May 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008590016X.

Ikeda:1980:FPD

[IY80] Saburo Ikeda and Takashi Yokoi. Fish population dynamics under nutrient enrichment — a case of the East Seto Inland Sea. Ecological Modelling, 10(3–4):141–165, August 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380080900575.

Jacobsen:1987:CWP

[Jac87] Ole Stig Jacobsen. Coastal wetlands: Proceedings of the First Great Lakes Coastal Wetlands Colloquium, 5–7 November 1984, East Lansing, MI. Harold H. Prince and Frank D’Itri (Editors). Lewis Publishers, Inc., Chelsea, MI, U. S. A., 1985. Marketed and distributed by John Wiley & Sons Limited, Chichester, Great Britain, 286 pp., £36.75. ISBN 0-87371-052-5. Ecological Modelling, 38(3–4):309–310, October 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087901037.

Jameson:1987:MMH

[Jam87] D. A. Jameson. Mathematical methods in human geography and planning: A. G. Wilson and R. J. Bennett. Wiley, Chichester, 1985. 411 pp., sterling 29.95. ISBN 0-471-10521-X. Ecological Modelling, 37(3–4):324–325, July 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900366.

Jameson:1987:SOM

[JB87] Donald A. Jameson and E. Thomas Bartlett. Selection of optimal management strategies based on stochastic dynamic ecological models. Ecological Modelling, 36(1–2):5–13, April 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900548.

Janecek:1989:MSP

[JBL^{+}89] A. Janecek, G. Benderoth, M. K. B. Lüdeke, J. Kindermann, and G. H. Kohlmaier. Model of the seasonal and perennial carbon dynamics in deciduous-type forests controlled by climatic variables. Ecological Modelling, 49(1–2):101–124, December 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008990046X.

Johnson:1985:NAM

[JCN85] Douglas H. Johnson, Michael J. Conroy, and James D. Nichols. The need for accuracy in modelling: an example.

Ecological Modelling, 30(1–2):157–161, October 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900419.

Jensen:1982:IOT

[Jen82] A. L. Jensen. Impact of a once-through cooling system on the yellow perch stock in the Western Basin of Lake Erie. Ecological Modelling, 15(2):127–144, March 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900576.

Jensen:1986:CUF

[Jen86] A. L. Jensen. Contaminant uptake by fish and the potential for transfer to humans modelled over time. Ecological Modelling, 32(4):281–290, July 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008690092X.

Jensen:1987:SME

[Jen87] A. L. Jensen. Simple models for exploitative and interference competition. Ecological Modelling, 35(1–2):113–121, February 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900937.

Jensen:1989:SPL

[Jen89] A. L. Jensen. Simulation of the potential for life history components to regulate walleye population size. Ecological Modelling, 45(1):27–41, February 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900987.

Jorgensen:1986:SLA

[JG86] S. E. Jørgensen and Takeshi Goda. Scope and limit in the application of ecological models to environmental management — I–VI. Ecological Modelling, 32(1–3):237–240, June 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008690030X.

Jager:1988:SEI

[JG88] Henriette I. Jager and Robert H. Gardner. A simulation experiment to investigate food web polarization. Ecological Modelling, 41(1–2):101–116, April 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900488.

Jiagang:1988:TMP

[Jia88] Liu Jiagang. A theoretical model of the process of rainfall interception in forest canopy. Ecological Modelling, 42(2):111–123, August 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088901111.

Jorgensen:1981:PEE

[JJKNM81] S. E. Jørgensen, L. A. Jørgensen, L. Kamp-Nielsen, and H. F. Mejer. Parameter estimation in eutrophication modelling. Ecological Modelling, 13(1–2):111–129, June 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900090.

Jorgensen:1986:VPB

[JKNC^{+}86] Sven Erik Jørgensen, Lars Kamp-Nielsen, Torben Christensen, Jørgen Windolf-Nielsen, and Brynjolf Westergaard. Validation of a prognosis based upon a eutrophication model. Ecological Modelling, 32(1–3):165–182, June 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900244.

Jorgensen:1986:EGE

[JKNJ86] Sven Erik Jørgensen, Lars Kamp-Nielsen, and L. A. Jørgensen. Examination of the generality of eutrophication models. Ecological Modelling, 32(4):251–266, July 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900906.

Jorgensen:1982:CSC

[JKNM82] S. E. Jørgensen, L. Kamp-Nielsen, and H. F. Mejer. Comparison of a simple and a complex sediment phosphorus model.

Ecological Modelling, 16(2–4):99–124, August 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900047.

Jolankai:1984:AOM

[Jol84] Géza Jolánkai. Aquatische ökosystem mmodellierung und simulation: Milan Straškraba and Albrecht Gnauck. VEB Gustav Fischer Verlag Jena, Jena, German Democratic Republic, 1983. 279 pp., DM58.00. Ecological Modelling, 26(3–4):316–318, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900772.

Jorgensen:1980:CME

[Jør80a] Sven Erik Jørgensen. A coastal marine ecosystem, simulation and analysis: J. N. Kremer and S. W. Nixon. Springer-Verlag, Berlin–Heidelberg–New York, 80 figs., 217 pp., US $32.00. Ecological Modelling, 10(1):63–64, June 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900794.

Jorgensen:1980:MES

[Jør80b] Sven Erik Jørgensen. Marsh-estuarine systems simulation: Marsh-Estuarine systems simulation. Belle W. Baruch Library in Marine Science No. 8. Richard F. Dame (Editor), The University of South Carolina Press, 1979, 108 figs, 260 pp. Ecological Modelling, 11(3):227–228, December 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900848.

Jorgensen:1980:MMW

[Jør80c] Sven Erik Jørgensen. Mathematical modelling in water pollution control: A. James (Editor). John Wiley and Sons, 1978, Chichester — New York — Brisbane — Toronto, U. S. $31.00. Ecological Modelling, 11(3):228–229, December 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008090085X.

Jorgensen:1980:EMR

[Jør80d] Sven Erik Jørgensen. Ecological modelling — a review of editorial policy. Ecological Modelling, 11(2):79–80, November 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380080900873.

Jorgensen:1981:CFW

[Jør81a] L. A. Jørgensen. Comparison of forest water and energy exchange models. Proceedings of an IUFRO workshop, Uppsala (Sweden), September 24–30, 1978: Swen Halldin (Editor). ISEM, Copenhagen and Elsevier, Amsterdam, 1978, 269 pp. Ecological Modelling, 11(4):294, February 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008190065X.

Jorgensen:1981:BWT

[Jør81b] Sven Erik Jørgensen. Basic water treatment for application worldwide: by George smethurst. Thomas telford, London, 1979, 216 pp., illustrated. Ecological Modelling, 12(3):210– 211, April 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900594.

Jorgensen:1981:SAR

[Jør81c] Sven Erik Jørgensen. Simulation of assimilation, respiration and transpiration of crops: C. T. de Wit et al. Wageningen Centre for Agricultural Publishing and Documentation, Wageningen, 1978, 141 pp., 36 figs. Ecological Modelling, 11(4):291–292, February 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900636.

Jorgensen:1981:TSE

[Jør81d] Sven Erik Jørgensen. Theoretical systems ecology: Efraim Halfon (Editor). Academic Press, New York–San Francisco– London, 1979, 516 pp., illustrated, U.S. $43.00, ISBN 0-12-318750-8. Ecological Modelling, 12(3):209–210, April 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900582.

Jorgensen:1982:SEH

[Jør82a] S. E. Jørgensen. Systems ecology, by H. H. Shugart and R. V. O’Neill (Editors). Bechmark Papers in Ecology/9, Dowden, Hutchinson and Ross, Stroudsberg, PA, 1979. 368 pp., illustrated, US $29.50, ISBN 0-87933-347-2. Ecological Modelling, 15(1):79–80, February 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900692.

Jorgensen:1982:EN

[Jør82b] S. E. Jørgensen. Editorial note. Ecological Modelling, 17(2):181, October 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900540.

Jorgensen:1982:EEW

[Jør82c] Sven Erik Jørgensen. Ecological effects of waste water: E. B. Welch, with hydrographic characteristics by T. Lindell. Cambridge University Press, 1980, 377 pp., 157 figs., £20.00 (hb), £7.25 (pb), ISBN 0-521-22495-0 (hb)/0-521-29525-4 (pb). Ecological Modelling, 15(2):185–186, March 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900618.

Jorgensen:1982:E

[Jør82d] Sven Erik Jørgensen. Editorial. Ecological Modelling, 14 (3–4):153, January 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900151.

Jorgensen:1983:EEW

[Jør83] Sven Erik Jørgensen. Ecological effects of waste water: E. B. Welch, with hydrographic characteristics by T. Lindell. 1980, 337 pp., 157 figs., ISBN: 0-521-22495-0. Ecological Modelling, 18(3–4):307–308, April 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900212.

Jorgensen:1984:E

[Jør84a] S. E. Jørgensen. Editorial. Ecological Modelling, 21(4):i–ii, March 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900590.

Jorgensen:1984:Ia

[Jør84b] S. E. Jørgensen. Introduction. Ecological Modelling, 22(1–4): vii, April 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900024.

Jorgensen:1984:Ib

[Jør84c] S. E. Jørgensen. Introduction. Ecological Modelling, 23 (1–2):vii–viii, May 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084901157.

Jorgensen:1984:PET

[Jør84d] S. E. Jørgensen. Parameter estimation in toxic substance models. Ecological Modelling, 22(1–4):1–11, April 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900036.

Jorgensen:1986:E

[Jør86a] S. E. Jørgensen. Editorial. Ecological Modelling, 33(2–4):83–84, October 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900323.

Jorgensen:1986:PEI

[Jør86b] Sven Erik Jørgensen. Perspective on environmental impact assessment: Brian D. Clark and Alexander Gilad, Ronald Bisset and Paul Tomlinson (Editors), Reidel, Dordrecht, The Netherlands, 1984. 520 pp., Dfl. 180.00. ISBN 90-277-1753-2. Ecological Modelling, 34(1–2):134–135, November 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900840.

Jorgensen:1986:RSS

[Jør86c] Sven Erik Jørgensen. Remote sensing of shelf sea hydrodynamics: Proc. 15th International Liège Colloquium on Ocean Hydrodynamics. Jacques C.J. Nihoul (Editor). Elsevier Oceanography Series, 38. Elsevier, Amsterdam/Oxford/New York/Tokyo, 1984. xii + 354 pp., Dfl.180. ISBN 0-44-42314-1. Ecological Modelling, 33(1):80, September 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086901109.

Jorgensen:1986:RTV

[Jør86d] Sven Erik Jørgensen. The role of terrestrial vegetation in the global carbon cycle: Measurement by remote sensing: G. M. Woodwell (Editor). SCOPE 23, Wiley, Chichester, Great Britain, 1984. 274 pp., £27.50. ISBN 0-471-90262-4. Ecological Modelling, 34(1–2):133–134, November 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900839.

Jorgensen:1986:DEC

[Jør86e] Sven Erik Jørgensen. Disturbance and ecosystems: Components of response: H. A. Mooney and M. Godron (Editors), Ecological Studies, 44. Springer, Berlin, 1983. xvi + 292 pp., + 82 figures, Cloth DM 128.00 ISBN 3-540-12454-3. Ecological Modelling, 34(1–2):136–137, November 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900852.

Jorgensen:1986:SDM

[Jør86f] Sven Erik Jørgensen. Structural dynamic model. Ecological Modelling, 31(1–4):1–9, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900517.

Jorgensen:1987:AAB

[Jør87a] S. E. Jørgensen. Advances in applied biology: vol. 9: T. H. Coaker (Editor). Academic Press Inc., Orlando, FL., U. S. A., 1983. 338 pp., US $59.50. ISBN 0-12-040909-7. Ecological Modelling, 36(3–4):317, May 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900743.

Jorgensen:1987:AME

[Jør87b] S. E. Jørgensen. Advances in microbial ecology, vol. 8: K. C. Marshall (Editor). Plenum, New York, 1985. 307 pp., US$45.00. ISBN 0-306-41877-0. Ecological Modelling, 37(3– 4):317–318, July 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380087900329.

Jorgensen:1987:DSR

[Jør87c] S. E. Jørgensen. Databases in systematics: R. Allkin and F. A. Bisby (Editors). Systematics Association Special Volume Series, Vol. 26. Academic Press, London/Orlando, FL, 1984. 352 pp., £29.50. ISBN 0-12-053040-6. Ecological Modelling, 35(1–2):149, February 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900962.

Jorgensen:1987:ENR

[Jør87d] S. E. Jørgensen. Ecology of natural resources: F. Ramade. Wiley, Chichester, Great Britain, 1984. 231 pp., Paperback £14.95/Cloth £34.00/US $49.00. ISBN Paperback 0-471-90625-5/Cloth 0-471-90104-0. Ecological Modelling, 36 (3–4):318–319, May 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380087900755.

Jorgensen:1987:PBL

[Jør87e] S. E. Jørgensen. Pollination biology: Leslie Real (Editor). Academic Press, Orlando, FL, 1983. 330 pp., hardcover US$49.00/softcover US$35.00. ISBN 0-12-583980-4/0-12-583982-0. Ecological Modelling, 37(3–4):318–320, July 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900330.

Jorgensen:1987:PHR

[Jør87f] S. E. Jørgensen. Principles of health risk assessment: Prentice-Hall Inc., Englewood Cliffs, NJ, U. S. A., 1985. 417 pp. ISBN 0-13-709197-4. Ecological Modelling, 36(3–4):319–320, May 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900767.

Jorgensen:1988:UME

[Jør88] Sven Erik Jørgensen. Use of models as experimental tool to show that structural changes are accompanied by increased exergy. Ecological Modelling, 41(1–2):117–126, April 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008890049X.

Joyeux:1985:TCE

[Joy85] Roselyne Joyeux. Fur trade in Canada: an econometric analysis. Ecological Modelling, 27(1–2):139–152, March 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900298.

Johnson:1987:MPM

[JSC87] Douglas H. Johnson, Donald W. Sparling, and Lewis M. Cowardin. A model of the productivity of the mallard duck. Ecological Modelling, 38(3–4):257–275, October 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087901001.

Jernigan:1980:LSM

[JT80] R. W. Jernigan and C. P. Tsokos. A linear stochastic model for phytoplankton production in a marine ecosystem. Ecological Modelling, 10(1):1–12, June 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900757.

Jenq:1983:LPM

[JUGH83] Tzay-Rong Jenq, Christopher G. Uchrin, Marvin L. Granstrom, and Shing-Fu Hsueh. A linear program model for pointnonpoint source control decisions: Theoretical development.

Ecological Modelling, 19(4):249–262, September 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008390042X.

Kercher:1984:APS

[KA84] J. R. Kercher and M. C. Axelrod. Analysis of SILVA: a model for forecasting the effects of SO_{2} pollution and fire on western coniferous forests. Ecological Modelling, 23(1–2):165–184, May 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901248.

Kamykowski:1981:SSC

[Kam81] Daniel Kamykowski. The simulation of a Southern California red tide using characteristics of a simultaneously-measured internal wave field. Ecological Modelling, 12(4):253–265, May 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900429.

Karlson:1981:SSG

[Kar81] Ronald H. Karlson. A simulation study of growth inhibition and predator resistance in Hydractinia echinata. Ecological Modelling, 13(1–2):29–47, June 1981. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380081900065.

Karlson:1985:COI

[Kar85] Ronald H. Karlson. Competitive overgrowth in interactions among sessile colonial invertebrates: a comparison of stochastic and phenotypic variation. Ecological Modelling, 27(3–4):299– 312, April 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900092.

Kaufmann:1987:BME

[Kau87] Robert Kaufmann. Biophysical and Marxist economics: Learning from each other. Ecological Modelling, 38(1–2):91–105, September 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900469.

Kawanishi:1986:NAFa

[Kaw86a] Hiroshi Kawanishi. Numerical analysis of forest temperature. I. Diurnal variations. Ecological Modelling, 33(2–4):315–327, October 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900463.

Kawanishi:1986:NAFb

[Kaw86b] Hiroshi Kawanishi. Numerical analysis of forest temperature. II. Seasonal variations. Ecological Modelling, 33(2–4):329–340, October 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900475.

Kalmaz:1981:MMC

[KB81] Ekrem V. Kalmaz and JoséL. Barbieri. Mathematical modeling and computer simulation of radioactive and toxic chemical species dispersion in porous media in the vacinity of uranium recovery operations. Ecological Modelling, 13(3):159–181, August 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900508.

Karlson:1984:CDL

[KB84] Ronald H. Karlson and Leo W. Buss. Competition, disturbance and local diversity patterns of substratum-bound clonal organisms: a simulation. Ecological Modelling, 23(3):243–255, June 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901030.

Kreikenbohm:1988:BC

[KB88] R. Kreikenbohm and E. Bohl. Bistability in the chemostat. Ecological Modelling, 43(3–4):287–301, November 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900099.

Koschel:1987:MAE

[KBPR87] Rainer Koschel, Jürgen Benndorf, Gottfried Proft, and Frieder Recknagel. Model-assisted evaluation of alternative hypotheses to explain the self-protection mechanism of lakes due to calcite precipitation. Ecological Modelling, 39(1–2):59–65, November 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900135.

Kozerski:1988:ILE

[KBSM88] Hans-Peter Kozerski, Horst Behrendt, Günter Schellenberger, and Volker Mohaupt. Investigations of the lake ecosystem model emsy by means of the simulation system sonches. Ecological Modelling, 41(3–4):193–200, June 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900270.

Kozerski:1982:TLA

[KD82] Hans-Peter Kozerski and Marie Dvořáková. A three-layer aquatic ecosystem model: sediment submodel tests and some simulations. Ecological Modelling, 17(2):147–156, October 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900515.

Klihamer:1983:IPB

[KD83] Peter G. L. Klihamer and Tom J. De Jong. Is it profitable for biennials to live longer than two years? Ecological Modelling, 20(2–3):223–232, November 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008390008X.

Kempf:1984:ROO

[KDC84] J. Kempf, L. Duckstein, and J. Casti. Relaxation oscillations and other non-michaelian behavior in a slow-fast phytoplankton growth model. Ecological Modelling, 23(1–2):67– 90, May 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901194.

Kowalik:1984:WTS

[KE84] Piotr J. Kowalik and Henrik Eckersten. Water transfer from soil through plants to the atmosphere in willow energy forest. Ecological Modelling, 26(3–4):251–284, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900723.

Kalmaz:1985:PKM

[KEK85] E. E. Kalmaz, A. H. Eraslan, and K. H. Kim. A preliminary kinetics model predicting concentration variations of hypobromous acid and bromate in ozonated marine water. Ecological Modelling, 29(1–4):315–326, September 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900596.

Kercher:1983:CFS

[Ker83] James R. Kercher. Closed-form solutions to sensitivity equations in the frequency and time domains for linear models of ecosystems. Ecological Modelling, 18(3–4):209–221, April 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380083900133.

Krauthamer:1987:SMT

[KGG87] Judith T. Krauthamer, William E. Grant, and Wade L. Griffin. A sociobioeconomic model: the Texas inshore shrimp fishery. Ecological Modelling, 35(3–4):275–307, March 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087901165.

Kawasaki:1986:ESN

[KGHM86] S. Kawasaki, M. J. Gromiec, K. Hanaki, and J. Matsumoto. Effect of seawater on nitrification by attached biofilm. Ecological Modelling, 32(1–3):183–190, June 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900256.

Knudsen:1987:UCS

[KH87a] Guy R. Knudsen and George W. Hudler. Use of a computer simulation model to evaluate a plant disease biocontrol agent. Ecological Modelling, 35(1–2):45–62, February 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900901.

Kutas:1987:ELR

[KH87b] Tibor Kutas and Sándor Herodek. Effects of load reductions on the water quality of a large shallow lake. Ecological Modelling, 39(1–2):85–99, November 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900159.

Kadlec:1988:MNB

[KH88] Robert H. Kadlec and David E. Hammer. Modeling nutrient behavior in wetlands. Ecological Modelling, 40(1):37–66, January 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088901019.

King:1988:EOM

[KHF88] David A. King, Allen S. Heagle, and Richard B. Flagler. Evaluation of an ozone × moisture stress interaction model for soybean. Ecological Modelling, 41(3–4):269–279, June 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900324.

Katerji:1986:MDP

[KHMBD86] Nader Katerji, Marc Hallaire, Yvette Menoux-Boyer, and Brigitte Durand. Modelling diurnal patterns of leaf water potential in field conditions. Ecological Modelling, 33 (2–4):185–203, October 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380086900402.

Kishi:1986:PDR

[KI86] Michio Kishi and Saburo Ikeda. Population dynamics of ‘red tide’ organisms in eutrophicated coastal waters — numerical experiment of phytoplankton bloom in the East Seto Inland Sea, Japan. Ecological Modelling, 31(1–4):145–174, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008690061X.

King:1987:MPI

[Kin87] David A. King. A model for predicting the influence of moisture stress on crop losses caused by ozone. Ecological Modelling, 35(1–2):29–44, February 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900895.

Kirchner:1989:TZI

[Kir89] Thomas B. Kirchner. TIME-ZERO: the integrated modeling environment. Ecological Modelling, 47(1–2):33–52, September 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089901087.

Kremer:1982:TTL

[KK82] James N. Kremer and Patricia Kremer. A three-trophic level estuarine model: Synergism of two mechanistic simulations. Ecological Modelling, 15(2):145–157, March 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900588.

Koschel:1984:MUV

[KK84] Rainer Koschel and Peter Kasprzak. Model use and verification of ecological parameters for an oligotrophic lake (Lake Stechlin, G.D.R.). Ecological Modelling, 26(1–2):97–101, December 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900966.

Krajewski:1982:GTR

[KKG82] Witold F. Krajewski, Andrzej K. Kraszewski, and William J. Grenney. A graphical technique for river water temperature predictions. Ecological Modelling, 17(3–4):209–224, December 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900321.

Kauppi:1986:AFS

[KKP^{+}86] Pekka Kauppi, Juha Kämäri, Maximilian Posch, Lea Kauppi, and Egbert Matzner. Acidification of forest soils: Model development and application for analyzing impacts of acidic deposition in Europe. Ecological Modelling, 33(2–4):231–253, October 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900426.

Krysanova:1989:SMC

[KMRV89] V. Krysanova, A. Meiner, J. Roosaare, and A. Vasilyev. Simulation modelling of the coastal waters pollution from agricultural watershed. Ecological Modelling, 49(1–2):7–29, December 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900410.

Knijnenburg:1984:CUI

[KMW84] Arnd Knijnenburg, Eberhard Matthäus, and Volker Wenzel. Concept and usage of the interactive simulation system for ecosystems sonches. Ecological Modelling, 26(1–2):51–76, December 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900942.

Kohlmaier:1989:CSN

[Koh89] G. H. Kohlmaier. Climate shocks: Natural and anthropogenic: K. Ya. Kondratyev. Wiley, New York, 1988. 296 pp. $47.50 ISBN 0-471-83019-4. Ecological Modelling, 47(3–4):314–316, September 15, 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900100.

Kalceva:1982:OME

[KOSS82] Rosica Kalčeva, Jiří Outrata, Zdeněk Schindler, and Milan Straškraba. An optimization model for the economic control of reservoir eutrophication. Ecological Modelling, 17(2):121–128, October 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900497.

Kuuluvainen:1989:SWT

[KP89] Timo Kuuluvainen and Timo Pukkala. Simulation of withintree and between-tree shading of direct radiation in a forest canopy: effect of crown shape and sun elevation. Ecological Modelling, 49(1–2):89–100, December 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900458.

Kindlmann:1982:NSS

[KR82] Pavel Kindlmann and Marcel Rejmánek. Number of species at stable equilibrium of complex model ecosystems. Ecological Modelling, 16(2–4):85–90, August 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900023.

Khanbilvardi:1986:MSE

[KR86] R. M. Khanbilvardi and A. S. Rogowski. Modeling soil erosion, transport and deposition. Ecological Modelling, 33 (2–4):255–268, October 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380086900438.

Kremer:1983:EIP

[Kre83] James N. Kremer. Ecological implications of parameter uncertainty in stochastic simulation. Ecological Modelling, 18 (3–4):187–207, April 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380083900121.

Kaluzny:1985:SEC

[KS85] Stephen Kaluzny and Gordon Swartzman. Simulation experiments comparing alternative process formulations using a factorial design. Ecological Modelling, 28(3):181–200, June 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900821.

Kmet:1989:GBG

[KS89a] T. Kmeť and M. Straškraba. Global behavior of a generalized aquatic ecosystem model. Ecological Modelling, 45(2):95– 110, April 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900860.

Krivan:1989:AGR

[KS89b] Vlastimil Křivan and Jaromír Seďa. Application of a guaranteed regression model to trophic interaction in an aquatic system. Ecological Modelling, 49(1–2):1–6, December 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900409.

Kohlmaier:1984:DDD

[KSB^{+}84] G. H. Kohlmaier, E. O. Siré, H. Bröhl, W. Kilian, U. Fischbach, M. Plöchl, T. Müller, and Jiang Yunsheng. Dramatic development in the dying of German spruce-fir forests: In search of possible cause-effect relationships. Ecological Modelling, 22(1–4):45–65, April 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900061.

Kozerski:1984:TCE

[KSBM84] Hans-Peter Kozerski, Günther Schellenberger, Horst Behrendt, and Volker Mohaupt. Testing of a complex ecological model for shallow water bodies. Ecological Modelling, 26(1–2):103–113, December 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900978.

Kot:1988:CCI

[KST^{+}88] M. Kot, W. M. Schaffer, G. L. Truty, D. J. Graser, and L. F. Olsen. Changing criteria for imposing order. Ecological Modelling, 43(1–2):75–110, October 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900749.

Kundu:1982:ISA

[KSW82] Satyansu S. Kundu, Gaylord V. Skogerboe, and Wynn R. Walker. Improvement and sensitivity analysis of the corngro model. Ecological Modelling, 16(2–4):209–239, August 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900096.

Kmet:1989:VMM

[KT89] Tibor Kmeť and Dezider Tóth. Verification of the mathematical model of nitrogen circulation with and without light access. Ecological Modelling, 46(3–4):135–146, August 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900148.

Kleeper:1988:DCC

[KvdK88] O. Kleeper and J. P. G. van de Kamer. A definition of the consistency of the carbon budget of an ecosystem, and its application to the oosterschelde estuary, S.W. Netherlands. Ecological Modelling, 42(3–4):217–232, September 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900580.

Kirchner:1984:VPS

[KW84] Thomas B. Kirchner and F. Ward Whicker. Validation of pathway, a simulation model of the transport of radionuclides through agroecosystems. Ecological Modelling, 22(1–4):21–44, April 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008490005X.

Kawamoto:1989:SMP

[KWSM89] H. Kawamoto, S. M. Woods, R. N. Sinha, and W. E. Muir. A simulation model of population dynamics of the rusty grain beetle, Cryptolestes ferrugineus in stored wheat. Ecological Modelling, 48(1–2):137–157, October 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900641.

Kachi:1986:SMD

[KYTS86] N. Kachi, Y. Yasuoka, T. Totsuka, and K. Suzuki. A stochastic model for describing revegetation following forest cutting: an application of remote sensing. Ecological Modelling, 32(1–3):105–117, June 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900207.

Logofet:1984:MMCa

[LA84a] D. O. Logofet and G. A. Alexandrov. Modelling of matter cycle in a mesotrophic bog ecosystem I. Linear analysis of carbon environs. Ecological Modelling, 21(4):247–258, March 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900620.

Logofet:1984:MMCb

[LA84b] D. O. Logofet and G. A. Alexandrov. Modelling of matter cycle in a mesotrophic bog ecosystem II. Dynamic model and ecological succession. Ecological Modelling, 21(4):259–276, March 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900632.

Lum:1981:ICE

[LAF81] Leighton W. K. Lum, Neal E. Armstrong, and E. Gus Fruh. Incorporation of chemical equilibrium, physico-chemical reactions and the simulation of inorganic qualities in a reservoir ecologic model. Ecological Modelling, 14(1–2):95–123, November 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900168.

Laitner:1987:RCN

[Lai87] Skip Laitner. Resource constraints: the need for community management in economic development strategies. Ecological Modelling, 38(1–2):159–170, September 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900500.

Lassey:1983:CSM

[Las83] Keith R. Lassey. Conceptually simple mathematical models of filtration (and exchange) processes. Ecological Modelling, 18(1):1–26, January 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380083900728.

See erratum [Ano83h].

Longstaff:1984:EMI

[LC84] B. C. Longstaff and W. R. Cuff. An ecosystem model of the infestation of stored wheat by Sitophilus oryzae: a reappraisal. Ecological Modelling, 25(1–3):97–119, October 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900358.

Legovic:1984:MMB

[Leg84a] Tarzan Legović. Mathematical models in biological oceanography: T. Platt, K. H. Mann and R. E. Ulanowicz (Editors). UNESCO Press, Paris, France, 1982. 156 pp., FF50.00. ISBN 92-3-101922-8. Ecological Modelling, 23(4):341–342, July 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084901285.

Legovic:1984:MBC

[Leg84b] Tarzan Legović. The major biochemical cycles and their interactions: R. Bolin and R. B. Cook (Editors). SCOPE 21. John Wiley and Sons, Chichester, Great Britain, 1983. xxi + 532 pp., £37.00. ISBN 0-471-10522-8. Ecological Modelling, 26(3–4):313–314, December 15, 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900747.

Legovic:1987:RIJ

[Leg87] Tarzan Legović. A recent increase in jellyfish populations: a predator–prey model and its implications. Ecological Modelling, 38(3–4):243–256, October 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900998.

Legovic:1989:RFO

[Leg89a] Tarzan Legović. The role of freshwater outflow in coastal marine ecosystems: Stig Skreslet (Editor). NATO ASI Series, Ecological Sciences, 7. Springer, Berlin, 1986. xi + 453 pp., hardcover, DM198.00. ISBN 3-540-16089-2. Ecological Modelling, 45(2):151–152, April 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900896.

Legovic:1989:PFW

[Leg89b] Tarzan Legović. Predation in food webs. Ecological Modelling, 48(3–4):267–276, November 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900513.

Leiler:1986:MCG

[Lei86] Igor Leiler. Material circulation and growth — with special reference to pollution problems. Ecological Modelling, 31(1–4):125–131, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380086900591.

Leohle:1983:ETC

[Leo83] Craig Leohle. Evaluation of theories and calculation tools in ecology. Ecological Modelling, 19(4):239–247, September 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380083900418.

Levine:1988:DFI

[Lev88] Stephen H. Levine. A dynamic formulation for inputoutput modelling of exploited ecosystems. Ecological Modelling, 44(1–2):143–151, December 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900865.

Lefkovitch:1985:SCH

[LF85] L. P. Lefkovitch and Lenore Fahrig. Spatial characteristics of habitat patches and population survival. Ecological Modelling, 30(3–4):297–308, December 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900729.

Lhotka:1987:WMM

[Lho87] L. Lhotka. Water management models in practice: a case study of the Aswan high dam: Dale Whittington and Giorgio Guariso. Developments in Environmental Modelling, Vol. 2. Elsevier, Amsterdam/Oxford/New York, 1983. 246 pp., Dfl. 150.00. ISBN 0-444-42156-4. Ecological Modelling, 35(3–4):309–310, March 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087901177.

Lauenroth:1986:EAN

[LHSS86] W. K. Lauenroth, H. W. Hunt, D. M. Swift, and J. S. Singh. Estimating aboveground net primary production in grasslands: a simulation approach. Ecological Modelling, 33(2–4):297–314, October 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900451.

Lijklema:1987:MCS

[Lij87] L. Lijklema. Methods of computing sedimentation in lakes and reservoirs: Stevan Bruk (Rapporteur). Unesco, Paris, 1985, 224 pp. Ecological Modelling, 38(3–4):310–312, October 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087901049.

Limburg:1985:ICE

[Lim85] Karin E. Limburg. Increasing complexity and energy flow in models of food webs. Ecological Modelling, 29(1–4):5–25, September 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900444.

Limburg:1982:CEF

[LJZ82] Karin E. Limburg, Ann-Mari Jansson, and James Zucchetto. A coastal ecosystem fisheries minimodel for the island of Gotland, Sweden. Ecological Modelling, 17(3–4):271–295, December 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900357.

Leps:1987:MDS

[LK87] Jan Lepš and Pavel Kindlmann. Models of the development of spatial pattern of an even-aged plant population over time. Ecological Modelling, 39(1–2):45–57, November 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900123.

Lorber:1984:MAW

[LMR84] M. N. Lorber, J. W. Mishoe, and P. R. Reddy. Modeling and analysis of waterhyacinth biomass. Ecological Modelling, 24(1–2):61–77, August 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900553.

Loehle:1985:OSS

[Loe85] Craig Loehle. Optimal stocking for semi-desert range: a catastrophe theory model. Ecological Modelling, 27(3–4):285–297, April 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900080.

Loehle:1987:AAI

[Loe87a] Craig Loehle. Applying artificial intelligence techniques to ecological modeling. Ecological Modelling, 38(3–4):191–212, October 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900974.

Loehle:1987:ECE

[Loe87b] Craig Loehle. Errors of construction, evaluation, and inference: a classification of sources of error in ecological models. Ecological Modelling, 36(3–4):297–314, May 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008790072X.

Loehr:1987:IHC

[Loe87c] J. Loehr. Impact of the hydrodynamic conditions on the primary production in an impounded river. Ecological Modelling, 39(3–4):227–245, December 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900020.

Loehle:1988:RPE

[Loe88] Craig Loehle. Robust parameter estimation for nonlinear models. Ecological Modelling, 41(1–2):41–54, April 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900439.

Loehle:1989:CTE

[Loe89] Craig Loehle. Catastrophe theory in ecology: a critical review and an example of the butterfly catastrophe. Ecological Modelling, 49(1–2):125–152, December 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900471.

Logofet:1981:ACD

[Log81] Dmitry O. Logofet. On the algorithmic congruence of a difference model and a differential one. Ecological Modelling, 12(4):297–299, May 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380081900454.

Logofet:1987:LMB

[Log87] Dmitri O. Logofet. Linear models in biology: Linear systems analysis with biological applications: M. R. Cullen. Ellis Horwood Limited, Chichester, Great Britain, 1985. 213 pp., £10.50. ISBN 0-85312-905-3. Ecological Modelling, 38(3–4):313–316, October 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087901062.

Loman:1984:CTM

[Lom84] Jon Loman. Comparing two models of alloregulated predators. Ecological Modelling, 24(3–4):293–303, September 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900474.

Loman:1985:ITS

[Lom85] Jon Loman. Influence of territoriality on the stability and coexistence of competing predators a simulation study.

Ecological Modelling, 27(1–2):95–108, March 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900274.

Loman:1986:UOI

[Lom86] Jon Loman. Use of overlap indices as competition coefficients: Tests with field data. Ecological Modelling, 34(3–4):231–243, December 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900062.

Loman:1988:APD

[Lom88] Jon Loman. Alternative prey that decreases vole population cyclicity: a simulation study based on field data. Ecological Modelling, 40(3–4):265–310, March 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008890021X.

Longstaff:1986:PPC

[Lon86] B. C. Longstaff. Pest and pathogen control: Strategic tactical and policy models: G. R. Conway (Editor). International Series on Applied Systems Analysis, 13. Wiley, Great Britain, 1984. xii + 488 pp., £32.00. ISBN 0-471-90349-3. Ecological Modelling, 32(4):311–313, July 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900955.

Longstaff:1987:PDR

[Lon87] Barry C. Longstaff. Population dynamics of rabies in wildlife: P. J. Bacon (Editor). Academic Press, London/Orlando, FL, U. S. A., 1985. xvii + 358 pp., £53.00/US$68.00 (also available in paperback £27.50/US$29.95). ISBN 0-12-071350-0. Ecological Modelling, 37(3–4):320–322, July 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900342.

Lonergan:1988:TMU

[Lon88a] Stephen C. Lonergan. Theory and measurement of unequal exchange: a comparison between a Marxist approach and an energy theory of value. Ecological Modelling, 41(1–2):127–145, April 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900506.

Longstaff:1988:TMM

[Lon88b] B. C. Longstaff. Temperature manipulation and the management of insecticide resistance in stored grain pests: a simulation study for the rice weevil, Sitophilus oryzae. Ecological Modelling, 43(3–4):303–313, November 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900105.

Levins:1983:SRA

[LP83] Richard A. Levins and Murl W. Parker. Short run aspects of a genetic control program for Heliothis virescens (Lepidoptera: Noctuidae). Ecological Modelling, 19(3):213–220, July 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900558.

Legovic:1984:HPP

[LP84] Tarzan Legović and Goran Perić. Harvesting population in a periodic environment. Ecological Modelling, 24(3–4):221–229, September 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900425.

Linder:1987:AET

[LPV87] Ernst Linder, G. P. Patil, and Douglas S. Vaughan. Application of event tree risk analysis to fisheries management. Ecological Modelling, 36(1–2):15–28, April 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008790055X.

Legovic:1982:PHS

[LR82] T. Legović and S. Rinaldi. Periodic harvesting strategies for predator–prey systems: a special case. Ecological Modelling, 14(3–4):179–191, January 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900175.

Liepmann:1985:DGS

[LS85] Dorian Liepmann and Gregory Stephanopoulos. Development and global sensitivity analysis of a closed ecosystem model. Ecological Modelling, 30(1–2):13–47, October 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900353.

Logofet:1986:AAE

[LS86a] D. O. Logofet and Y. M. Svirezhev. Averaging and aggregation in ecological models: an attempt at a non-linear approach. Ecological Modelling, 34(3–4):217–229, December 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900050.

Lorda:1986:STA

[LS86b] Ernesto Lorda and Saul B. Saila. A statistical technique for analysis of environmental data containing periodic variance components. Ecological Modelling, 32(1–3):59–69, June 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900165.

Lhotka:1987:CME

[LS87] Ladislav Lhotka and Milan Straškraba. Combinatorial model of ecosystem dynamics. Ecological Modelling, 39(1–2):181–200, November 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900202.

Lam:1989:WAM

[LSSF89] D. C. L. Lam, D. A. Swayne, J. Storey, and A. S. Fraser. Watershed acidification models using the knowledge-based systems approach. Ecological Modelling, 47(1–2):131–152, September 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089901130.

Luckyanov:1983:AVS

[LSV83] N. K. Luckyanov, Yu. M. Svirezhev, and O. V. Voronkova. Aggregation of variables in simulation models of water ecosystems. Ecological Modelling, 18(3–4):235–240, April 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900157.

Lister:1982:PSG

[LU82a] Andrew Lister and Michael B. Usher. Population systems: a general introduction: by Alan A. Berryman. Plenum Press, New York and London, 1981, xv + 222 pp., $16.95 or £10.71, ISBN 0-306-40589-X. Ecological Modelling, 16(1):77–78, July 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900746.

Logofet:1982:SSM

[LU82b] D. O. Logofet and N. B. Ulianov. Sign stability in model ecosystems: a complete class of sign-stable patterns. Ecological Modelling, 16(2–4):173–189, August 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900072.

Luckyanov:1984:LAS

[Luc84] N. K. Luckyanov. Linear aggregation and separability of models in ecology. Ecological Modelling, 21(1–2):1–12, January 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900218.

Luckyanov:1988:IAT

[Luc88] N. K. Luckyanov. Iterative approximation of trophic functions. Ecological Modelling, 40(1):1–10, January 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900981.

Leggett:1981:RIM

[LW81] Richard W. Leggett and Lynn R. Williams. A reliability index for models. Ecological Modelling, 13(4):303–312, September 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008190034X.

Ludwig:1982:OHI

[LW82] D. Ludwig and C. J. Walters. Optimal harvesting with imprecise parameter estimates. Ecological Modelling, 14(3–4): 273–292, January 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900230.

Larsen:1988:FWC

[LW88] Lawrence C. Larsen and William A. Williams. Fitting De wit competition models with general nonlinear regression programs. Ecological Modelling, 41(1–2):147–150, April 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900518.

Murphy:1982:UHF

[MA82] Charles E. Murphy and Jorge Ares. The uptake of hydrogen fluoride by a forest. Ecological Modelling, 15(3):265–285, April 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900308.

Maas:1988:URS

[Maa88] Stephan J. Maas. Use of remotely-sensed information in agricultural crop growth models. Ecological Modelling, 41 (3–4):247–268, June 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380088900312.

Mahamah:1988:SSA

[Mah88] Dintie S. Mahamah. Simplified sensitivity analysis applied to a nutrient-biomass model. Ecological Modelling, 42(2):103–109, August 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008890110X.

Malanson:1984:LLM

[Mal84] George P. Malanson. Linked Leslie matrices for the simulation of succession. Ecological Modelling, 21(1–2):13–20, January 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008490022X.

Malanson:1985:SCB

[Mal85] George P. Malanson. Simulation of competition between alternative shrub life history strategies through recurrent fires. Ecological Modelling, 27(3–4):271–283, April 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900079.

Mangel:1982:AFD

[Man82] Marc Mangel. Aggregation and fishery dynamics: Multiple time scales, times to extinction, and random environments. Ecological Modelling, 15(3):191–209, April 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900266.

Martens:1987:CLV

[Mar87] Bernd Martens. Connectance in linear and Volterra systems. Ecological Modelling, 35(3–4):157–163, March 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087901098.

Mauersberger:1982:LGL

[Mau82a] Peter Mauersberger. Logistic growth laws for phyto- and zooplankton. Ecological Modelling, 17(2):57–63, October 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900424.

Mauersberger:1982:RPP

[Mau82b] Peter Mauersberger. Rates of primary production, respiration and grazing in accordance with the balances of energy and entropy. Ecological Modelling, 17(1):1–10, September 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900837.

Mauersberger:1987:LGA

[Mau87] P. Mauersberger. Lake Gårdsjön — an acid forest lake and its catchment: F. Andersson and B. Olsson (Editors). Ecological Bulletins, 37. Publ. House Swedish Research Councils, Stockholm, 1985. 336 pp., US$44.00. ISBN 91-86344-25-0. Ecological Modelling, 38(3–4):312–313, October 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087901050.

Mauersberger:1988:TDM

[Mau88] P. Mauersberger. Three-dimensional models of marine and estuarine dynamics: Jaques C. Nihoul and Bruno M. Jamart (Editors). Proc. 18th Int. Liège Colloq. Ocean Hydrodynamics. Elsevier Oceanography Series, 45. Elsevier, Amsterdam, 1987. xii + 630 pp., Dfl. 270.00. ISBN 0-444-42794-5. Ecological Modelling, 40(1):77–78, January 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088901032.

Messer:1984:LEK

[MB84] Jay J. Messer and Patrick L. Brezonik. Laboratory evaluation of kinetic parameters for lake sediment denitrification models. Ecological Modelling, 21(4):277–286, March 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900644.

McKay:1988:EMC

[MB88] R. J. McKay and J. S. Bradley. Erlangian models for capturerecapture data. Ecological Modelling, 40(3–4):161–173, March 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900166.

Matthies:1989:EEE

[MBM^{+}89] M. Matthies, R. Brüggemann, B. Münzer, G. Schernewski, and S. Trapp. Exposure and ecotoxicity estimation for environmental chemicals (E4CHEM): Application of fate models for surface water and soil. Ecological Modelling, 47(1–2):115–130, September 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089901129.

Morioka:1986:BWE

[MC86] Tohru Morioka and Satoshi Chikami. Basin-wide ecological fate model for management of chemicals hazard. Ecological Modelling, 31(1–4):267–281, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900682.

McClanahan:1986:SDV

[McC86] T. R. McClanahan. Seed dispersal from vegetation islands. Ecological Modelling, 32(4):301–309, July 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900943.

Messia:1984:CSS

[MDS84] Maria Grazia Messia, Piero De Mottoni, and Elisabetta Santi. On a competitive system subject to switching predation. Ecological Modelling, 24(1–2):9–24, August 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900528.

Mehandjiev:1984:TAP

[Meh84] Marin R. Mehandjiev. Thermodynamics of accumulation processes applied to ecological modelling. Ecological Modelling, 26(1–2):17–20, December 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900899.

Mejer:1983:DPE

[Mej83] Henning Mejer. Diffusion processes in environmental systems: J. Crank, N. R. McFarlane, J. C. Newby, G. D. Paterson and J. B. Pedley. Macmillan Press, 1981, xi + 160 pp., £20.00, ISBN: 0-333-30721-6. Ecological Modelling, 19(3): 225, July 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900595.

Muraoka:1986:BMP

[MF86] Kohji Muraoka and Takehiko Fukushima. On the box model for prediction of water quality in eutrophic lakes. Ecological Modelling, 31(1–4):221–236, May 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900657.

Montague:1982:CHE

[MFG82] Clay L. Montague, Willard R. Fey, and David M. Gillespie. A causal hypothesis explaining predator–prey dynamics in Great Salt Lake, Utah. Ecological Modelling, 17(3–4):243–270, December 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900345.

Matsuoka:1986:EML

[MGN86] Yuzuru Matsuoka, Takeshi Goda, and Masaaki Naito. An eutrophication model of Lake Kasumigaura. Ecological Modelling, 31(1–4):201–219, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900645.

Makela:1986:SGM

[MH86] Annikki Mäkelä and Pertti Hari. Stand growth model based on carbon uptake and allocation in individual trees. Ecological Modelling, 33(2–4):205–229, October 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900414.

Morris:1984:TLB

[MHB84] James T. Morris, R. A. Houghton, and Daniel B. Botkin. Theoretical limits of belowground production by Spartina alterniflora: an analysis through modelling. Ecological Modelling, 26(3–4):155–175, December 15, 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900681.

Michaelides:1985:SMF

[Mic85] S. C. Michaelides. A simulation model of the fungus Phytophthora infestans (Mont) de Bary. Ecological Modelling, 28(1–2):121–137, May 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380085900171.

Midgley:1984:AEM

[Mid84a] D. C. Midgley. Application of ecological modelling in environmental management, part A: S. E. Jørgensen (Editor). Developments in Environmental Modelling, 4A. Elsevier Science Publishers, Amsterdam/Oxford/New York, 1983. 735 pp., US $125.00/Dfl. 295.00, ISBN 0-444-42155-6. Ecological Modelling, 23(3):269–272, June 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008490108X.

Midgley:1984:HMS

[Mid84b] D. C. Midgley. Hydrologic modeling of small watersheds: C. T. Haan, H. P. Johnson and D. L. Brakensiek (Editors). American Society of Agricultural Engineers, American Technical Publishers Ltd., Hitchin, Herts., Great Britain, 1982. 533 pp., £25.50. ISBN 0-916150-44-5. Ecological Modelling, 23(4):342– 344, July 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084901297.

Matsuda:1981:MAE

[MIK81] Ikuo Matsuda, Masumi Ishikawa, and Yoichi Kaya. Macroscopic approach to the environment in an industry allocation model. Ecological Modelling, 12(1–2):69–82, March 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900259.

Mikhailovsky:1986:LCA

[Mik86] George E. Mikhailovsky. The law of congruous attraction and the structure of zooplankton communities. Ecological Modelling, 34(1–2):83–98, November 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900803.

Mitsch:1987:FEM

[Mit87] William J. Mitsch. Freshwater Ecosystems: Modelling and simulation: Milan Straškraba and Albrecht H. Gnauck. Developments in Environmental Modelling, Vol. 8. Elsevier, Amsterdam/Oxford/New York/Tokyo, 1985. 309 pp., Dfl. 210.00. ISBN 0-444-99567-6. Ecological Modelling, 35(3–4):310–311, March 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087901189.

Miyanaga:1986:MSF

[Miy86] Yoichi Miyanaga. Modelling of stratified flow and eutrophication in reservoirs. Ecological Modelling, 31(1–4):133–144, May 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900608.

Magara:1986:CAA

[MK86] Yasumoto Magara and Shoichi Kunikane. Cost analysis of the adverse effects of algal growth in water bodies on drinking water supply. Ecological Modelling, 31(1–4):303–313, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900712.

Marsili-Libelli:1984:AMA

[ML84] Stefano Marsili-Libelli. Aggregated model of activated sludge processes. Ecological Modelling, 24(3–4):171–190, September 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900395.

Mohn:1987:RBM

[MM87] R. K. Mohn and R. J. Miller. A ration-based model of a seaweed-sea urchin community. Ecological Modelling, 37(3– 4):249–267, July 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380087900287.

Moore:1989:MGE

[Moo89]  Andrew D. Moore.  On the maximum growth equa-

tion used in forest gap simulation models. Ecological Modelling, 45(1):63–67, February 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089901002.

Mangel:1983:MMA

[MP83] Marc Mangel and Richard E. Plant. Multiseasonal management of an agricultural pest. I: Development of the theory. Ecological Modelling, 20(1):1–19, October 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900285.

McLellan:1986:HCS

[MR86] A. R. McLellan and C. M. Rowland. A honeybee colony swarming model. Ecological Modelling, 33(2–4):137–148, October 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900372.

Muetzelfeldt:1989:UPI

[MRBU89] R. Muetzelfeldt, D. Robertson, A. Bundy, and M. Uschold. The use of Prolog for improving the rigour and accessibility of ecological modelling. Ecological Modelling, 46(1–2):9– 34, July 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900677.

Mulholland:1987:AAA

[MRE87] Robert J. Mulholland, John S. Read, and William R. Emanuel. Asymptotic analysis of airborne fraction used to validate global carbon models. Ecological Modelling, 36(1–2):139–152, April 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900639.

Majkowski:1981:MSA

[MRM81] Jacek Majkowski, Joanne M. Ridgeway, and Donald R. Miller. Multiplicative sensitivity analysis and its role in development of simulation models. Ecological Modelling, 12(3):191–208, April 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900570.

Mauersberger:1987:TAG

[MS87a] Peter Mauersberger and Milan Straškraba. Two approaches to generalized ecosystem modelling: Thermodynamic and cybernetic. Ecological Modelling, 39(1–2):161–169, November 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900184.

Mendoza:1987:YPM

[MS87b] Guillermo A. Mendoza and Janes Siahaya. Yield prediction models for Pinus merkusii plantations in Indonesia. Ecological Modelling, 36(3–4):181–194, May 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900676.

MacNeil:1985:SAG

[MSH85] M. D. MacNeil, J. W. Skiles, and J. D. Hanson. Sensitivity analysis of a general rangeland model. Ecological Modelling, 29(1–4):57–76, September 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008590047X.

Mack:1987:MMP

[MSR87] T. P. Mack, J. W. Smith, and R. B. Reed. A mathematical model of the population dynamics of the lesser cornstalk borer, Elasmopalpus lignosellus. Ecological Modelling, 39(3–4):269–286, December 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900044.

Morrison:1985:SFR

[MTC85] K. A. Morrison, N. Thérien, and B. Coupal. Simulating fish redistribution in the LG-2 reservoir after flooding. Ecological Modelling, 28(1–2):97–111, May 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900158.

Murtaugh:1988:ULR

[Mur88] Paul A. Murtaugh. Use of logistic regression in modelling prey selection by Neomysis mercedis. Ecological Modelling, 43(3–4):225–233, November 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900051.

Maclean:1980:SWE

[MW80] David A. Maclean and Ross W. Wein. Simulation of wildfire effects on the nitrogen cycle of a Pinus banksiana ecosystem in New Brunswick, Canada. Ecological Modelling, 10(3–4):167– 192, August 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900587.

North:1988:CSS

[NBF88] Philip M. North, Alex W. Boddy, and Duncan R. Forrester. A computer simulation study of stochastic models to investigate the population dynamics of the screech owl (Otus asio) under increased mortality. Ecological Modelling, 40(3–4):233– 263, March 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900208.

Nemchenko:1987:CMM

[Nem87] Olga Nemchenko. A compartmental model of metabolite utilization for plant growth. Ecological Modelling, 39(1–2):17–32, November 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008790010X.

Nishida:1986:IAP

[NHTO86] N. Nishida, S. Hiratsuka, T. Tanaka, and T. Okino. An interactive algorithm for the parameter estimation of complex systems and its application to an ecological modelling of an actual Japanese lake. Ecological Modelling, 32(1–3):85–93, June 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900189.

Nihoul:1981:DEP

[Nih81] Jacques C. J. Nihoul. Diffusion and Ecological Problems: Mathematical Models: Akira Okubo. Springer-Verlag, 1980, 254 pp., 114 figs., DM 78.00/U.S. $46.10 (cloth), ISBN 3-540-09620-5. Ecological Modelling, 13(4):313, September 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900351.

Nihoul:1982:PES

[Nih82] J. C. J. Nihoul. Principles of environmental science and technology: edited by S. E. Jørgensen and I. Johnsen, Studies in Environmental Science 14, Elsevier, Amsterdam, 1981, 516 pp., US$53.25/Dfl.125.00, ISBN 0-444-99721-0. Ecological Modelling, 16(1):79–80, July 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900758.

Nihoul:1983:ASO

[Nih83a] Jacques C. J. Nihoul. Analytical solutions of the onedimensional connective-dispersive solute transport equation: M. Th. van Genuchten and W. J. Alves. U.S. Department of Agriculture, Technical Bulletin No. 1661, 1982, 151 pp. Ecological Modelling, 19(3):222–223, July 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900571.

Nihoul:1983:MJR

[Nih83b] Jacques C. J. Nihoul. Modelling: John N. R. Jeffers, Outline Studies in Ecology. Chapman and Hall, London, New York, 1982, 80 pp., 9 fig., 8 tables, Paperback £2.75, ISBN: 0-412-24360-1. Ecological Modelling, 19(3):234, July 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900649.

Nihoul:1984:ESA

[Nih84a] Jacques C. J. Nihoul. Environmental systems analysis and management: S. Rinaldi (Editor). Proceedings of the IFIP WG 7.1 Working Conference on Environmental Systems Analysis and Management, Rome, Italy, 28–30 September, 1981. North-Holland Publ. Co., Amsterdam and New York, 1982. 750 pp., US $93.00/Dfl. 200.00, ISBN 0-444-86406-7. Ecological Modelling, 23(3):272–273, June 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084901091.

Nihoul:1984:SEI

[Nih84b] Jacques C. J. Nihoul. Systems ecology, An Introduction: Howard T. Odum. John Wiley, West Sussex, Great Britain, 1983. 644 pp., 702 figures, £47.50. ISBN 0-471-65277-6. Ecological Modelling, 23(4):346–348, July 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084901315.

Nihoul:1984:NLM

[Nih84c] Jacques C. J. Nihoul. A non-linear mathematical model for the transport and spreading of oil slicks. Ecological Modelling, 22(1–4):325–339, April 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900188.

Nihoul:1984:AEM

[Nih84d] Jaques C. J. Nihoul. Application of ecological modelling in environmental management, part B: S. E. Jørgensen and W. J. Mitsch (Editors). Developments in Environmental Modelling, 4B. Elsevier Science Publishers, Amsterdam/ Oxford/New York/Tokyo, 1983. vii + 438 pp., Dfl. 175.00/ US$67.50 (U.S.A. and Canada). ISBN 0-444-42247-1. Ecological Modelling, 26(3–4):320–321, December 15, 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900796.

Nihoul:1985:SLE

[Nih85] Jacques C. J. Nihoul. The siting of liquefied energy gas facilities in four countries: Howard C. Kunreuther and Joanne Linnerooth et al. Springer-Verlag, Berlin/Heidelberg/ New York, 1983. xviii + 290 pp., Cloth DM 49.00/ approximately US$19.10. ISBN 3-540-12804-2. Ecological Modelling, 28(3):238–239, June 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900900.

Nyholm:1984:MHM

[NNP84] Niels Nyholm, Tue K. Nielsen, and Knud Pedersen. Modeling heavy metals transport in an Arctic fjord system polluted from mine tailings. Ecological Modelling, 22(1–4):285– 324, April 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900176.

Nokoe:1984:RSA

[NO84] Sagary Nokoe and Julius A. Okojie. Relationship of stand attributes of some plantation mahoganies with estimated Weibull parameters. Ecological Modelling, 24(3–4):231–240, September 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900437.

North:1985:CMS

[Nor85] Philip M. North. A computer modelling study of the population dynamics of the screech owl (Otus asio). Ecological Modelling, 30(1–2):105–143, October 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900390.

Norgaard:1987:EMD

[Nor87] Richard B. Norgaard. Economics as mechanics and the demise of biological diversity. Ecological Modelling, 38(1–2):107–121, September 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900470.

Nakanishi:1986:SMB

[NUK86] Hiroshi Nakanishi, Masao Ukita, and Yasuharu Kawai. Study on the modelling of the behavior of phosphorus released from sediments. Ecological Modelling, 31(1–4):105–123, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008690058X.

Nuttall:1989:SPD

[Nut89] Roger M. Nuttall. Simulated population dynamics of a stored-products’ pest (Ptinus tectus, coleoptera). Ecological Modelling, 48(3–4):291–313, November 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900537.

Outcalt:1982:MAT

[OA82] Samuel I. Outcalt and Harold L. Allen. Modeling the annual thermal regime of lake ohrid, Yugoslavia, using daily weather data. Ecological Modelling, 15(2):165–184, March 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900606.

Oker-Blom:1989:RBR

[OBPK89] Pauline Oker-Blom, Timo Pukkala, and Timo Kuuluvainen. Relationship between radiation interception and photosynthesis in forest canopies: effect of stand structure and latitude. Ecological Modelling, 49(1–2):73–87, December 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900446.

ONeill:1989:MNL

[ODP^{+}89] R. V. O’Neill, D. L. DeAngelis, J. J. Pastor, B. J. Jackson, and W. M. Post. Multiple nutrient limitations in ecological models. Ecological Modelling, 46(3–4):147–163, August 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008990015X.

Odum:1983:MPE

[Odu83] Howard T. Odum. Maximum power and efficiency: a rebutttal. Ecological Modelling, 20(1):71–82, October 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900327.

Olufeagba:1981:MCD

[OF81] B. J. Olufeagba and R. H. Flake. Modelling and control of dissolved oxygen in an estuary. Ecological Modelling, 14 (1–2):79–94, November 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380081900156.

Ohuchi:1986:MLT

[OFMS86] A. Ohuchi, J. Fukuoka, A. Miyakoshi, and M. Suzuki. Modelling of the lower trophic levels of a marine ecosystem and its example of short-period variations of chlorophyll and nutrient in harima-nada. Ecological Modelling, 32(1–3):149–163, June 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900232.

ONeill:1982:PCS

[OGC82] R. V. O’Neill, R. H. Gardner, and J. H. Carney. Parameter constraints in a stream ecosystem model: Incorporation of a priori information in Monte Carlo error analysis. Ecological Modelling, 16(1):51–65, July 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900722.

Oliver:1988:OMB

[OL88a] J. Douglas Oliver and Tarzan Legović. Okefenokee marshland before, during and after nutrient enrichment by a bird rookery. Ecological Modelling, 43(3–4):195–223, November 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008890004X.

Ouimet:1988:PAM

[OL88b] Chantal Ouimet and Pierre Legendre. Practical aspects of modelling ecological phenomena using the cusp catastrophe. Ecological Modelling, 42(3–4):265–287, September 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900610.

Oi:1986:ACS

[OMA^{+}86] Ko Oi, Sadaaki Miyamoto, Osamu Abe, Atsuo Katsuya, and Kazuhiko Nakayama. Analysis of cognitive structures of environment of local residents through word association methods. Ecological Modelling, 32(1–3):29–41, June 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900141.

Okojie:1985:SDM

[ON85] J. A. Okojie and S. Nokoe. Stand development of mixtures of species with varying tolerances to shade. Ecological Modelling, 30(1–2):71–82, October 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900377.

Onstad:1988:PDT

[Ons88] David W. Onstad. Population-dynamics theory: the roles of analytical, simulation, and supercomputer models. Ecological Modelling, 43(1–2):111–124, October 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900750.

Ogawa:1986:MIE

[OSJ^{+}86] Hisashi Ogawa, Kiichiro Sato, Nobuo Jo, Kageyu Noro, and Kenzaburo Tsuchiya. Modelling of industrial ecological systems for evaluation of health services. Ecological Modelling, 31(1–4):329–339, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900736.

Ostojski:1987:MMS

[Ost87] Mieczysław Stefan Ostojski. Mathematical model of sewage treatment plant operation. Ecological Modelling, 39(1–2): 67–83, November 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380087900147.

Olson:1985:WPM

[OSW85] Richard L. Olson, Peter J. H. Sharpe, and Hsin-I Wu. Whole-plant modelling: a continuous-time Markov (CTM) approach. Ecological Modelling, 29(1–4):171–187, September 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900523.

Packard:1988:WCE

[Pac88] Jane M. Packard. Wildlife conservation evaluation: M. B. Usher (Editor). Chapman and Hall, London, 1986. 394 pp., £30.00 (hardback). ISBN 0-412-26750-0. Ecological Modelling, 41(3–4):327–328, June 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900361.

Paloheimo:1988:EMP

[Pal88] J. E. Paloheimo. Estimation of marine production from size spectrum. Ecological Modelling, 42(1):33–44, July 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900907.

Paling:1989:RMH

[Pal89] W. A. J. Paling. Risk management and hazardous waste — Implementation and the dialectics of credibility: Brian Wynne. Springer, Berlin, Federal Republic of Germany, 1987. Hardcover, 447 pp., DM98.00. ISBN 3-540-18243-8. Ecological Modelling, 47(3–4):313–314, September 15, 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900094.

Parker:1982:DDP

[Par82] Richard A. Parker. Differential dispersion in planktonic food chains with constant inputs. Ecological Modelling, 15(1):75–78, February 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900680.

Parker:1986:SDC

[Par86] R. A. Parker. Simulating the development of chlorophyll maxima in the Celtic Sea. Ecological Modelling, 33(1):1–11, September 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086901043.

Patten:1983:LEE

[Pat83] Bernard C. Patten. Linearity enigmas in ecology. Ecological Modelling, 18(3–4):155–170, April 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900108.

Patten:1984:STF

[Pat84] Bernard C. Patten. System theory formulation of sitespecific water quality standards and protocols. Ecological Modelling, 23(4):313–340, July 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084901273.

Patten:1985:ECE

[Pat85] Bernard C. Patten. Energy cycling in the ecosystem. Ecological Modelling, 28(1–2):1–71, May 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900134.

Paton:1986:MMA

[Pat86] G. Paton. A matrix modelling approach to population growth systems involving multiple time delays. Ecological Modelling, 34(3–4):197–216, December 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900049.

Patten:1988:NTD

[Pat88] Bernard C. Patten. Network trophic dynamics: Reply to R. A. Herendeen. Ecological Modelling, 42(1):78–84, July 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900944.

Pickup:1986:RFM

[PC86] G. Pickup and V. H. Chewings. Random field modelling of spatial variations in erosion and deposition in flat alluvial landscapes in arid central Australia. Ecological Modelling, 33(2–4):269–296, October 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008690044X.

Portier:1981:CBR

[PE81] Kenneth M. Portier and Katherine Carter Ewel. Confidence bands for a regression-based, non-linear decomposition model. Ecological Modelling, 14(1–2):125–132, November 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008190017X.

Pearce:1987:FEE

[Pea87] David Pearce. Foundations of an ecological economics. Ecological Modelling, 38(1–2):9–18, September 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900421.

Plichta:1989:GAM

[PG89] W. Plichta and M. Gurtowski. A general analytical model of the process of humus mineralization and accumulation in soil. Ecological Modelling, 44(3–4):209–217, January 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900306.

Patten:1984:MCI

[PH84] Bernard C. Patten and Masahiko Higashi. Modified cycling index for ecological applications. Ecological Modelling, 25(1–3):69–83, October 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084900334.

Phillips:1985:SAD

[Phi85] Jonathan D. Phillips. Stability of artificially-drained lowlands: a theoretical assessment. Ecological Modelling, 27(1–2):69–79, March 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900250.

Phillips:1989:ESF

[Phi89] Jonathan D. Phillips. An evaluation of the state factor model of soil ecosystems. Ecological Modelling, 45(3):165–177, May 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008990080X.

Padilla:1988:MTF

[PLRV88] Francisco Padilla, Pierre Lafrance, Claude Robert, and Jean-Pierre Villeneuve. Modeling the transport and the fate of pesticides in the unsaturated zone considering temperature effects. Ecological Modelling, 44(1–2):73–88, December 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008890083X.

Patten:1982:WEO

[PM82] Bernard C. Patten and James H. Matis. The water environs of okefenokee swamp: an application of static linear environ analysis. Ecological Modelling, 16(1):1–50, July 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900710.

Paterson:1989:MIE

[PM89] Sally Paterson and Donald Mackay. A model illustrating the environmental fate, exposure and human uptake of persistent organic chemicals. Ecological Modelling, 47(1–2):85–114, September 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089901117.

Pearlstine:1985:MIR

[PMK85] Leonard Pearlstine, Henry McKellar, and Wiley Kitchens. Modelling the impacts of a river diversion on bottomland forest communities in the Santee River floodplain, South Carolina. Ecological Modelling, 29(1–4):283–302, September 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900572.

Podvintsev:1982:ISM

[Pod82] Yury Vitalievich Podvintsev. Information-statistical model of the ecological system of the Black Sea environment. Ecological Modelling, 17(2):175–180, October 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900539.

Pola:1985:NSF

[Pol85] Nancy B. Pola. Numerical simulation of fish migrations in the eastern Bering Sea. Ecological Modelling, 29(1–4):327–351, September 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900602.

Power:1984:NMS

[Pow84] James H. Power. A numerical method for simulating plankton transport. Ecological Modelling, 23(1–2):53–66, May 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084901182.

Prasad:1983:ISM

[PPH83] Shivaji Prasad, B. D. Patil, and C. R. Hazra. An introduction to systems and models. Ecological Modelling, 18(2):73–83, February 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900479.

Parton:1984:ABS

[PS84] W. J. Parton and J. S. Singh. Adapting a biomass simulation model to a tropical grassland. Ecological Modelling, 23(1–2):151–163, May 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084901236.

Polgar:1988:EPP

[PTS88] T. T. Polgar, M. A. Turner, and J. K. Summers. Effect of power plant entrainment on the population dynamics of the bay anchovy (Anchoa mitchilli). Ecological Modelling, 41(3–4):201– 218, June 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900282.

Park:1988:NMZ

[PU88] Seok S. Park and Christopher G. Uchrin. A numerical mixing zone model for water quality assessment in natural streams: Conceptual development. Ecological Modelling, 42(3–4):233–244, September 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900592.

Qiwu:1985:MMP

[QF85] Cui Qiwu and Lu Fengyong. A mathematical model of predation based upon the theory of nutrition kinetics. Ecological Modelling, 28(1–2):155–164, May 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900195.

Qiwu:1988:MMP

[QF88] Cui Qiwu and Lu Fengyong. A mathematical model of predation based upon the theory of nutrition kinetics 2. A nutrition structure of the predator population and its functional response to the prey. Ecological Modelling, 40(1):67–76, January 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088901020.

Reynolds:1985:PRP

[RA85] James F. Reynolds and Basil Acock. Predicting the response of plants to increasing carbon dioxide: a critique of plant growth models. Ecological Modelling, 29(1–4):107–129, September 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900493.

Racsko:1987:SIA

[Rac87] Peter Racsko. Stability and instability of an adaptive resource replacement strategy in an agroecological production system. Ecological Modelling, 35(3–4):165–174, March 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087901104.

Rago:1984:PFA

[Rag84] Paul J. Rago. Production forgone: an alternative method for assessing the consequences of fish entrainment and impingement losses at power plants and other water intakes. Ecological Modelling, 24(1–2):79–111, August 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900565.

Rastetter:1987:ACI

[Ras87] Edward B. Rastetter. Analysis of community interactions using linear transfer function models. Ecological Modelling, 36(1–2):101–117, April 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900615.

Roise:1981:AFK

[RBK81] Joseph P. Roise, David R. Betters, and Brian M. Kent. An approach to functionalizing key environmental factors: Forage production in rocky mountain aspen stands. Ecological Modelling, 14(1–2):133–146, November 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900181.

Rinaldo:1989:SSN

[RBM89] Andrea Rinaldo, Alberto Bellin, and Alessandro Marani. A study on solute NO_{3}–N transport in the hydrologic response by an MRF model. Ecological Modelling, 48(3–4):159–191, November 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900483.

Reckhow:1983:CWQ

[RC83] Kenneth H. Reckhow and Steven C. Chapra. Confirmation of water quality models. Ecological Modelling, 20(2–3):113–133, November 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900029.

Running:1988:GMF

[RC88] Steven W. Running and Joseph C. Coughlan. A general model of forest ecosystem processes for regional applications. I. Hydrologic balance, canopy gas exchange and primary production processes. Ecological Modelling, 42(2):125–154, August 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088901123.

Ray:1989:LVP

[RC89] S. Saha Ray and K. S. Chaudhuri. Lotka–Volterra preypredator model with harvesting and environmental perturbations. Ecological Modelling, 47(3–4):283–290, September 15, 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900069.

Recknagel:1984:CSA

[Rec84] Frieder Recknagel. A comprehensive sensitivity analysis for an ecological simulation model. Ecological Modelling, 26 (1–2):77–96, December 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380084900954.

Recknagel:1985:ASS

[Rec85] Frieder Recknagel. Analysis of structural stability of aquatic ecosystems as an aid for ecosystem control. Ecological Modelling, 27(3–4):221–234, April 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900043.

Reed:1982:SSH

[Ree82] William J. Reed. Sex-selective harvesting of Pacific salmon: a theoretically optimal solution. Ecological Modelling, 14 (3–4):261–271, January 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380082900229.

Reed:1983:MCM

[Ree83] Mark Reed. A multidimensional continuum model of fish behavior. Ecological Modelling, 20(4):311–322, December 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008390039X.

Reuss:1980:SSN

[Reu80] John O. Reuss. Simulation of soil nutrient losses resulting from rainfall acidity. Ecological Modelling, 11(1):15–38, October 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900691.

Reed:1989:SME

[RFCC89] Mark Reed, Deborah P. French, John Calambokidis, and James C. Cubbage. Simulation modelling of the effects of oil spills on population dynamics of Northern fur seals. Ecological Modelling, 49(1–2):49–71, December 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900434.

Reed:1987:SME

[RFJ87] Mark Reed, Deborah French, and Katherine Jayko. Simulation of marine ecosystem effects due to PCB waste incineration in the Gulf of Mexico. Ecological Modelling, 38(3–4):213–242, October 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900986.

Rykiel:1985:F

[RG85] Edward J. Rykiel and William E. Grant. Foreword. Ecological Modelling, 29(1–4):3, September 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900432.

Rykiel:1988:F

[RG88] Edward J. Rykiel and William E. Grant. Foreword. Ecological Modelling, 43(1–2):3, October 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900695.

Rabinovich:1985:SMM

[RHC85] Jorge E. Rabinovich, Maria J. Hernández, and Jorge L. Cajal. A simulation model for the management of vicuña populations. Ecological Modelling, 30(3–4):275–295, December 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900717.

Rexstad:1985:MST

[RI85] Eric Rexstad and George S. Innis. Model simplification — three applications. Ecological Modelling, 27(1–2):1–13, March 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900213.

Ritchie:1989:ESR

[Rit89] Jerry R. Ritchie. An expert system for a rangeland simulation model. Ecological Modelling, 46(1–2):91–105, July 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900719.

Rao:1982:DVS

[RJ82] P. S. C. Rao and R. E. Jessup. Development and verification of simulation models for describing pesticide dynamics in soils. Ecological Modelling, 16(1):67–75, July 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900734.

Rodriguez:1988:RRG

[RJ88] Abelardo Rodriguez and Donald A. Jameson. Rainfall risk in grazing management. Ecological Modelling, 41(1–2):85–100, April 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900476.

Reed:1988:NMB

[RJBL88] Mark Reed, Katherine Jayko, Ann Bowles, and Steve Leatherwood. Numerical models of bowhead and gray whale migration in Alaskan waters. Ecological Modelling, 44(1–2):1–42, December 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900804.

Rowland:1987:SCD

[RM87] C. M. Rowland and A. R. McLellan. Seasonal changes of drone numbers in a colony of the honeybee, Apis mellifera. Ecological Modelling, 37(3–4):155–166, July 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900238.

Rose:1989:DMC

[RMS89] Kenneth A. Rose, Richard I. McLean, and J. Kevin Summers. Development and Monte Carlo analysis of an oyster bioaccumulation model applied to biomonitoring data. Ecological Modelling, 45(2):111–132, April 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900872.

Roberts:1989:AFS

[Rob89] David W. Roberts. Analysis of forest succession with fuzzy graph theory. Ecological Modelling, 45(4):261–274, June 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900756.

Rossis:1986:RRM

[Ros86] George Rossis. Reductionism and related methodological problems in ecological modelling. Ecological Modelling, 34(3–4):289–298, December 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900098.

Rossi:1986:CLE

[RPM86] Giovanni Rossi, Guido Premazzi, and Giuseppe Marengo. Correlation of a lake eutrophication model to field experiments. Ecological Modelling, 34(3–4):167–189, December 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900025.

Reed:1984:LHA

[RRW84] K. L. Reed, K. A. Rose, and R. C. Whitmore. Latin hypercube analysis of parameter sensitivity in a large model of outdoor recreation demand. Ecological Modelling, 24(3–4):159–169, September 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900383.

Riley:1988:MDL

[RS88] Michael J. Riley and Heinz G. Stefan. Minlake: a dynamic lake water quality simulation model. Ecological Modelling, 43(3–4):155–182, November 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900026.

Racsko:1989:AMP

[RS89] P. Racsko and M. Semenov. Analysis of mathematical principles in crop growth simulation models. Ecological Modelling, 47(3–4):291–302, September 15, 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900070.

nek:1987:PDF

[RSG87] Marcel Rejmánek, James D. Smith, and Richard A. Goyer. Population dynamics of the forest tent caterpillar (Malacosoma disstria) in a water tupelo (Nyssa aquatica) forest: a simulation model. Ecological Modelling, 39(3–4):287–305, December 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900056.

Rose:1988:SIC

[RSKT88] Kenneth A. Rose, Gordon L. Swartzman, Andrew C. Kindig, and Frieda B. Taub. Stepwise iterative calibration of a multi-species phytoplankton-zooplankton simulation model using laboratory data. Ecological Modelling, 42(1):1–32, July 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900890.

Rauscher:1983:FMG

[RSS83] H. Michael Rauscher, David Wm. Smith, and Terry L. Sharik. A forest management gaming model of the nitrogen cycle in Appalachian upland oak forests. Ecological Modelling, 20(2–3):175–199, November 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900066.

Reeves:1989:ADM

[RU89] Stuart A. Reeves and Michael B. Usher. Application of a diffusion model to the spread of an invasive species: the coypu in Great Britain. Ecological Modelling, 47(3–4):217–232, September 15, 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900021.

Rubtsov:1983:MMD

[Rub83] V. V. Rubtsov. Mathematical model for development of leaf-eating insects (oakleaf roller taken as an example). Ecological Modelling, 18(3–4):269–289, April 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900182.

Rogowski:1987:MEM

[RW87] A. S. Rogowski and B. E. Weinrich. Modeling the effects of mining and erosion on biomass production. Ecological Modelling, 35(1–2):85–112, February 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900925.

Reinhardt:1989:AES

[RWJ89] Elizabeth Reinhardt, Alden H. Wright, and David H. Jackson. An advisory expert system for designing fire prescriptions. Ecological Modelling, 46(1–2):121–133, July 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900732.

Rykiel:1989:AIE

[Ryk89] Edward J. Rykiel. Artificial intelligence and expert systems in ecology and natural resource management. Ecological Modelling, 46(1–2):3–8, July 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900665.

Spiller:1986:MMS

[SA86] Gary B. Spiller and Allan N. D. Auclair. A mathematical model of seasonal and spatial variation in phosphorus concentrations in lake memphremagog, Quebec. Ecological Modelling, 34(3–4):143–166, December 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900013.

Shiyomi:1986:MEF

[SAT86] Masae Shiyomi, Tsuyoshi Akiyama, and Shigeo Takahashi. Modelling of energy flows and conversion efficiencies in a grassland ecosystem. Ecological Modelling, 32(1–3):119–135, June 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900219.

Sawaragi:1983:APM

[Saw83] Y. Sawaragi. Air pollution modeling and its application I: C. de Wispelaere (Editor), NATO — Challenges of Modern Society, Vol. I. Plenum Publishing, New York, 1981, xiv + 748 pp., illus., US $75.00, ISBN: 0-306-40820-1. Ecological Modelling, 19(3):223–224, July 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900583.

Sawaragi:1985:EPE

[SB85] Y. Sawaragi and N. Baba. Effects of pollutants at the ecosystem level: P. J. Sheehan, D. R. Miller and G. C. Butler. Scope,

22. Wiley, Chichester, Great Britain, 1984. vii + 433 pp., £31.90. ISBN 0-471-90204-7. Ecological Modelling, 28(3):233– 234, June 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900869.

Sala:1988:BEP

[SBL88] O. E. Sala, M. E. Biondini, and W. K. Lauenroth. Bias in estimates of primary production: an analytical solution. Ecological Modelling, 44(1–2):43–55, December 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900816.

Sklar:1985:DSS

[SCD85] Fred H. Sklar, Robert Costanza, and John W. Day. Dynamic spatial simulation modeling of coastal wetland habitat succession. Ecological Modelling, 29(1–4):261–281, September 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900560.

Schut:1985:MPE

[Sch85] H. E. Schut. Models for the physiological effects of short O_{3} exposures on plants. Ecological Modelling, 30(3–4):175–207, December 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900675.

Sciandra:1986:SMS

[Sci86] Antoine Sciandra. Study and modelling of a simple planktonic system reconstituted in an experimental microcosm. Ecological Modelling, 34(1–2):61–82, November 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900797.

Snodgrass:1983:TTM

[SD83] W. J. Snodgrass and P. J. Dillon. A test of two models of different levels of complexity for predicting changes of phosphorus concentration in a lake’s outflow. Ecological Modelling, 19(3):163–187, July 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900522.

Seip:1980:MMC

[Sei80] Knut L. Seip. A mathematical model of competition and colonization in a community of marine benthic algae. Ecological Modelling, 10(2):77–104, July 1980. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380080900654.

Seno:1988:ESP

[Sen88] Hiromi Seno. Effect of a singular patch on population persistence in a multi-patch system. Ecological Modelling, 43(3–4):271–286, November 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900087.

Senft:1989:HFM

[Sen89] Richard L. Senft. Hierarchical foraging models: Effects of stocking and landscape composition on simulated resource use by cattle. Ecological Modelling, 46(3–4):283–303, August 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900239.

Seymour:1989:APN

[Sey89] R. M. Seymour. Is Acanthaster planci a near-optimal predator? Ecological Modelling, 46(3–4):239–260, August 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900203.

Sharpe:1982:IEC

[SF82] David M. Sharpe and David E. Fields. Integrating the effects of climate and seed fall velocities on seed dispersal by wind: a model and application. Ecological Modelling, 17(3–4):297–310, December 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900369.

Samuel:1983:ESO

[SF83] Michael D. Samuel and Theodore C. Foin. Exploiting sea otter populations: a simulation analysis. Ecological Modelling, 20(4):297–309, December 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900388.

Suga:1986:MSR

[SFO^{+}86] Yuzo Suga, Tanehiro Futagami, Kaneo Okano, Sotaro Tanaka, and Toshiro Sugimura. A microcomputer system for remotely sensed image data on ecological environments. Ecological Modelling, 32(1–3):15–27, June 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008690013X.

Shukla:1989:DSR

[SFP^{+}89] J. B. Shukla, H. I. Freedman, V. M. Pal, O. P. Misra, M. Agarwal, and A. Shukla. Degradation and subsequent regeneration of a forestry resource: a mathematical model. Ecological Modelling, 44(3–4):219–229, January 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900318.

Shapiro:1984:MDO

[SFS84] Alexander P. Shapiro, Efim Ja. Frisman, and Elena J. Skaletskaya. Modelling dynamics and optimal exploitation of the population of the deer Cervus nippon. Ecological Modelling, 26(1–2):41–44, December 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900929.

Starfield:1989:RBE

[SFT89] A. M. Starfield, B. P. Farm, and R. H. Taylor. A rulebased ecological model for the management of an estuarine lake. Ecological Modelling, 46(1–2):107–119, July 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900720.

Shaffer:1985:SMP

[SG85] Phillip L. Shaffer and Harvey J. Gold. A simulation model of population dynamics of the codling moth, Cydia pomonella. Ecological Modelling, 30(3–4):247–274, December 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900705.

Sawyer:1985:SSD

[SH85] Alan J. Sawyer and Dean L. Haynes. Simulating the spatiotemporal dynamics of the cereal leaf beetle in a regional crop system. Ecological Modelling, 30(1–2):83–104, October 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900389.

Sawyer:1986:CLB

[SH86] Alan J. Sawyer and Dean L. Haynes. Cereal leaf beetle spatial dynamics: Simulations with a random diffusion model. Ecological Modelling, 33(2–4):89–99, October 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900347.

Shiyomi:1980:PAS

[Shi80] Masae Shiyomi. Predation-affected spatial pattern changes in a prey population. Ecological Modelling, 11(1):1–14, October 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008090068X.

Seiderer:1982:RLM

[SHL82] L. J. Seiderer, B. D. Hahn, and L. Lawrence. Rocklobsters, mussels and man: a mathematical model. Ecological Modelling, 17(3–4):225–241, December 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900333.

Sievanen:1988:MEP

[SHOP88] R. Sievänen, P. Hari, P. J. Orava, and P. Pelkonen. A model for the effect of photosynthate allocation and soil nitrogen on plant growth. Ecological Modelling, 41(1–2):55–65, April 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900440.

Schwarzenbach:1984:MCH

[SI84] R. P. Schwarzenbach and D. M. Imboden. Modelling concepts for hydrophobic organic pollutants in lakes. Ecological Modelling, 22(1–4):171–212, April 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900140.

Silvert:1982:TPE

[Sil82] William Silvert. The theory of power and efficiency in ecology. Ecological Modelling, 15(2):159–164, March 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008290059X.

Silvert:1989:MM

[Sil89] William Silvert. Modelling for managers. Ecological Modelling, 47(1–2):53–64, September 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089901099.

Sjoberg:1980:ZFQ

[Sjö80] Stig Sjöberg. Zooplankton feeding and queueing theory. Ecological Modelling, 10(3–4):215–225, August 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900605.

Summers:1980:SME

[SKMD80] J. K. Summers, W. M. Kitchens, H. N. McKellar, and R. F. Dame. A simulation model of estuarine subsystem coupling and carbon exchange with the sea. II. North Inlet model structure, output and validation. Ecological Modelling, 11(2):101–140, November 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900897.

Skogerboe:1985:MTE

[Sko85] Gaylord V. Skogerboe. Mediterranean-type ecosystems: the role of nutrients: F. J. Kruger, D. T. Mitchell and J. U. M. Jarvis (Editors). Ecological Studies, 43. Springer, Berlin/ Heidelberg, 1983. 552 pp., US$40.50. ISBN 3-540-12158-7. Ecological Modelling, 28(3):231–232, June 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900857.

Skogerboe:1986:AEU

[Sko86a] Gaylord V. Skogerboe. Agricultural ecosystems: Unifying concepts: R. Lowrance, B. R. Stinner and G. J. House (Editors). A Wiley-Interscience Publication, Wiley, Great Britain, 1984. x + 233 pp., £40.85. ISBN 0-471-87888-X. Ecological Modelling, 32(4):319–321, July 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086901006.

Skogerboe:1986:EAT

[Sko86b] Gaylord V. Skogerboe. Energy and agriculture: Their interacting futures: (Policy Implications of Global Models). M. Levy and J. L. Robindson (Editors). Harwood Academic Publishers for the United Nations University, London, 1984. xii + 371 pp., US $78.00. ISSN 0739-5671, Vol. 1/ISBN 3-7186-0187-7. Ecological Modelling, 32(4):321–322, July 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086901018.

Skogerboe:1986:MMA

[Sko86c] Gaylord V. Skogerboe. Mathematical models in agriculture: J. France and J. H. M. Thornley. Butterworth & Co. (Publishers) Ltd., London, 1984. xi + 335 pp. ISBN 0-408-10868-1. Ecological Modelling, 32(4):317–319, July 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900992.

Skogerboe:1986:PPM

[Sko86d] Gaylord V. Skogerboe. Pollutants in porous media: the unsaturated zone between soil surface and groundwater: B. Yaron, G. Dagan and J. Goldschmid (Editors). Springer, Berlin/ Heidelberg, 1984. xvi + 296 pp., DM139/US$51.90. ISBN 3-540-13179-5. Ecological Modelling, 33(1):77–79, September 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086901080.

Skogerboe:1986:EMM

[Sko86e] Gaylord V. Skogerboe. Ecological modelling, monitoring and implementation in irrigated agriculture. Ecological Modelling, 31(1–4):45–59, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900542.

Skogerboe:1988:SCA

[Sko88] Gaylord V. Skogerboe. Soil conservation: Assessing the national resources inventory: Committee on Conservation Needs and Opportunities, Board on Agriculture, National Research Council, National Academy Press, Washington, DC, U. S. A., 1986. Marketed and distributed by Wiley, Chichester, Great Britain. 114 pp., US$11.50 (paperback), ISBN 0-309-03675-5. Ecological Modelling, 42(1):85–86, July 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900956.

Svirezhev:1984:MMF

[SKV84] Yu. M. Svirezhev, V. P. Krysanova, and A. A. Voinov. Mathematical modelling of a fish pond ecosystem. Ecological Modelling, 21(4):315–337, March 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900668.

Samuels:1984:CSP

[SL84] William B. Samuels and Anthony Ladino. Calculations of seabird population recovery from potential oilspills in the mid-Atlantic region of the United States. Ecological Modelling, 21(1–2):63–84, January 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900255.

Shirazi:1984:MSB

[SLG84] Mostafa A. Shirazi, Bruce Lighthart, and James Gillett. A method for scaling biological response of soil microcosms. Ecological Modelling, 23(3):203–226, June 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084901017.

Summers:1981:SAE

[SM81] J. K. Summers and H. N. McKellar. A sensitivity analysis of an ecosystem model of estuarine carbon flow. Ecological Modelling, 13(4):283–301, September 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900338.

Steinmuller:1982:QAG

[SM82] Karlheinz Steinmüller and Eberhard Matthäus. Qualitative analysis of generalized Volterra models. Ecological Modelling, 17(2):91–106, October 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900461.

Smith:1980:TBE

[Smi80a] Richard Allmon Smith. The theoretical basis for estimating phytoplankton production and specific growth rate from chlorophyll, light and temperature data. Ecological Modelling, 10(3–4):243–264, August 1980. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380080900629.

Smith:1980:ASC

[Smi80b] T. P. Smith. Application of the simple charge-discharge metabolic model to transient behavior of experimental marine microcosms. Ecological Modelling, 10(1):13–29, June 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380080900769.

Smies:1983:SSBa

[Smi83a] Maarten Smies. Simulation of small bird populations. I. Development of a stochastic model. Ecological Modelling, 20 (4):259–277, December 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380083900364.

Smies:1983:SSBb

[Smi83b] Maarten Smies. Simulation of small bird populations. II. Reduced breeding in two British raptor species. Ecological Modelling, 20(4):279–296, December 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900376.

Smith:1987:CPG

[Smi87] Ian R. Smith. Control of population growth by events: an initial account. Ecological Modelling, 35(3–4):175–188, March 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087901116.

Steinhorst:1985:SDS

[SMN85] R. K. Steinhorst, P. Morgan, and L. F. Neuenschwander. A stochastic-deterministic simulation model of shrub succession. Ecological Modelling, 29(1–4):35–55, September 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900468.

Sueishi:1986:ACA

[SMR86] Tomitaro Sueishi, Tohru Morioka, and Christian Rouviere. Analysis and cartographical approach to the regional water utilization system in the Yodo River basin. Ecological Modelling, 31(1–4):315–327, May 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900724.

Somers:1988:ECC

[SO88] Greg L. Somers and Richard G. Oderwald. Estimating and constructing confidence intervals for spatial patterns between random and regular. Ecological Modelling, 44(1–2):57–72, December 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900828.

Somlyody:1982:WQM

[Som82] L. Somlyódy. Water-quality modelling: a comparison of transport-oriented and ecology-oriented approaches. Ecological Modelling, 17(3–4):183–207, December 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008290031X.

Somers:1984:RRM

[Som84] I. Somers. Renewable resource management: T. L. Vincent and J. M. Skowronski (Editors). Lecture Notes in Biomathematics, 40. Springer-Verlag, Berlin, Heidelberg, New York, 1981. xii + 236 pp., DM 28.50, approx. US$13.60. ISBN 3-540-10566-2. Ecological Modelling, 23(4):344–346, July 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084901303.

Swartzman:1984:SBE

[SR84] Gordon Swartzman and Kenneth A. Rose. Simulating the biological effects of toxicants in aquatic microcosm systems. Ecological Modelling, 22(1–4):123–134, April 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900115.

Stohlgren:1986:PML

[SR86] Thomas J. Stohlgren and Philip W. Rundel. A population model for a long-lived, resprouting chaparral shrub: Adenostoma fasciculatum. Ecological Modelling, 34(3–4):245–257, December 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900074.

Sarkar:1989:RHA

[SR89] A. K. Sarkar and A. B. Roy. Role of herbivore attack pattern in growth of plant populations. Ecological Modelling, 45(4):307– 316, June 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008990077X.

Skalski:1983:CPM

[SRM83] John R. Skalski, Douglas S. Robson, and Carrie L. Matsuzaki. Competing probabilistic models for catch-effort relationships in wildlife censuses. Ecological Modelling, 19(4):299–307, September 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380083900455.

Shapiro:1982:EOS

[SS82] Aleksandr Pavlovich Shapiro and Elena Iosifovna Skaletskaya. The existence of optimal strategy in the exploitation of limited biocenoses. Ecological Modelling, 17(2):113–120, October 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900485.

Sheppard:1987:SST

[SS87] Marsha I. Sheppard and S. C. Sheppard. A soil solute transport model evaluated on two experimental systems. Ecological Modelling, 37(3–4):191–206, July 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900251.

Saarenmaa:1988:AIM

[SSF^{+}88] H. Saarenmaa, N. D. Stone, L. J. Folse, J. M. Packard, W. E. Grant, M. E. Makela, and R. N. Coulson. An artificial intelligence modelling approach to simulating animal/habitat interactions. Ecological Modelling, 44(1–2):125–141, December 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900853.

Schaalje:1989:MIP

[SSJ89] G. Bruce Schaalje, R. L. Stinner, and D. L. Johnson. Modelling insect populations affected by pesticides with application to pesticide efficacy trials. Ecological Modelling, 47(3–4):233–263, September 15, 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900033.

Starfield:1981:SLC

[SSS81] A. M. Starfield, J. D. Shiell, and G. L. Smuts. Simulation of lion control strategies in a large game reserve. Ecological Modelling, 13(1–2):17–28, June 1981. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380081900053.

Stephenson:1988:PGP

[Ste88] D. Stephenson. Planning for groundwater protection: G. W. Page. Academic Press, Orlando, FL, U. S. A., 1987. 387 pp., US$45.00/£37.50. ISBN 0-12-543615-7. Ecological Modelling, 42(3–4):305, September 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900634.

Stocker:1981:OMW

[Sto81] Max Stocker. Optimization model for a wolf-ungulate system. Ecological Modelling, 12(3):151–172, April 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900557.

Stocker:1983:UPD

[Sto83] Max Stocker. Ungulate population dynamics and optimization models. Ecological Modelling, 18(2):121–139, February 1983. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380083900509.

Straskraba:1982:AEO

[Str82a] M. Straškraba. Aquatic ecosystems. An Operational research approach: Jan E. Beyer, University of Washington Press, Seattle and London, 1981, IX + 315 pp., US $25.00, ISBN: 295-95719-0. Ecological Modelling, 17(3–4):319–321, December 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900382.

Straskraba:1982:SSS

[Str82b] Milan Straškraba. Symposium: Simulation of systems in biology and medicine — Sisy, Praha (Czechoslovakia), 18–20 November 1980. Ecological Modelling, 17(2):53–56, October 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900412.

Straskraba:1983:CFC

[Str83] Milan Straškraba. Cybernetic formulation of control in ecosystems. Ecological Modelling, 18(2):85–97, February 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900480.

Straskraba:1984:EEM

[Str84a] M. Straškraba. Energy and ecological modelling: W. J. Mitsch, R. W. Bosserman and J. M. Klopatek (Editors). Developments in Environmental Modelling, Vol. 1. Proceedings of a symposium held from 20 to 23 April 1981 at Louisville, Kentucky. Elsevier, Amsterdam — Oxford — New York — Tokyo, 1981. 839 pp., US $139.50, Dfl. 300.-, ISBN 0-444-99731-8. Ecological Modelling, 23(3):265–267, June 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084901054.

Straskraba:1984:TSS

[Str84b] Milan Straškraba. Third symposium: Simulation of systems in biology and medicine — Sisy, Praha, Czechoslovakia, 22– 25 November 1982. Ecological Modelling, 26(1–2):3–7, December 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900875.

Straskraba:1987:PFN

[Str87a] Milan Straškraba. Positive feedback in natural systems: D. L. DeAngelis, W. Post, C. C. Travis. Springer-Verlag, Berlin/ Heidelberg/New York/Tokyo, 1986. xii + 290 pp., DM138. ISBN 3-540-15942-8. Ecological Modelling, 38(3–4):316–317, October 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087901074.

Straskraba:1987:EMF

[Str87b] Milan Straškraba. Ecological modelling at the fourth symposium ‘simulation of systems in biology and medicine’ — Sisy, Praha, Czechoslovakia, 12–14 November 1984. Ecological Modelling, 39(1–2):1–15, November 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900093.

Szyrmer:1987:TFE

[SU87] Janusz Szyrmer and Robert E. Ulanowicz. Total flows in ecosystems. Ecological Modelling, 35(1–2):123–136, February 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900949.

Sullivan:1988:EBC

[Sul88] Patrick J. Sullivan. Effect of boundary conditions, region length, and diffusion rates on a spatially heterogeneous predator–prey system. Ecological Modelling, 43(3–4):235–249, November 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900063.

Summers:1985:SMC

[Sum85] J. Kevin Summers. A simulation model of carbon and oxygen dynamics in a reservoir. Ecological Modelling, 28(4):279–309, August 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008590078X.

Summers:1989:SIE

[Sum89] J. Kevin Summers. Simulating the indirect effects of power plant entrainment losses on an estuarine ecosystem. Ecological Modelling, 49(1–2):31–47, December 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900422.

Seligman:1989:HPM

[SV89a] N. G. Seligman and H. Van Keulen. Herbage production of a Mediterranean grassland in relation to soil depth, rainfall and nitrogen nutrition: a simulation study. Ecological Modelling, 47(3–4):303–311, September 15, 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900082.

Svirezhev:1989:VML

[SV89b] Yuri M. Svirezhev and Alexey A. Voinov. Variability and management of large marine ecosystems: K. Sherman and L. M. Alexander (Editors). AAAS Selected Symposia Series, 99. Westview Press, Boulder, CO, 1986. xxvi + 319 pp., US$41.00. ISBN 0-8133-7145-7. Ecological Modelling, 45(2):153–154, April 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900902.

Svetlosanov:1985:PES

[Sve85] V. A. Svetlosanov. The problem of ecosystem stability and some applications of one of the stochastic methods in investigation of this problem. Ecological Modelling, 28(4):311–322, August 1985. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380085900791.

Svirezhev:1988:ETA

[Svi88a] Yuri M. Svirezhev. Ecosystem theory and application: Nicholas Polunin (Editor). Wiley, Chichester, Great Britain, 1986. xv + 445 pp., £49.95. ISBN 0-471-10274-1. Ecological Modelling, 41(1–2):153–154, April 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900531.

Svirezhev:1988:MEI

[Svi88b] Yuri M. Svirezhev. Mathematical ecology. An introduction: T. G. Hallam and S. A. Levin (Editors). Biomathematics, 17. Springer, Berlin/New York, 1986. xii + 457 pp., DM168.00. ISBN 3-540/0-387-13631-2. Ecological Modelling, 41(1–2):151– 153, April 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008890052X.

Svirezhev:1988:SAS

[Svi88c] Yuri M. Svirezhev. Systems analysis and simulation in wildlife and fisheries sciences: W. E. Grant. Wiley-Interscience, New York/Chichester, Great Britain, 1986. xiv + 338 pp., sterling 45.65. ISBN 0-471-89236-X. Ecological Modelling, 41(1–2):155– 156, April 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900543.

Svirezhev:1989:WMW

[Svi89] Yuri M. Svirezhev. Wetland modelling: W. J. Mitsch, M. Straškraba and S. E. Jørgensen (Editors). Developments in Environmental Modelling, 12. Elsevier, Amsterdam, 1988. ix + 227 pp., Dfl. 185.00. ISBN 0-444-42936-0. Ecological Modelling, 45(2):154–155, April 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900914.

Stocker:1984:DVU

[SW84] Max Stocker and C. J. Walters. Dynamics of a vegetationungulate system and its optimal exploitation. Ecological Modelling, 25(1–3):151–165, October 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900371.

Schmitt:1985:IER

[SW85] T. Schmitt and C. Wissel. Interdependence of ecological risk and economic profit in the exploitation of renewable resources. Ecological Modelling, 28(3):201–215, June 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900833.

Sauer:1984:IDR

[SWH84] R. H. Sauer, M. L. Warner, and W. T. Hinds. Indirect determination of rooting depth and permanent wilting point. Ecological Modelling, 21(1–2):109–124, January 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084900279.

Sharpe:1985:PBC

[SWPW85] Peter J. H. Sharpe, Joe Walker, Les K. Penridge, and Hsin-I Wu. A physiologically based continuous-time Markov approach to plant growth modelling in semi-arid woodlands. Ecological Modelling, 29(1–4):189–213, September 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900535.

Takeuchi:1986:DSE

[TA86] Yasuhiro Takeuchi and Norihiko Adachi. Dynamics and stability of ecological models. Ecological Modelling, 32(1–3):95– 104, June 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900190.

Taylor:1988:CPM

[Tay88] A. H. Taylor. Characteristic properties of models for the vertical distribution of phytoplankton under stratification. Ecological Modelling, 40(3–4):175–199, March 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900178.

Thorpe:1982:CSO

[TCL82] G. R. Thorpe, W. R. Cuff, and B. C. Longstaff. Control of Sitophilus oryzae infestation of stored wheat: an ecosystem model of the use of aeration. Ecological Modelling, 15(4):331– 351, June 1982. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380082900886.

Turner:1989:MEP

[TCS89] Monica G. Turner, Robert Costanza, and Fred H. Sklar. Methods to evaluate the performance of spatial simulation models. Ecological Modelling, 48(1–2):1–18, October 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900574.

Telegdi:1982:DMA

[TD82] László Telegdi and László Dávid. Description of the multi-annual eutrophication process of shallow lakes by a discrete time watershed development model. Ecological Modelling, 16(2–4):241–250, August 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900102.

Teramoto:1986:ICP

[Ter86a] Ei Teramoto. Island colonization and population structure of species pool. Ecological Modelling, 31(1–4):61–68, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900554.

Tercafs:1986:CSA

[Ter86b] R. Tercafs. A computer system to assist optimalization of land management. Ecological Modelling, 31(1–4):355–363, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008690075X.

Tamura:1980:IDL

[TH80] Hiroyuki Tamura and Efraim Halfon. Identification of a dynamic lake model by the group method of data handling: an application to Lake Ontario. Ecological Modelling, 11(2):81–100, November 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900885.

Tung:1988:PDC

[TH88] Yeou-Koung Tung and Wade E. Hathhorn. Probability distribution for critical do location in streams. Ecological Modelling, 42(1):45–60, July 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900919.

Tung:1989:DCL

[TH89] Yeou-Koung Tung and Wade E. Hathhorn. Determination of the critical locations in a stochastic stream environment. Ecological Modelling, 45(1):43–61, February 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900999.

Thomann:1984:PCE

[Tho84] Robert V. Thomann. Physio-chemical and ecological modeling the fate of toxic substances in natural water systems. Ecological Modelling, 22(1–4):145–170, April 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900139.

Tamura:1985:EEM

[TI85] Hiroyuki Tamura and Tsutomu Ishida. Environmentaleconomic models for total emission control of regional environmental pollution — input-output approach. Ecological Modelling, 30(3–4):163–173, December 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900663.

Takada:1986:SDD

[TI86] Takenori Takada and Yoh Iwasa. Size distribution dynamics of plants with interaction by shading. Ecological Modelling, 33(2–4):173–184, October 1986. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380086900396.

Tyagi:1983:PAL

[TN83] N. K. Tyagi and V. V. Dhruva Narayana. Planning for alkali land reclamation under rainfall uncertainty. Ecological Modelling, 20(4):243–258, December 1983. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380083900352.

Trenbath:1989:MBM

[Tre89] B. R. Trenbath. MIDAS, A bioeconomic model of a dryland farm system: R. S. Kingwell and D. J. Pannell (Editors). Pudoc, Wageningen, The Netherlands, 1987. Paperback, 207 pp., Dfl. 65.00/US$32.50. ISBN 90-220-0913-0. Ecological Modelling, 45(2):155–157, April 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900926.

Trigo:1987:FRA

[Tri87] Nuri Trigo. FLEX-REFLEX approach to ecological modeling. Ecological Modelling, 36(1–2):65–72, April 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900585.

Taylor:1987:HMP

[TSJ87] J. A. Taylor, R. W. Simpson, and A. J. Jakeman. A hybrid model for predicting the distribution of sulphur dioxide concentrations observed near elevated point sources. Ecological Modelling, 36(3–4):269–296, May 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900718.

Takasaki:1986:SDO

[TSOS86] Mitsuru Takasaki, Atsuhisa Sato, Mitsumasa Okada, and Ryuichi Sudo. A study on dissolved oxygen budgets in natural and artificial lakes. Ecological Modelling, 31(1–4):283– 292, May 1986. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380086900694.

Tyagi:1988:MST

[Tya88] N. K. Tyagi. Managing salinity through conjunctive use of water resources. Ecological Modelling, 40(1):11–24, January 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900993.

Uchmanski:1982:SME

[Uch82] Janusz Uchmański. Simulation model of energy flow through an animal organism: General form and applications. Ecological Modelling, 17(2):83–90, October 1982. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008290045X.

Ulanowicz:1988:IHL

[Ula88] Robert E. Ulanowicz. On the importance of higher-level models in ecology. Ecological Modelling, 43(1–2):45–56, October 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900725.

Umbach:1989:SES

[Umb89] Eberhard Umbach. Socio-economic systems as causal factors in the dynamics of ecosystems. Ecological Modelling, 46(3–4):305– 310, August 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900240.

Uchrin:1984:CSF

[UO84] Christopher G. Uchrin and William Ollinger. Continuous system and finite cell models: a statistical comparison. Ecological Modelling, 25(1–3):85–96, October 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900346.

Urchin:1984:MTP

[Urc84] Christopher G. Urchin. Modeling transport processes and differential accumulation of persistent toxic organic substances in groundwater systems. Ecological Modelling, 22(1–4):135– 143, April 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900127.

Upadhyaya:1988:RSO

[URC88] S. K. Upadhyaya, R. H. Rand, and J. R. Cooke. Role of stomatal oscillations on transpiration, assimilation and wateruse efficiency of plants. Ecological Modelling, 41(1–2):27–40, April 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900427.

Uri:1980:EUO

[Uri80] Noel D. Uri. Estimating undiscovered oil resources: an engineering approach. Ecological Modelling, 10(2):67–75, July 1980. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380080900642.

Usher:1981:GMR

[Ush81a] Michael B. Usher. Gradient modeling: Resource and fire management: Stephen R. Kessell. Springer-Verlag, Berlin– Heidelberg–New York, 1979, xv + 432 pp., U. S. $43.80, ISBN 3-540-90379-8. Ecological Modelling, 12(1–2):133–134, March 1981. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380081900284.

Usher:1981:GAT

[Ush81b] Michael B. Usher. The golden age of theoretical ecology: 1923–1940. Lecture notes in biomathematics, Vol. 22: Francesco M. Schuda and James R. Ziegler (Editors). Springer Verlag, Berlin — Heidelberg — New York, 1978, xii + 490 pp., illustrated, U.S. $21.50, ISBN 3-540-08769-9. Ecological Modelling, 11(4):292–293, February 1981. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900648.

Usher:1984:AAB

[Ush84a] M. B. Usher. Advances in Applied Biology: Vol. VI. T. H. Coaker (Editor). Academic Press, London/New York/Toronto/ Sydney/San Francisco, 1981. x + 332 pp., £25/US $51.50, ISBN 0-12-040906-2. Ecological Modelling, 23(3):268–269, June 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380084901078.

Usher:1984:AER

[Ush84b] M. B. Usher. Advances in ecological research, vol. 13: A. Macfadyen and E. D. Ford (Editors). Academic Press, London/New York, 1983. 382 pp., £27.00/ US$45.00. ISBN 0-12-013913-8. Ecological Modelling, 26(3–4):321–322, December 15, 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900802.

Usher:1984:PML

[Ush84c] M. B. Usher. Problems in management of locally abundant wild mammals: Peter A. Jewell and Sidney Holt (Editors). Academic Press, New York/London, 1982. xiv + 361 pp., £17.20/US $26.00 ISBN 0-12-385280-3. Ecological Modelling, 26(3–4):323–324, December 15, 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900826.

Usher:1984:SMM

[Ush84d] M. B. Usher. Soil microbiology: a model of decomposition and nutrient cycling: O. L. Smith. CRC Press, Boca Raton, FL, U. S. A., 1982. 273 pp., US$79.50/ US$90.00 (outside U. S. A.). ISBN 0-8493-5952-X. Ecological Modelling, 26(3–4):322–323, December 15, 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900814.

Usher:1985:PBP

[Ush85] Michael B. Usher. Population biology. Proceeding of the international conference held at the University of Alberta, 22– 30 June 1982, Edmonton, Alta., Canada: H. I. Freedman and C. Strobeck (Editors). Lecture Notes in Biomathematics, 52. Springer, Berlin/Heidelberg/New York/Tokyo, 1983. 440 pp., DM62.00/ approximately US$24.00. ISBN 3-540-12677-5/0-387-12677-5. Ecological Modelling, 28(3):235, June 1985. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380085900870.

Vandermeer:1984:EBA

[VAH^{+}84] John Vandermeer, Robert Ambrose, Michael Hansen, Hugh McGuinness, Ivette Perfecto, Cruz Phillips, Peter Rosset, and Brian Schultz. An ecologically-based approach to the design of intercrop agroecosystems: an intercropping system of soybeans and tomatoes in Southern Michigan. Ecological Modelling, 25(1–3):121–150, October 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008490036X.

VanStraalen:1987:TAS

[Van87] Nico M. Van Straalen. Turnover of accumulating substances in populations with weight-structure. Ecological Modelling, 36(3–4):195–209, May 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900688.

Vardavas:1987:MSV

[Var87a] Ilias Mihail Vardavas. Modelling the seasonal variation of net all-wave radiation flux and evaporation in a tropical wetdry region. Ecological Modelling, 39(3–4):247–268, December 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087900032.

Vardavas:1987:SMR

[Var87b] Ilias Mihail Vardavas. A simple model for rapidly computing terrestrial flux, solar flux and global mean surface temperature. Ecological Modelling, 35(3–4):189–210, March 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380087901128.

Vardavas:1988:SWB

[Var88a] Ilias Mihail Vardavas. A simple water balance daily rainfallrunoff model with application to the tropical Magela Creek catchment. Ecological Modelling, 42(3–4):245–264, September 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900609.

Varis:1988:TSA

[Var88b] Olli Varis. Temporal sensitivity of Aphanizomenon flos-aquae dominance — a whole-lake simulation study with input perturbations. Ecological Modelling, 43(3–4):137–153, November 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900014.

Vardavas:1989:FST

[Var89a] Ilias Mihail Vardavas. A Fibonacci search technique for model parameter selection. Ecological Modelling, 48(1–2):65–81, October 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900604.

Vardavas:1989:WBM

[Var89b] Ilias Mihail Vardavas. A water budget model for the tropical magela flood plain. Ecological Modelling, 46(3–4):165–194, August 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089900161.

Vavilin:1982:GMA

[Vav82] Vasilij A. Vavilin. Generalized model of aerobic biological treatment. Ecological Modelling, 17(2):157–173, October 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380082900527.

VanderPloeg:1987:IRE

[VBV87] S. W. F. Van der Ploeg, L. C. Braat, and W. F. J. Van Lierop. Integration of resource economics and ecology. Ecological Modelling, 38(1–2):171–190, September 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087900512.

Vertinsky:1987:EMR

[Ver87] Ilan Vertinsky. An ecological model of resilient decision making: an application to the study of public and private sector decision making in Japan. Ecological Modelling, 38(1–2):141–158, September 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900494.

Vincent:1982:GTA

[VG82] Thomas L. Vincent and Jonathan E. Gayek. A game theoretic analyses for renewable resource management. Ecological Modelling, 14(3–4):213–227, January 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900199.

VanStraten:1982:EAG

[VH82] Gerrit Van Straten and Sandor Herodek. Estimation of algal growth parameters from vertical primary production profiles. Ecological Modelling, 15(4):287–311, June 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900862.

Vidal:1989:ICS

[Vid89] R. V. V. Vidal. International conference on the systems analysis approach to environmental and natural resources management in the Baltic region, 26–29 September 1988, Gdańsk, Poland. Ecological Modelling, 45(2):161–162, April 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438008990094X.

Virtanen:1986:TDW

[VKDS86] Markku Virtanen, Jorma Koponen, Kim Dahlbo, and Juha Sarkkula. Three-dimensional water-quality-transport model compared with field observations. Ecological Modelling, 31 (1–4):185–199, May 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380086900633.

Volohonsky:1983:SLK

[VKS83] H. Volohonsky, A. Kaplanovsky, and S. Serruya. Storms on Lake Kinneret: Observations and mathematical model. Ecological Modelling, 18(2):141–153, February 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900510.

VanDuin:1989:MPO

[VL89] E. H. S. Van Duin and L. Lijklema. Modelling photosynthesis and oxygen in a shallow, hypertrophic lake. Ecological Modelling, 45(4):243–260, June 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900744.

Vlad:1988:HPS

[Vla88a] Marcel Ovidiu Vlad. A harvesting problem in structured population dynamics. Ecological Modelling, 41(3–4):229–246, June 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900300.

Vlad:1988:NNM

[Vla88b] Marcel Ovidiu Vlad. A new nonlinear model for the growth of age-structured populations living in patchy habitats. Ecological Modelling, 43(3–4):251–269, November 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900075.

Volohonsky:1981:FPZ

[Vol81] H. Volohonsky. Feeding process of zooplankton in terms of information flow. Ecological Modelling, 12(1–2):1–10, March 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380081900211.

Volohonsky:1982:FEF

[Vol82] Henri Volohonsky. Free energy flow in the process of collecting material for body build up of aquatic organisms. Ecological Modelling, 15(4):313–329, June 1982. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380082900874.

Volohonsky:1985:FBP

[Vol85] Henri Volohonsky. Form-building potencies of photons and the structural dynamics of ecosystems. Ecological Modelling, 28(1–2):139–154, May 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900183.

Volohonsky:1986:EMC

[Vol86] Henri Volohonsky. Ecosystem’s memory in the context of structural dynamics. Ecological Modelling, 33(1):59–75, September 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086901079.

Voinov:1984:MME

[VS84] A. A. Voinov and Yu. M. Svirezhev. A minimal model of eutrophication in freshwater ecosystems. Ecological Modelling, 23(4):277–292, July 1984. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438008490125X.

Voinov:1987:QME

[VT87] A. A. Voinov and A. P. Tonkikh. Qualitative model of eutrophication in macrophyte lakes. Ecological Modelling, 35(3–4):211– 226, March 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438008790113X.

Vavilin:1983:MCS

[VV83] V. A. Vavilin and V. B. Vasiliev. On the mechanism of consumption of simple subtrates by particles of activated sludge. Ecological Modelling, 18(1):27–34, January 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438008390073X.

Vavilin:1980:MBO

[VVK80] V. A. Vavilin, V. B. Vasiliev, and S. S. Kuzmin. Modelling the biochemical oxidation processes in artificial conditions of waste treatment plants. Ecological Modelling, 10(2):105–137, July 1980. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380080900666.

Vyhnalek:1987:IBA

[Vyh87] Vojtěch Vyhnálek. Interactions between algae and zooplankton in a continuous cultivation system. Ecological Modelling, 39(1–2):33–43, November 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900111.

Wahlgren:1981:ADA

[Wah81] Ulf Wahlgren. Age distributions and age-class equations for simple population-growth and prey-predator models. Ecological Modelling, 12(3):135–149, April 1981. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380081900545.

Walker:1985:CLC

[Wal85] Robert Walker. A comment on linear control problems in the theory of renewable resource exploitation. Ecological Modelling, 30(3–4):309–315, December 1985. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380085900730.

Walker:1987:RDR

[Wal87] Robert Walker. Regional development and renewable resource exploitation. Ecological Modelling, 37(3–4):303–316, July 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900317.

Weiner:1981:DNE

[WC81] Jacob Weiner and Paul T. Conte. Dispersal and neighborhood effects in an annual plant competition model. Ecological Modelling, 13(3):131–147, August 1981. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008190048X.

Wlosinski:1985:EWQ

[WC85] Joseph H. Wlosinski and Carol Desormeau Collins. Evaluation of a water quality model (CE-QUAL-R1) using data from a small Wisconsin reservoir. Ecological Modelling, 29(1–4):303–313, September 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900584.

Warwick:1988:EMR

[WC88]  John J. Warwick and William G. Cale.  Estimating

model reliability using data with uncertainty. Ecological Modelling, 41(3–4):169–181, June 1988. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380088900257.

Wellings:1988:PCO

[Wel88] Paul W. Wellings. Pest control: Operations and systems analysis in fruit fly management: M. Mangel, J. R. Carey and R. E. Plant (Editors). NATO ASI Ecological Sciences Series, 11. Springer, Berlin/New York, 1986.

465 pp., Hardcover, DM178.00. ISBN 0-387-16088-4. Ecological Modelling, 43(3–4):317–319, November 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900129.

Wilkie:1988:SML

[WF88] David S. Wilkie and John T. Finn. A spatial model of land use and forest regeneration in the Ituri forest of northeastern Zaire. Ecological Modelling, 41(3–4):307–323, June 1988. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380088900348.

Wallach:1989:MSE

[WG89] D. Wallach and B. Goffinet. Mean squared error of prediction as a criterion for evaluating and comparing system models. Ecological Modelling, 44(3–4):299–306, January 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380089900355.

Wagner:1984:BMS

[WGSC84] Terence L. Wagner, James A. Gagne, Peter J. H. Sharpe, and Robert N. Coulson. A biophysical model of southern pine beetle, Dendroctonus frontalis Zimmermann (Coleoptera: Scolytidae), development. Ecological Modelling, 21(1–2):125–147, January 1984. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380084900280.

Watanabe:1986:IBM

[WH86] Masataka Watanabe and Akira Harashima. Interaction between motile phytoplankton and Langmuir circulation. Ecological Modelling, 31(1–4):175–183, May 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900621.

Woolhouse:1987:TMMa

[WH87a] M. E. J. Woolhouse and R. Harmsen. A transition matrix model of seasonal changes in mite populations. Ecological Modelling, 37(3–4):167–189, July 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438008790024X.

Woolhouse:1987:TMMb

[WH87b] M. E. J. Woolhouse and R. Harmsen. A transition matrix model of the population dynamics of a two-prey — two-predator acarid complex. Ecological Modelling, 39(3– 4):307–323, December 1987. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380087900068.

Woolhouse:1989:TMM

[WH89] M. E. J. Woolhouse and R. Harmsen. A transition matrix model of European red mite (Panonychus ulmi) population dynamics in a managed apple orchard. Ecological Modelling, 46(3–4):269–282, August 1989. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380089900227.

White:1983:FMT

[WHB83] Gary C. White, Thomas E. Hakonson, and Kenneth V. Bostick. Fitting a model of tritium uptake by honey bees to data. Ecological Modelling, 18(3–4):241–251, April 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900169.

White:1984:MSMa

[Whi84a] E. G. White. A multispecies simulation model of grassland producers and consumers I. Validation. Ecological Modelling, 24(1–2):137–157, August 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900589.

White:1984:MSMb

[Whi84b] E. G. White. A multispecies simulation model of grassland producers and consumers II. Producers. Ecological Modelling, 24(3–4):241–262, September 1984. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380084900449.

White:1984:MSMc

[Whi84c] E. G. White. A multispecies simulation model of grassland producers and consumers III. Consumers. Ecological Modelling, 24(3–4):263–279, September 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900450.

Williams:1982:OSC

[Wil82] Byron K. Williams. Optimal stochastic control in natural resource management: framework and examples. Ecological Modelling, 16(2–4):275–297, August 1982. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380082900126.

Wu:1987:STD

[WMP^{+}87] H. Wu, K. W. J. Malafant, L. K. Pendridge, P. J. H. Sharpe, and J. Walker. Simulation of two-dimensional point patterns: Application of a lattice framework approach. Ecological Modelling, 38(3–4):299–308, October 1987. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380087901025.

Wroblewski:1981:MTT

[WO81] J. S. Wroblewski and James J. O’Brien. On modeling the turbulent transport of passive biological variables in aquatic ecosystems. Ecological Modelling, 12(1–2):29–44, March 1981. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380081900235.

Wolaver:1983:RHT

[Wol83] Thomas G. Wolaver. The role of herbivory in terrestrial ecosystem dynamics. Ecological Modelling, 18(3–4):223–233, April 1983. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380083900145.

Woolhouse:1988:DBT

[Woo88] M. E. J. Woolhouse. On the dynamical behaviour of transition matrix population models. Ecological Modelling, 42(1): 61–74, July 1988. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380088900920.

Worton:1987:RMH

[Wor87] B. J. Worton. A review of models of home range for animal movement. Ecological Modelling, 38(3–4):277–298, October 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087901013.

Wu:1987:SPA

[WRSZ87] Hsin-I Wu, Edward J. Rykiel, Peter J. H. Sharpe, and Guangzhou Zou. A statistical physics approach to nearest neighbor distibution for individuals of finite size. Ecological Modelling, 36(1–2):73–87, April 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900597.

Wheeler:1980:TSM

[WS80] G. L. Wheeler and M. J. Sale. TELOC: a simulation model of the effects of lead on corn (zea maize) growth. Ecological Modelling, 11(2):141–150, November 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900903.

Wong:1989:FGS

[WSML89] Isaac Wong, D. A. Swayne, C. R. Murthy, and D. C. L. Lam. Fast graphical simulations of spills and plumes for application to the great lakes. Ecological Modelling, 47(1–2):161–173, September 1989. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380089901154.

Wu:1985:EFT

[WSWP85] Hsin-I Wu, Peter J. H. Sharpe, Joe Walker, and Les K. Penridge. Ecological field theory: a spatial analysis of resource interference among plants. Ecological Modelling, 29(1–4):215–243, September 1985. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380085900547.

Wulff:1980:ACS

[Wul80] Fredrik V. Wulff. Animal community structure and energy budget calculations of a Daphnia magna (Straus) population in relation to the rock pool environment. Ecological Modelling, 11(3):179–225, December 1980. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380080900836.

Wulff:1986:SMI

[Wul86] Fredrik Wulff. Synthesis and modelling of intermittent estuaries. A case study from planning to evaluation: W. R. Cuff and M. Tomczak, Jr. (Editors) Lecture Notes on Coastal and Estuarine Studies, 3. Springer, Berlin/Heidelberg/New York/ Tokyo, 1983. viii + 302 pp., DM49.80 soft cover. ISBN 3-540-12681-3. Ecological Modelling, 33(1):79–80, September 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086901092.

Wolfe:1986:CSM

[WZE86] John R. Wolfe, Ronald D. Zweig, and David G. Engstrom. A computer simulation model of the solar-algae pond ecosystem. Ecological Modelling, 34(1–2):1–59, November 1986. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380086900785.

Yearsley:1987:MCD

[YL87] John R. Yearsley and Dennis P. Lettenmaier. Model complexity and data worth: an assessment of changes in the global carbon budget. Ecological Modelling, 39(3–4):201–226, December 1987. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380087900019.

Young:1987:UBE

[YW87] J. H. Young and Linda J. Willson. Use of Bose–Einstein statistics in population dynamics models of arthropods. Ecological Modelling, 36(1–2):89–99, April 1987. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380087900603.

Zhi-Bin:1986:PEN

[ZBBLCSTL86] Cheng Zhi-Bin, Li Bai-Lian, Zou Chao-Shun, and Que Tai-Lin. Parameter estimation of a nonlinear population model with two parameters, growth of a yeast population as an example. Ecological Modelling, 34(3–4):191–196, December 1986. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380086900037.

Zingales:1984:CMU

[ZMRB84] Franco Zingales, Alessandro Marani, Andrea Rinaldo, and Giuseppe Bendoricchio. A conceptual model of unit-mass response function for nonpoint source pollutant runoff. Ecological Modelling, 26(3–4):285–311, December 15, 1984. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380084900735.

Zucchetto:1983:EFH

[Zuc83] James Zucchetto. Energy and the future of human settlement patterns: Theory, models and empirical considerations. Ecological Modelling, 20(2–3):85–111, November 1983. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380083900017.

Zucchetto:1988:EDT

[Zuc88] J. Zucchetto. Environmental decisionmaking in a transboundary region: Alison Rieser, Judith Spiller and David VanderZwang (Editors). Lecture Notes on Coastal and Estuarine Studies, 20, Springer, Berlin, Federal Republic of Germany, 1986. 209 pp. DM48.00. ISBN 3-540-96446-0. Ecological Modelling, 42(3–4):306–307, September 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380088900646.

Zucchetto:1989:CMP

[Zuc89a] J. Zucchetto. A climate modelling primer: A. Henderson-Sellers and K. McGuffie. Wiley, Chichester, Great Britain, 1987. xv + 217 pp., hardcover, £28.50. ISBN 0-471-91462-2. Ecological Modelling, 45(1):69–70, February 1989. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089901014.

Zucchetto:1989:ANS

[Zuc89b] J. Zucchetto. Agricultural nonpoint source pollution: Model selection and application: A. Giorgini and F. Zingales (Editors). Developments in Environmental Modelling, 10. International Society for Ecological Modelling, Copenhagen/ Elsevier, Amsterdam, 1986. 409 pp., Dfl.190. ISBN 0-444-99505-6. Ecological Modelling, 45(1):70–71, February 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089901026.